### v1

# Inflation with vector fields revisited: heavy entropy perturbations and primordial black holes

Abstract
We revisit inflation coupled with vector fields employing kinetic coupling in the comoving gauge. It is known that there is a cumulative effect $IN^{2}$ on curvature power spectrum. For a large number of e-foldings $N$, this contribution is so significant that it could violate observational constraints when the ratio of kinetic energy between vector fields and inflaton $I$ is not extremely small. In this paper, we explore the regime where $I\gg 1$, a realm that has not been extensively explored due to the limitations of perturbative methods. We found that the entropy perturbation becomes heavy in this regime and the cumulative effect decays away on super-horizon scales. Consequently, the power spectrum retains its scale invariance in the decoupling limit. By straightforwardly integrating out the heavy modes near horizon-crossing, we derive a low-energy effective field theory describing a massless adiabatic perturbation with an imaginary speed of sound $c_{s}^{2}=-1/3$. Namely, the inflation with vector fields presents a potential mechanism for generating primordial black holes.
1 Introduction
Inflationary physics has proven successful in phenomenologically explaining the large-scale structure of our universe. The simplest inflationary model, single-field inflation, predicts nearly scale-invariant, Gaussian, and adiabatic primordial perturbations. These predictions are confirmed by Cosmic Microwave Background (CMB) observations [1, 2, 3]. However, the single-field model encounters theoretical challenges. These inflationary models are expected to be effective field theories derived from UV completed theories, such as string theory. Despite the anticipated suppression of higher-dimension irrelevant operators by the UV cutoff, they play a crucial role at the low-energy scale during inflation. The UV sensitivity of inflation gives rise to issues like the $\eta$-problem [4, 5] and the super-Planckian problem of a single large field [6, 7, 5], disrupting the flatness of the scalar field potential.
The study of introducing extra fields during inflation has attracted attention to evade the issues of single-field inflation. In particular, the study of multi-scalar fields exhibiting strongly non-geodesic motion has been extensively examined in recent years [8, 9, 10, 11, 12, 13, 14]. In contrast, our focus is on U(1) gauge fields, commonly known as vector fields, as additional fields during inflation. This choice stems not only from the ubiquity of gauge fields as essential components of the standard model (SM) and beyond but also from their rich phenomenological implications in cosmology [15, 16, 17, 18, 19]. The breaking of the conformal coupling between vector fields and gravity is necessary to generate sufficiently strong vector fields [19]. Analogous to the non-geodesic motion of scalar fields, vector fields should exhibit background dynamics and participate alongside scalar fields in driving inflation. The inflationary phase undergoes a transition to new attractors if the potential is steep enough [20], which is very similar to the hyperbolic inflation [8]. This kind of models was firstly constructed to explain the statistical anisotropy in the CMB [21, 22, 23, 24]. The vector field is coupled to gravity through $f^{2}(\phi)F_{\mu\nu}F^{\mu\nu}$. After suitable choices of $f(\phi)$, the dynamics of background vector fields provides an effective potential for the slow-roll scalar field.
Under this new attractor, the power spectrum of curvature perturbation is sourced by the vector field, yielding an additional contribution
$\mathcal{P}_{\zeta}=\mathcal{P}_{\zeta}^{(0)}\left[1+\Theta(\theta)IN^{2}\right],$ (1.1)
where $N\equiv\ln{(-k\tau_{e})}$ represents the e-folding number of observed mode elapsed since horizon-crossing and $I=2\epsilon(\rho_{A}/V)$ is the fraction kinetic energy of the vector field. For one U(1) vector field, $\Theta(\theta)=24\sin^{2}\theta$, where $\theta$ is the angle between mode and direction of the vector field [22, 25]. The term $IN^{2}$ accountable for the amplitude of statistical anisotropy and should be rigorously constrained by CMB observations to $I<10^{-7}$ for a sufficiently lengthy inflationary period $N\sim 60$ [22, 25].The exceedingly small parameter implies the presence of a very small but non-zero vector field, potentially giving rise to fine-tuning challenges [26, 27]. Studies have revealed that employing multiple vector fields can mitigate the anisotropy, eventually converging to an isotropic attractor [24]. For the isotropic configuration of a triad of vector fields, $\Theta(\theta)=16$, relaxing the constraint on $I$ [28]. However, it still needs to be constrained to $I<10^{-5}$due to the tilt of the power spectrum caused by the e-folding number $N$. This is a cumulative effect, i.e., the curvature is sourced by non-zero vector fields and evolves on super-horizon scales.
However, this does not imply that results for large $I$ are undesirable. If we require vector fields to handle a sufficiently large field displacement, indicating that $I$ is not a small parameter, it appears that the second term in (1.1) becomes significant and violates the adiabatic and scale-invariant initial conditions. However, this result lacks reliability because the perturbative method breaks down in this regime. It has been demonstrated that for $I\gtrsim 0.2$, the calculation by using in-in formalism starts to deviate from the numerical results and it is difficult to track the dynamics near and outside the horizon analytically [29, 30]. Numerical findings indicate that the power spectrum does not actually evolve on super-horizon scales [30, 20]. In this paper, we revisit this dilaton-gauge coupling model in comoving gauge and thoroughly explore the $I\gg 1$ case with the assistance of effective field theory (EFT). We will show, the entropy perturbation is super heavy in this regime and the theory can be described by a low-energy EFT of a massless adiabatic perturbation. $IN^{2}$ does not occur in this regime; instead, it is replaced by an exponentially enhanced power spectrum, dependent solely on the parameter $I$. This enhancement also provides a mechanism for generating primordial black holes.
2 Inflation with vector fields
We are considering the spatially flat Friedmann-Lemaitre-Robertson-Walker (FLRW) universe, which is characterized by the scale factor $a(t)$ and the Hubble variable is $H(t)=\dot{a}/a$. The Arnowitt-Deser-Misner(ADM) formalism of the metric is [31]
$ds^{2}=-N^{2}dt^{2}+q_{ij}\left(N^{i}dt+dx^{i}\right)\left(N^{j}dt+dx^{j}\right),$ (2.1)
where $N$ is the lapse function, $N^{i}$ is the shift vector and $g_{ij}$ is the metric of hypersurface. We consider the model of inflaton and $U(1)$ gauge fields, which is called dilaton-gauge inflation in this paper
$S=\int d^{4}x\sqrt{-{g}}\left[-\frac{1}{2}\partial^{\mu}\phi\partial_{\mu}\phi-\frac{1}{4}f_{ab}(\phi)F^{a}_{\ \mu\nu}F^{b\mu\nu}\right],$ (2.2)
where $f_{ab}$ is the kinetic coupling and only depends on the inflaton.
2.1 Background
We consider one dilaton field $\phi$ as inflaton and a triplet of $U(1)$ fields and take an isotropic configuration of background evolution, which can be realized by the choice of $U(1)$ fields as following [32, 33]
[…]
The identification of the spatial indices and internal indices of the field space allows us to achieve the internal gauge transformation of three $U(1)$ gauge fields by the $O(3)$ rotation of the three spatial dimensions in real space. We mention that the discussion doesn’t rely on the number of the $U(1)$ fields. If we have only one $U(1)$ field and kinetic coupling with dilaton, there may be anisotropy left in our universe when the energy density of $U(1)$ gauge field is not so small [22, 25]. Although we will see, the theoretical parameterization of power spectrum is different when this energy density is large enough, we still consider isotropic case so that we can only treat scalar modes at linear order.
Under this configurations the Friedmann equations are given by
$M_{\text{pl}}^{2}H^{2}\left(3-\epsilon_{\phi}-\frac{3}{2}\epsilon_{A}\right)=V(\phi),\ \ \ \ \ \ \ \epsilon=\epsilon_{\phi}+\epsilon_{A},$ (2.4)
where $\epsilon\equiv-\dot{H}/H^{2}$, $\epsilon_{\phi}\equiv\dot{\phi}^{2}/(2M_{\text{pl}}^{2}H^{2})$ and $\epsilon_{A}\equiv f^{2}\dot{\mathbb{A}}^{2}/(a^{2}M_{\text{pl}}^{2}H^{2})$. It’s also useful to define $\dot{\sigma}^{2}\equiv\dot{\phi}^{2}/2+f^{2}\dot{\mathbb{A}}^{2}/a^{2}$ as the “kinetic” energy of the matter fields. Then another gravitational equation can be written as
$H^{2}M_{\text{pl}}^{2}\epsilon=\dot{\sigma}^{2}.$ (2.5)
The equations of motion of the matter fields are
$\ddot{\phi}+3H\dot{\phi}+V_{\phi}-3f_{\phi}f\frac{\dot{\mathbb{A}}^{2}}{a^{2}}=0,\ \ \ \ \ \ \ \ddot{\mathbb{A}}+\left(2\frac{\dot{f}}{f}+H\right)\dot{\mathbb{A}}=0.$ (2.6)
In this paper we assume that $\dot{\phi}<0$ and $\dot{\mathbb{A}}>0$. We also define a useful variable, the ratio of the kinetic energy of $U(1)$ fields and dilaton [^1] The parameter $I$ in [28] is related to $h$ in our paper as $I\simeq 2h^{2}$.
$h\equiv\sqrt{\frac{\epsilon_{A}}{2\epsilon_{\phi}}}.$ (2.7)
But here the ratio $h$ needn’t to be a small parameter. In fact, we will consider the larger $h\gg 1$ on small scales for PBH’s production. We will also define the slow-roll parameter $\epsilon_{h}\equiv\dot{h}/(Hh)$, $\eta_{h}\equiv\dot{\epsilon}_{h}/(H\epsilon_{h})$, $\eta_{\phi}\equiv\ddot{\phi}/(H\dot{\phi})$ and $\eta_{A}\equiv\dot{\epsilon}_{A}/(H\epsilon_{A})$. Taking time derivative on the Friedmann equation one can obtain
$V_{\phi}=M_{\text{pl}}H^{2}\left(6\epsilon+2\epsilon_{\phi}\eta_{\phi}+\frac{3}{2}\epsilon_{A}\eta_{A}-3\epsilon\epsilon_{A}\right).$ (2.8)
Inserting $V_{\phi}$ into the equation of motion of $\phi$ we obtain
$\frac{f_{\phi}}{f}M_{\text{pl}}\sqrt{2\epsilon_{\phi}}=2\left(1+\frac{1}{4}\eta_{A}-\frac{1}{2}\epsilon\right).$ (2.9)
2.2 Quadratic action in comoving gauge
Thanks to the isotropic background and we will work in the quadratic action and the linear fluctuations of the theory hence the vector and tensor fluctuations are decoupled from scalar ones at linear order. After using the helicity decomposition the scalar parts of the $U(1)$ fields can be written as
[…]
We do not distinguish indices $a$ and $i$ here because we identified the spatial rotation symmetry with the internal global $O(3)$ symmetry of the space of the gauge fields. The gauge fields enjoys local $U(1)$ symmetry $A^{a}_{\ \mu}\rightarrow A^{a}_{\ \mu}+\partial_{\mu}\rho^{a}$, where $\rho^{a}$ is a arbitrary function. One can fix the scalar gauge freedom of $U(1)$ fields by $\mathbb{M}=0$ [20]. The $\mathbb{Y}$ is the constrain d.o.f and we will eliminate it latter after fixing the gauge freedoms of gravity.
The scalar fluctuations of gravity and scalar fields can be written as
$N=1+\alpha,\ \ \ \ N_{i}=\partial_{i}\beta,\ \ \ \ g_{ij}=a^{2}(t)e^{2\zeta}\delta_{ij},\ \ \ \ \phi=\phi(t)+\delta\phi.$ (2.11)
There are four gauge freedoms of the time and spatial coordinate choice. If we have scalar field, in addition to vector fields we still have one physical d.o.f in the theory. The spatial flat gauge is usually used to study the dilaton-gauge theory. In this gauge, this physical d.o.f is the fluctuation of the scalar field $\delta\phi$. On the other hand, the “comoving” gauge is also used in the two-scalar fields [34, 35]. In this gauge, the adiabatic fluctuation is set to zero. The physical d.o.f are the curvature and entropy fluctuations.
In this paper we adopt the similar comoving gauge for the dilaton-gauge model. The adiabatic fluctuation is given by $Q_{\sigma}\equiv-\sqrt{2}\dot{\sigma}\delta u$, where $\delta u$ is the velocity potential given by $\delta T^{0}_{i}\equiv(\rho+p)\partial_{i}\delta u$. For the isotropic one-form gauge fields, we have $\delta T^{0}_{\ i}=-\dot{\phi}\partial_{i}\delta\phi-(2f^{2}\dot{\mathbb{A}}/a^{2})\partial_{i}\delta\mathbb{A}$ [20]. Here $\rho$ and $p$ are the total energy density and pressure of the system and in our model we have $\rho+p=\dot{\phi}^{2}+2f^{2}\dot{\mathbb{A}}^{2}/a^{2}=2\dot{\sigma}^{2}$. It’s useful to define the variable $\delta Q=\sqrt{2}f\delta\mathbb{A}/a$. Then the adiabatic fluctuation is given by
$Q_{\sigma}\equiv\frac{\dot{\phi}/\sqrt{2}}{\dot{\sigma}}\delta\phi+\frac{f\dot{\mathbb{A}}/a}{\dot{\sigma}}\delta Q.$ (2.12)
We choose the comoving gauge
$Q_{\sigma}=0.$ (2.13)
This is the unitary gauge in the gauge theory, whose Nambu-Goldstone boson is eaten by the metric, as we will see in the next section. Then we have fixed all gauge freedom of the theory. The d.o.f $\mathbb{U}$ in the vector fields is the magnetic fluctuation and decouple with any other modes in our configuration. It only contribute to the isocurvature fluctuation hence we ignore it in the following discussion. We also define the entropy fluctuations. Defining [36, 28]
$\cos{\mathbb{\vartheta}}\equiv\frac{\dot{\phi}/\sqrt{2}}{\dot{\sigma}},\ \ \ \ \ \ \ \ \ \ \sin{\vartheta}\equiv\frac{f\dot{\mathbb{A}}/a}{\dot{\sigma}},$ (2.14)
The adiabatic fluctuation reads $Q_{\sigma}=\cos{\vartheta}\delta\phi+\sin{\vartheta}\delta Q$. This fluctuation follows the direction of the classical “trajectory” of the matter fields. The entropy fluctuation is defined by
$\mathcal{F}\equiv\cos{\vartheta}\delta Q-\sin{\vartheta}\delta\phi,$ (2.15)
which is similar to the entropy fluctuation in the two-scalar fields. But here the entropy fluctuations is not equivalent to the isocurvature one because we still have another contribution $\mathbb{U}$, which is decoupled with $\zeta$ and $\mathcal{F}$.
After imposing the comoving gauge, the equation of motion of the non-dynamical d.o.f $\alpha$ and $\mathbb{Y}$ can be derived and the solutions read
$\alpha=$ $\frac{\dot{\zeta}}{H},$ (2.16)
$\partial^{2}\mathbb{Y}=$ $\dot{\mathbb{A}}\left(\zeta-\frac{\dot{\zeta}}{H}\right)-\frac{a}{\sqrt{2}f}\left(\frac{\mathcal{F}}{\sqrt{1+2h^{2}}}\right)^{.}$
$-\frac{a}{\sqrt{2}f}\left(H-\frac{\dot{f}}{f}+4f_{\phi}\frac{\dot{\mathbb{A}}}{a}h\right)\frac{\mathcal{F}}{\sqrt{1+2h^{2}}}.$ (2.17)
The d.o.f $\beta$ is totally eliminated from the quadratic action when we use the constrain equation of $\alpha$ hence we don’t need to write down its specific expression. Then after using the background equations (2.6) and (2.9) to eliminate the $V$, $V_{\phi}$ and $f_{\phi}$, and inserting the constrain equations into the quadratic Lagrangian, we obtain
$\mathcal{L}^{(2)}=$ $a^{3}M_{\text{pl}}^{2}\epsilon\left(\dot{\zeta}^{2}-\frac{1}{a^{2}}(\partial\zeta)^{2}-m_{\sigma}^{2}\zeta^{2}\right)+\frac{a^{3}}{2}\left(\dot{\mathcal{F}}^{2}-\frac{1}{a^{2}}(\partial\mathcal{F})^{2}-m_{s}^{2}\mathcal{F}^{2}\right),$
$-2a^{3}\dot{\sigma}h\left(\mathcal{A}\dot{\zeta}+\mathcal{B}H\zeta\right)\mathcal{F},$ (2.18)
where the mass $m_{\sigma}$, $m_{s}$ and the coupling $\mathcal{A}$, $\mathcal{B}$ are
$\ \ \ \ \mathcal{A}=4,\ \ \ \ \ \ \ \mathcal{B}=\frac{16h^{2}}{1+2h^{2}},\ \ \ \ \ \ \ m_{\sigma}^{2}=\frac{16h^{2}}{1+2h^{2}}H^{2},$
$m_{s}^{2}=\frac{8h^{4}-40h^{2}-4}{1+2h^{2}}H^{2}-\frac{12h^{4}-8h^{2}-2}{1+2h^{2}}\frac{f_{\phi\phi}}{f}M_{\text{pl}}^{2}H^{2}\epsilon_{\phi},$ (2.19)
where we have discard slow-roll suppressed terms in the slow-low limit and also assumed $|V_{\phi\phi}|/H^{2}\ll 1$. This is the decoupling limit $\epsilon\to 0$ in the EFT. Then $\epsilon_{\phi}$ and $\epsilon_{A}$ are also vanishing because they are positive quantities.
We find that the curvature and entropy fluctuations are both massive. The mass term of $\zeta$ is only dependent on $h$. For the mass term of $\mathcal{F}$, the second term can be regards as the contribution from the curved field space, whose curvature is given by $R_{\text{fs}}=-f_{\phi\phi}/f$. If we assume $h=0$ initially, where we are in the single-field inflation regime, the entropy mass $m_{s}^{2}$ is still non-vanishing and have large negative contribution if the curvature of field space is negatively large enough. In the slow-roll regime, the second derivative of $f$ in the entropy mass becomes $f_{\phi\phi}/f\simeq f_{\phi}^{2}/f^{2}$. Then using (2.9) the mass square of the entropy fluctuation reads
$m_{s}^{2}\simeq-\frac{8h^{2}\left(3+2h^{2}\right)}{1+2h^{2}}H^{2}.$ (2.20)
Therefore the entropy fluctuation has imaginary mass. This is similar to the so-called geometrical destabilization [11, 12], which is firstly studied in the multi-scalar inflation and then in the inflation with multiple vector fields [20]. The differences are that the curvature fluctuation $\zeta$ is massive and we have an additional coupling $\zeta\mathcal{F}$. Hence we next investigate the mode solutions of this Lagrangian.
2.3 Evolution of the adiabatic mode
We are interested in the super-horizon evolution of the adiabatic mode. We can solve the system from the equations of motion, which is shown in spatially flat gauge in [20]. Here we directly find the solutions from the symmetries of Lagrangian. The quadratic action (2.2) in the long-wavelength limit, where the momentum terms are ignored, can be reduced to
$\mathcal{L}^{(2)}=a^{3}M_{\text{pl}}^{2}\epsilon\dot{\zeta}^{2}+\frac{a^{3}}{2}\dot{\mathcal{F}}^{2}-\frac{a^{3}}{2}m_{s}^{2}\left(\mathcal{F}+2\dot{\sigma}h\frac{\mathcal{A}\dot{\zeta}+\mathcal{B}H\zeta}{m_{s}^{2}}\right)^{2},$ (2.21)
where we obtained the $\zeta\dot{\zeta}$ term through integral by parts. We see that this Lagrangian is invariant under shift transformation
$\zeta\to\zeta+\zeta_{0},\ \ \ \ \ \ \ \mathcal{F}\to\mathcal{F}-\frac{2\dot{\sigma}h\mathcal{B}H}{m_{s}^{2}}\zeta_{0}.$ (2.22)
This symmetry reveals that $\zeta$ and $\mathcal{F}$ have constant solutions obeying
$\mathcal{F}_{0}=-\frac{2\dot{\sigma}h\mathcal{B}H}{m_{s}^{2}}\zeta_{0}.$ (2.23)
These constant solutions exist for any value of $h$. And it’s obvious that these mode solutions are not the correction $IN^{2}$ in (1.1). We next explore the Lagrangian for small $h$.
[…]
$\zeta=\zeta^{(0)}+\zeta^{(1)}+\cdots,\ \ \ \ \ \ \ \ \ \mathcal{F}=\mathcal{F}^{(0)}+\mathcal{F}^{(1)}+\cdots$ (2.24)
We see in Lagrangian (2.2), the mass terms are $\mathcal{O}(h^{2})$, the $\mathcal{A}$ term is $\mathcal{O}(h)$ and the $\mathcal{B}$ term is $\mathcal{O}(h^{3})$. If we only consider super-horizon solutions up to $\mathcal{O}(h)$, the leading-order Lagrangian can be written as
$\mathcal{L}^{(2)}=$ $\frac{a^{3}}{2}\left(\sqrt{2\epsilon}M_{\text{pl}}\dot{\zeta}-\sqrt{2}\mathcal{A}Hh\mathcal{F}\right)^{2}+\frac{a^{3}}{2}\dot{\mathcal{F}}^{2}+\mathcal{O}(h^{2}).$ (2.25)
We found at leading order of $h$, the Lagrangian is invariant under
$\dot{\zeta}\to\dot{\zeta}+\dot{\zeta}^{(1)},\ \ \ \ \ \ \ \mathcal{F}\to\mathcal{F}+\frac{\sqrt{\epsilon}M_{\text{pl}}}{\mathcal{A}hH}\dot{\zeta}^{(1)}$ (2.26)
The first order $\dot{\zeta}^{(1)}$ is sourced from zero-th order of entropy perturbations $\mathcal{F}^{(0)}$. This source from entropy modes is a cumulative effect hence is dependent on the e-folding number after horizon-crossing
[…]
where $N_{c}$ is the time of horizon-crossing and $N_{k}\equiv N_{e}-N_{c}$ is the e-folding number elapsed since horizon-crossing. For small $h$ this is a small correction to the adiabatic mode. The zero-th order entropy perturbation is decoupled with the adiabatic one and massless in the slow-roll limit. Hence the zeroth-order power spectrum of entropy perturbation is simple $\mathcal{P}_{\mathcal{F}}^{(0)}=H^{2}/(4\pi^{2})$. Then the leading-order correction to the power spectrum of curvature perturbation can be estimated to
$\delta\mathcal{P}_{\zeta}\simeq 32h^{2}N^{2}\cdot\frac{H^{2}}{8\pi^{2}\epsilon M_{\text{pl}}^{2}},\ \ \ \ \ (h\ll 1)$ (2.28)
which is consistent with the result calculated through in-in formalism [28]. These correction is proportional to square of e-folding number $N\sim 60$, which is a large number for inflation. Hence for anisotropic configuration of vector fields, there is very strong constrain from CMB on the parameter $h$. The extreme small $h$ may lead to fine-tuning problem [26, 27]. However, these modes is heavy when $h$ is large enough hence dilute rapidly during inflation. To see this, we next explore the large $h$ case.
3 EFT under massive entropy mode
In the EFT formalism, the physical scalar d.o.f at low-energy scale is the Nambu-Goldstone (NG) boson particle coming from breaking the time translation symmetry $t\to t+\xi^{0}$ of the de Sitter space by a clock field $\phi(t)$, i.e., the homogeneous inflaton. One can choose the slice of $t$ that the inflaton has no excitation all the time, i.e. the unitary gauge. In this gauge, the only dynamical scalar d.o.f at low-energy is the scalar part of graviton $g_{ij}=a^{2}(t)e^{2\zeta}\delta_{ij}$, where the scalar $\zeta(t,\boldsymbol{x})$ is the spatial curvature of the time slice. The EFT of a single scalar field in the unitary gauge can be written as [38]
$\mathcal{L}_{\text{EFT}}=\sqrt{-g}\left[\frac{M_{\text{pl}}^{2}}{2}R+M_{\text{pl}}^{2}\dot{H}g^{00}-M_{\text{pl}}^{2}\left(3H^{2}+\dot{H}\right)+\cdots\right]+\mathcal{L}_{\text{n-u}},$ (3.1)
where the dots stand for the higher-order terms of $(1+g^{00})$ and we have ignored the extrinsic curvature fluctuation $\delta K_{\mu\nu}$. $\mathcal{L}_{\text{n-u}}$ is the non-universal parts and for $\omega^{2}\ll M^{2}+p^{2}$ [39]
$\mathcal{L}_{\text{n-u}}=\sqrt{-g}\sum_{n=2}\frac{M_{n}^{4}}{n!}\left[(1+g^{00})\frac{M^{2}}{M^{2}-\tilde{\partial}^{2}}\right]^{n-1}(1+g^{00})+\cdots,$ (3.2)
where $M$ is the mass of heavy fields and ~$\tilde{\partial}\equiv\partial/a$. These operators are non-local generally due to the hierarchy between the frequency and momentum when the speed of sound is small $c_{s}\ll 1$. There exist regime in energy scales that the dispersion is non-linear [39, 40]. If $\omega\sim p$ the momentum is also $p\ll M$ and we can ignore the non-local operators.
In order to transform to the spatial flat gauge $g_{ij}=a^{2}(t)\delta_{ij}$, one can perform a time diffeomorphisms $t\to\tilde{t}=t-\pi(t,\boldsymbol{x})$. Then at linear-order the NG boson is related to the curvature fluctuation as $\zeta=-H\pi$. The time component of the metric transforms as
$g^{00}=\frac{\partial x^{0}}{\partial\tilde{x}^{\mu}}\frac{\partial x^{0}}{\partial\tilde{x}^{\nu}}\tilde{g}^{\mu\nu}=\left(1+\dot{\pi}\right)^{2}\tilde{g}^{00}+2\left(1+\dot{\pi}\right)\partial_{i}\pi\tilde{g}^{0i}+\tilde{g}^{ij}\partial_{i}\pi\partial_{j}\pi$ (3.3)
In the low-energy scale $\omega\sim p\ll M$, we can write down the action in the re-parameterization of time. Dropping the tildes for simplicity the Lagrangian is given by
$\mathcal{L}_{\text{EFT}}=$ $\sqrt{-g}\Bigg{\{}\frac{M_{\text{pl}}^{2}}{2}R-M_{\text{pl}}^{2}\left[3H^{2}(t+\pi)+\dot{H}(t+\pi)\right]$
$+M_{\text{pl}}^{2}\dot{H}(t+\pi)\left[\left(1+\dot{\pi}\right)^{2}g^{00}+2\left(1+\dot{\pi}\right)\partial_{i}\pi g^{0i}+\partial^{i}\pi\partial_{i}\pi\right]+$
$+\frac{M_{2}^{4}}{2}\left[1+(1+\dot{\pi})^{2}g^{00}+2(1+\dot{\pi})\partial_{i}\pi g^{0i}+\partial^{i}\pi\partial_{i}\pi)\right]^{2}+\cdots\Bigg{\}}$ (3.4)
Expanding the action up to second order of fluctuations, the constrain of the lapse function is solved as $\alpha=\epsilon H\pi$ [41]. In the dilaton-gauge theory, from the quadratic action in the spatial flat gauge (A), constrain $\alpha$ is given by $\alpha\propto Q_{\sigma}$, which means that the unitary gauge is the vanishing adiabatic fluctuation at linear order.
In the decoupling limit, the NG boson is decoupled with gravity and becomes a dynamical scalar mode. For the inflation, this decoupling limit is the slow-roll limit $\epsilon\to 0$. Then the quadratic Lagrangian of NG boson reduces to
$\mathcal{L}_{\text{EFT}}=a^{3}M_{\text{pl}}^{2}H^{2}\frac{\epsilon}{c_{s}^{2}}\left[\dot{\pi}^{2}-\frac{c_{s}^{2}}{a^{2}}(\partial\mathcal{\pi})^{2}\right],$ (3.5)
where the speed of sound is defined by
$\frac{1}{c_{s}^{2}}\equiv 1-\frac{2M_{2}^{4}}{M_{\text{pl}}^{2}\dot{H}}.$ (3.6)
The modification of the speed of sound is provided by the higher operator $M_{2}$ and we will show soon that this can be obtained by integrating out the heavy fluctuation at low energy in the dilaton-gauge model.
3.1 Modified dispersion relation
From the quadratic Lagrangian (2.2) we can write down the linear equations of motion
$\ddot{\zeta}+\left(3+\eta\right)H\dot{\zeta}+\left(\frac{k^{2}}{a^{2}}+m_{\sigma}^{2}\right)\zeta=$ $\frac{1}{a^{3}\epsilon}\left(a^{3}\frac{\dot{\sigma}h\mathcal{A}}{M_{\text{pl}}^{2}}\mathcal{F}\right)^{.}-\frac{\dot{\sigma}hH\mathcal{B}}{\epsilon M_{\text{pl}}^{2}}\mathcal{F},$
¨$\ddot{\mathcal{F}}+3H\dot{\mathcal{F}}+\left(\frac{k^{2}}{a^{2}}+m_{s}^{2}\right)\mathcal{F}=$ $-2\dot{\sigma}h\left(\mathcal{A}\dot{\zeta}+\mathcal{B}H\zeta\right).$ (3.7)
We are interested in the decoupling limit where the nontrivial solution $\zeta=\text{constant}$ is permitted hence we drop out the small slow-roll parameters. Then we have $\mathcal{A}=4$ and $\mathcal{B}=m_{\sigma}/H^{2}$. We regard $\dot{\sigma}$ and $h$ as constants. To find the dispersion relation we study the short-wavelength limit where the Hubble friction terms can be disregarded. We have noted that there is also a mass term for curvature fluctuation. The upper bound of the mass is $m_{\sigma}^{2}\lesssim 8H^{2}$, i.e., the mass of $\zeta$ is at most of the same order as the energy scale of inflation. In the short-wavelength limit the contribution of $m_{\sigma}$ term can also be ignored. Then the equations of motion read
$\ddot{\zeta_{c}}+p^{2}\zeta_{c}=$ $\sqrt{2}hH\mathcal{A}\dot{\mathcal{F}},$
¨$\ddot{\mathcal{F}}+p^{2}\mathcal{F}+m_{s}^{2}\mathcal{F}=$ $-\sqrt{2}hH\mathcal{A}\dot{\zeta}_{c},$ (3.8)
where $p=k/a$ is the physical momentum and $\zeta_{c}\equiv M_{\text{pl}}\sqrt{2\epsilon}\zeta$ is the canonical variable. The schematic solutions of these equations can be written as [42]
$\zeta_{c}=\zeta_{+}e^{i\omega_{+}t}+\zeta_{-}e^{i\omega_{-}t},$
$\mathcal{F}=\mathcal{F}_{+}e^{i\omega_{+}t}+\mathcal{F}_{-}e^{i\omega_{-}t},$ (3.9)
where the two frequencies $\omega_{\pm}$ are given by
$\omega_{\pm}^{2}=\frac{m_{s}^{2}}{2c_{s}^{2}}+p^{2}\pm\frac{m_{s}^{2}}{2c_{s}^{2}}\sqrt{1+\frac{4\left(1-c_{s}^{2}\right)p^{2}}{m_{s}^{2}c_{s}^{-2}}}$ (3.10)
and the speed of sound is given by
$\frac{1}{c_{s}^{2}}=1+\frac{32H^{2}h^{2}}{m_{s}^{2}}.$ (3.11)
The modification of the dispersion relation of fluctuations is provided by the interaction with $\dot{\zeta}{\mathcal{F}}$. The heavy entropy fluctuation implies the hierarchy of the two frequencies $\omega_{-}\ll\omega_{+}$, which requires $4(1-c_{s}^{2})p^{2}\ll m_{s}^{2}c_{s}^{-2}$. Hence we have a cut-off scale of EFT (3.5) on momentum $p_{\text{UV}}$ and the EFT is only valid below this scale. The UV scale is given by
$p_{\text{UV}}^{2}=\frac{|m_{s}^{2}c_{s}^{-2}|}{4\left(1-c_{s}^{2}\right)}.$ (3.12)
In the low energy scales, the two scalar modes, which are corresponded to two frequencies, oscillate coherently and we integrate out the heavy mode to eliminate the high-frequency mode from the dynamics of the curvature fluctuation.
3.2 Integrating out the heavy $\mathcal{F}$
The fluctuation is heavy for $h\gg 1$ and we also have $|m_{s}^{2}|\gg m_{\sigma}^{2}$ as we have mentioned before. For positive $m_{s}^{2}$, the operator from coupling $\dot{\zeta}\mathcal{F}$ would reduce the speed of sound of the fluctuation [42]. However from (2.20) the mass square is negative due to the contribution from kinetic coupling $f(\phi)$. We can ignored the slow-roll suppressed terms and the entropy mass is given by (2.20). Then the speed of sound in this decoupling limit is
$\frac{1}{c_{s}^{2}}=1-\frac{4\left(1+2h^{2}\right)}{3+2h^{2}}.$ (3.13)
The speed of sound $c_{s}^{2}\simeq-1/3$ is imaginary for heavy entropy mass $h\gg 1$. The imaginary speed of sound has been studied in the two-scalar inflation recently [12, 14, 43]. And the entropy mass-square in the two-scalar models is also imaginary hence is tachyonic before the sound horizon crossing $p^{2}c_{s}^{2}\sim H^{2}$. The curvature fluctuation sourced by the entropy fluctuation also experiences transient instability before horizon crossing. After the horizon-crossing, the dynamics of fluctuations are influenced by the background. In other words, the mass term $m_{\sigma}$ and the Hubble terms become important outside the horizon.
Figure 1: Evolution of entropy perturbation $\mathcal{F}$ for $h=10$. An instability before horizon-crossing and constant evolution on super-horizon are presented. The scales are arbitrary.
We have noted that the speed of sound is $|c_{s}|\sim\mathcal{O}(1)$ for large $h$ hence the non-linear dispersion regime is narrower, and we can directly investigate the linear dispersion regime. In the heavy limit $|m_{s}^{2}|\gg\omega^{2}$ we can expand the operator
$\frac{1}{m_{s}^{2}-\square}=\frac{1}{m_{s}^{2}}+\frac{\square}{m_{s}^{4}}+\cdots,\ \ \ \ \ \ \text{where}\ \ \square\equiv-\frac{\partial^{2}}{\partial t^{2}}-3H\frac{\partial}{\partial t}+\frac{\partial^{2}}{a^{2}}.$ (3.14)
Then the leading-order of the linear equation of motion of $\mathcal{F}$ is given by
$\mathcal{F}_{\text{LO}}=-\frac{2\dot{\sigma}h(\mathcal{A}\dot{\zeta}+\mathcal{B}H\zeta)}{m_{s}^{2}}.$ (3.15)
We show the consistent of the leading-order contribution of $\mathcal{F}$ in Figure 1 for large enough $h$. The evolution of curvature perturbation on large scales is sourced by non-adiabatic perturbation $\dot{\zeta}=-\delta P_{\text{nad}}H/(\rho+p)$ [44, 45, 46]. Then from the leading-order solution of entropy perturbation, the non-adiabatic mode can be represented as
$\delta P_{\text{nad,LO}}=\frac{m_{s}^{2}\dot{\sigma}}{Hh\mathcal{A}}\mathcal{F}_{\text{LO}}+\frac{2\dot{\sigma}^{2}\mathcal{B}}{\mathcal{A}}\zeta.$ (3.16)
The adiabatic initial condition ensures that the r.h.s. should be relatively small enough on large scales. We can see that the entropy fluctuation doesn’t decay outside the horizon, but is proportional to the curvature fluctuation if $\zeta=\text{constant}$ on super-horizon.
We can find that in the EFT of single field, we have only this constant curvature perturbation on large scales. In the slow-roll limit the entropy mass $m_{s}$ can be calculated by the (2.20). From (3.11) the EFT is only valid under scale
$p_{\text{UV}}=\frac{3}{4}|m_{s}|.$ (3.17)
Then after inserting the equation (3.15) into the quadratic action (2.2) we obtain
$\mathcal{L}^{(2)}_{\text{LO}}=a^{3}M_{\text{pl}}^{2}\frac{\epsilon}{c_{s}^{2}}\left[\dot{\zeta}^{2}-\frac{c_{s}^{2}}{a^{2}}\left(\partial\zeta\right)^{2}\right],$ (3.18)
where $c_{s}$ is given by (3.13). We have used the integration by parts to eliminate the $\dot{\zeta}\zeta$ terms in the action. Then we found that the mass term of curvature fluctuation is cancelled and $\zeta$ is conserved on super-horizon scales.
Figure 2: The adiabaticity of super-horizon ($|k\tau_{e}|=10^{-15}$) curvature perturbation. As $h$ increases from very small values, the perturbations become non-adiabatic. However, when $h$ exceeds $\mathcal{O}(0.1)$, the perturbation quickly becomes adiabatic.
Different from multi-scalar inflation, where the non-decaying non-adibatic modes will source to curvature modes so that $\zeta$ evolves on super-horizon. In dilation-gauge inflation, we have constant $\zeta$ for non-decaying $\mathcal{F}$. In other words, $\mathcal{F}$ doesn’t play the role of non-adiabatic mode in dilaton-gauge inflation. For mid-value $h\lesssim\mathcal{O}(0.1)$, where $|m_{s}|\sim H$, the sub-leading terms in $\delta P_{\text{nad}}$ become important and the curvature perturbation evolves on super-horizon. These non-adiabatic contributions are from the cumulative modes (2.27). We show this non-adiabatic occurrence for mild $h\lesssim\mathcal{O}(0.1)$ in Figure 2.
One of the most important findings from this result is that the cumulative modes (2.27) decay away for large $h$. Because for large $h$, the modes (3.15) on super-horizon are exactly the constant solutions of the adiabatic mode (2.23). The decaying modes can be also obtained through solving the equation of motion (3.1), which has been studied in spatially flat gauge in [20]. The decaying mode is $\zeta\propto(-k\tau)^{(3-\sqrt{9-96h^{2}})/2}$. For small $h$ we can expand this solution and find that the leading correction is proportional to e-folding number $\ln(-k\tau)=N$, which is the solution (2.27).
4 Primordial black holes
The remained adiabatic mode experiences a exponential growth before horizon-crossing due to the imaginary speed of sound. Therefore the power spectrum is enhanced exponentially. If vector fields switch on only at late time, The power spectrum is enhanced on small scales so the formations of PBHs in radiation era is possible [47, 48]. And these PBHs can be one of candidates of the dark matter [49, 50, 51] and supermassive black holes in our universe [52].
We approach to the PBHs without any specific models ($V(\phi)$ and $f(\phi)$). We here only consider the switch-on of the vector fields at late time $N_{f}$ during inflation and see the features of power spectrum on small scales for generating PBHs. We choose the horizon-crossing time of the CMB scale $k_{\text{CMB}}=0.05\text{Mpc}^{-1}$ as $N_{\text{CMB}}=0$. Assuming that $h$ has peak at $N_{f}$. The curvature modes with different scales that cross the horizon at different time have different contributions to the power spectrum. We can define the corresponding scale
$k_{f}|c_{s}|=a(N_{f})H(N_{f}).$ (4.1)
The modes $k\gg k_{f}$ that cross the horizon far after the switch-off time experienced nothing hence the power spectrum is the same as the single-field inflation $\mathcal{P}_{\zeta}^{(0)}=H^{2}/(8\pi\epsilon M_{\text{pl}}^{2})$ on small scales.
On the other hand, the modes $k\ll k_{f}$ that cross the horizon far before the switch-on time will be impacted by the entropy fluctuation on super-horizon. The entropy fluctuation doesn’t decay outside the horizon hence always transfer to the curvature fluctuation. Hence on these large scales, the power spectrum is shifted so different from the single-field one.
The interesting scales are the modes $k\sim k_{f}$ that cross the horizon near the switch-on time $N_{f}$ have large enhancements due to the large instability of the entropy fluctuation. Hence we expect that the power spectrum of the curvature fluctuation has exponential enhancements near the scale $k_{f}$. We can see from Figure 1 that the mode takes a few e-folding to grow before frozen-in outside the horizon. Hence there are two kinds of power spectrum, the broad ($\delta N\gtrsim\ln{(h_{0})}$) and the sharp ($\delta N\lesssim\ln{(h_{0})}$) switch-on of $h(N)$ [53]. For the shape case, the curvature modes stop growing before crossing the horizon.
4.1 Board case
For this case, we have enough e-folding of $h(N)$ for exponential growth of the modes. The EFT is valid at time $N_{k}$ satisfying
$\frac{k}{a(N_{v})}=\frac{3}{4}|m_{s}(N_{v})|.$ (4.2)
Then the most contributions to the power spectrum are the modes with $N_{v}\sim N_{f}$. We here consider the Gaussian profile of the function
$h_{G}(N)=h_{0}e^{-(N-N_{f})^{2}/(2\delta N^{2})},$ (4.3)
where the profile is controlled by two parameters $(h_{0},\delta N)$. On super-horizon, the entropy modes do not decay and still couple to the curvature ones. Hence we should discuss large-scale modes $k\ll k_{f}$ and small-scale modes $k\sim k_{f}$ separately.
Modes with $\boldsymbol{k\gtrsim k_{f}}$
To characterize the peak of power spectrum with these modes, it’s also useful to introduce a dimensionless parameter $\mathcal{I}$ [43, 54]. The quantization of the curvature fluctuation with imaginary speed of sound is quite different from the real one [43]. The difference is that unlike the Bunch-Davies vacuum, the negative frequency modes are non-zero for initial state of EFT due to the different commutation conditions of the mode functions. Taking the super-horizon limit, the leading contribution of the power spectrum is
$\Delta_{\zeta}(k)\equiv\frac{\mathcal{P}_{\zeta}(k)}{\mathcal{P}_{\zeta}^{(0)}(k)}\simeq e^{2\mathcal{I}(h)}\big{|}_{N_{k}},$ (4.4)
where the exponential factor is given by (see Appendix B for derivation)
$\mathcal{I}(h)\simeq\sqrt{2}\pi\left(2-\sqrt{\frac{1+6h^{2}}{1+2h^{2}}}\right)h.$ (4.5)
For modes who start to experiences transient instability earlier, the amplitude will be larger after crossing the horizon. And we need to use a more UV EFT to describe these modes on deeper sub-horizon scales.
Figure 3: The numerical and analytical results (4.4) of power spectrum for the Gaussian case, where $h_{0}=8$ and $\delta N=2$.
The scale of peak $k_{p}$ is the one that has maximal $h$ at time when the modes is growing at $N_{v}$, i.e., $N_{p_{k}}=N_{v}$. Then we have
$k_{p}=k_{f}\cdot\frac{3}{4}\sqrt{\frac{8h_{0}^{2}\left(3+2h_{0}^{2}\right)}{1+2h_{0}^{2}}}.$ (4.6)
We show the numerical and analytical results of our calculation in the Figure 3. One thing we should mention is that (4.4) is not valid for all range of $h(N)$ because the computation is under the vaild of EFT so the mass term of curvature is small compared to the heavy entropy one (or other, large $h(N)$). Hence this estimation work more well for $|m_{s}^{2}(N)|\gtrsim 8H^{2}$ in our Gaussian-profile model.
Modes with $\boldsymbol{k\ll k_{f}}$
As Figure 3 shown, for the l.h.s of the peak, where the switch occurs after crossing the horizon, the power spectrum is governed by the long range contributions. Hence we can see a shift on the large scale due to the coupling with entropy fluctuation on large scales. We numerically found the amplitude of power spectrum on large scale as the function of $h$ and $\delta N$
$\Delta_{\zeta}^{\text{(LS)}}=\left(-0.00585h\cdot\delta N+0.0537h+0.333\delta N+1.95\right)^{2}.$ (4.7)
The amplitude of large-scale curvature perturbation depends on $h$ and $\delta N$ linearly and the amplitude of power spectrum $\Delta_{\zeta}^{\text{(LS)}}\gtrsim 10$, i.e., large-scale amplitude is always greater than the one of single-field inflation.
4.2 Sharp case
For very small $\delta N$ the Gaussian profile can be approximated to a delta function. For analytically computable, we use the top-hat profile
$h_{\text{TH}}=h_{0}\left[\theta(t-t_{1})-\theta(t-t_{2})\right].$ (4.8)
The wave number $k_{f}\simeq He^{H(t_{1}+t_{2})/2}$ is the scale that the modes cross the horizon during the $h\neq 0$. After the switch-off, the curvature and entropy modes are decoupled. Then the solutions of curvature modes are [55]
$\hat{\zeta}_{c}(t,\boldsymbol{k})=$ $\Big{[}E_{\zeta}u_{k}(t)+F_{\zeta}u_{k}^{*}(t)\Big{]}\hat{a}_{\zeta}(\boldsymbol{k})$
$+\Big{[}G_{\zeta}u_{k}(t)+H_{\zeta}u_{k}^{*}(t)\Big{]}\hat{a}_{\mathcal{F}}(\boldsymbol{k})+\text{h.c.}(-\boldsymbol{k}),$ (4.9)
where the annihilation and creation operators of two scalar modes satisfying the commutation relations $[\hat{a}_{\alpha}(\boldsymbol{k}),\hat{a}^{\dagger}_{\beta}(\boldsymbol{k}^{\prime})]=(2\pi)^{3}\delta_{\alpha\beta}\delta(\boldsymbol{k}-\boldsymbol{k}^{\prime})$ , $[\hat{a}^{\dagger}_{\alpha}(\boldsymbol{k}),\hat{a}^{\dagger}_{\beta}(\boldsymbol{k}^{\prime})]=[\hat{a}_{\alpha}(\boldsymbol{k}),\hat{a}_{\beta}(\boldsymbol{k}^{\prime})]=0$ and $u_{k}$ are the mode functions with Bunch-Davies initial conditons
$u_{k}(t)=\frac{iH}{\sqrt{2k^{3}}}\Big{[}1+ik\tau(t)\Big{]}e^{-ik\tau(t)}.$ (4.10)
Then the power spectrum of curvature perurtbation is given
$\Delta_{\zeta}=|E_{\zeta}-F_{\zeta}|^{2}+|G_{\zeta}-H_{\zeta}|^{2}.$ (4.11)
The non-trivial period is the $h_{\text{TH}}\neq 0$ (which is denoted as region-II), where two kinds of modes strongly couple to each other. We also discuss large-scale modes $k\ll k_{f}$ and small-scale modes $k\sim k_{f}$ separately.
Figure 4: The numerical and analytical results (C) of power spectrum for the top-hat case, where $\delta N=0.4$ and $h_{0}=24$. The envelope (4.14) is shown as orange dashed curve.
Modes with $\boldsymbol{k\gtrsim k_{f}}$
We simply use the conformal time as the function of $t$: $\tau(t)=-e^{-Ht}/H$. During $t_{1}$ and $t_{2}$ with $\delta t\equiv t_{2}-t_{1}\ll H^{-1}$, one can discard the Hubble friction of the equation of motions. After $t>t_{1}$, the coupling is switch-on. The duration of $h\neq 0$ is too short $\delta t\ll H^{-1}$ that we can ignore the evolution of the scale factor. Hence we can ignore the Hubble frictions and curvature mass. Then we can use the equations of motion (3.1) and the WKB-type solutions
$\hat{\zeta}_{c}^{\text{II}}(t,\boldsymbol{k})=\sum_{\pm}$ $\Big{(}A_{\pm}^{\zeta}e^{i\omega_{\pm}t}+B_{\pm}^{\zeta}e^{-i\omega_{\pm}t}\Big{)}\hat{a}_{\zeta}(\boldsymbol{k})+\Big{(}C_{\pm}^{\zeta}e^{i\omega_{\pm}t}+D_{\pm}^{\zeta}e^{-i\omega_{\pm}t}\Big{)}\hat{a}_{\mathcal{F}}(\boldsymbol{k})$
[…]
^$\hat{\mathcal{F}}^{\text{II}}(t,\boldsymbol{k})=\sum_{\pm}$ $\Big{(}A_{\pm}^{\mathcal{F}}e^{i\omega_{\pm}t}+B_{\pm}^{\mathcal{F}}e^{-i\omega_{\pm}t}\Big{)}\hat{a}_{\zeta}(\boldsymbol{k})+\Big{(}C_{\pm}^{\mathcal{F}}e^{i\omega_{\pm}t}+D_{\pm}^{\mathcal{F}}e^{-i\omega_{\pm}t}\Big{)}\hat{a}_{\mathcal{F}}(\boldsymbol{k})$
[…]
where the frequencies are given by (3.10). From the equations of motion, we have relations of the coefficients
$A_{\pm}^{\mathcal{F}}(C_{\pm}^{\mathcal{F}})=i\frac{\omega_{\pm}^{2}-p^{2}}{\sqrt{2}h\mathcal{A}H\omega_{\pm}}A_{\pm}^{\zeta}(C_{\pm}^{\zeta}),$
$B_{\pm}^{\mathcal{F}}(D_{\pm}^{\mathcal{F}})=-i\frac{\omega_{\pm}^{2}-p^{2}}{\sqrt{2}h\mathcal{A}H\omega_{\pm}}B_{\pm}^{\zeta}(D_{\pm}^{\zeta}).$
During the short region-II, where the scale factor is nearly invariant, we have $a=a(N_{f})=k_{f}/H$ hence $p=(k/k_{f})H$ in the above equations. We have eight unknown coefficients $A_{\pm}^{\zeta}$, $B_{\pm}^{\zeta}$, $C_{\pm}^{\zeta}$ and $D_{\pm}^{\zeta}$ in the region-II and we have eight boundary conditions at $t_{1}$ hence all these coefficients can be completely solved. Then at $t_{2}$ we also have eight boundary condictions and the coefficients of (C) can be solved to (see Appendiex C)
$E_{\zeta}$ $\simeq\frac{\left(S_{-}+\kappa\right)^{2}}{4\kappa S_{-}}e^{i2\kappa\sinh(\delta N/2)-iS_{-}\delta N},$
$F_{\zeta}$ $\simeq-\frac{S_{-}^{2}-\left(\kappa+i\right)^{2}}{4k}e^{i2\kappa\cosh(\delta N/2)}\cdot i2\sin\left(S_{-}\delta N\right),$
$G_{\zeta}$ $\simeq\sqrt{2}h\frac{\left(\kappa^{2}+iS_{-}+\kappa S_{-}\right)\left(i-\kappa-S_{-}\right)}{\kappa S_{-}S_{+}^{2}}e^{i2\kappa\sinh(\delta N/2)-iS_{-}\delta N},$
$H_{\zeta}$ $\simeq\sqrt{2}h\frac{\left(\kappa^{2}+iS_{-}+\kappa S_{-}\right)\left(i+\kappa-S_{-}\right)}{\kappa S_{-}S_{+}^{2}}e^{i2\kappa\cosh(\delta N/2)-iS_{-}\delta N},$ (4.13)
where $\kappa\equiv k/k_{f}$, $\delta N\equiv H(t_{2}-t_{1})$ and $S_{\pm}\equiv\omega_{\pm}/H$. We show the agreement between numerical and completed analytical results (see appendix C) in Figure 4. The power spectrum has exponential enhancement near $k_{f}$ and also rapidly oscillate during these scales. We found $E_{\zeta}$ and $F_{\zeta}$ are dominated so the power spectrum can be simply written as $\Delta_{\zeta}\simeq\Delta_{\text{env}}\times\left(\text{rapid oscillation}\right)$, where the envelope is
$\Delta_{\text{env}}=\left|\frac{e^{-i2S_{-}\delta N}\left(\kappa^{2}-|S_{-}|^{2}\right)}{4|S_{-}|^{2}}\right|$ (4.14)
and the rapid oscillation is characterized by $\text{Exp}\left(i2\kappa e^{\pm\delta N/2}\right)$, which is only related to the period of non-zero $h_{\text{tp}}$. The enhancement of power spectrum is provided by low-frequency $\omega_{-}$ and period $\delta N$ together. The scale of peak can be nearly determined by the exponential factor in $\Delta_{\text{env}}$, which is estimated by
$k_{p}=k_{f}\cdot\frac{h_{0}\sqrt{30+104h_{0}^{2}+56h_{0}^{4}}}{2(1+2h_{0}^{2})}.$ (4.15)
These calculations are only valid near scale $k\gtrsim k_{f}$ and for $k\ll k_{f}$ $\Delta_{\zeta}=1$, which miss the matching with the numerical results.
Modes with $\boldsymbol{k\ll k_{f}}$
Figure 5: In the top-hat case, the analytical result of large-scale power spectrum as function of $\delta N$ (4), where $h_{0}=24$. The envelope (4.19) is shown as orange dashed curve.
On large scales, the momentum terms in equations of motion can be ignored and the mass terms and all coupling terms become important. The equations of motion can be solved as
$\hat{\zeta}_{c}^{\text{II}}(\tau,\boldsymbol{k})=$ $\sum_{i=\zeta,\mathcal{F}}\left[c_{0}^{i}+c_{1}^{i}x^{3}+c_{+}^{i}x^{1+p_{+}}+c_{-}^{i}x^{1+p_{-}}\right]\hat{a}_{i}(\boldsymbol{k})+\text{h.c.}(-\boldsymbol{k}),$
^$\hat{\mathcal{F}}^{\text{II}}(\tau,\boldsymbol{k})=$ $\sum_{i=\zeta,\mathcal{F}}\left[d_{0}^{i}+d_{1}^{i}x^{3}+d_{+}^{i}x^{1+p_{+}}+d_{-}^{i}x^{1+p_{-}}\right]\hat{a}_{i}(\boldsymbol{k})+\text{h.c.}(-\boldsymbol{k}),$ (4.16)
where $x\equiv-k\tau$ and the coefficients satisfy
$d_{0}^{i}=\frac{2\sqrt{2}h}{3+2h^{2}}c_{0}^{i},$ $\ \ \ \ \ d_{1}^{i}=-\frac{1}{\sqrt{2}h}c_{1}^{i},$
$d_{+}^{i}=\frac{p_{-}(1+2h^{2})-2}{2\sqrt{2}h(3+4h^{2})}c_{+}^{i},$ $\ \ \ \ \ d_{-}^{i}=\frac{p_{+}(1+2h^{2})-2}{2\sqrt{2}h(3+4h^{2})}c_{-}^{i}.$ (4.17)
We have eight unknown coefficients $c_{0}^{i}$, $c_{1}^{i}$, $c_{+}^{i}$ and $c_{-}^{i}$ ($i=\zeta,\mathcal{F}$) in the region-II and eight boundary conditions at $t_{1}$ hence all these coefficients can be completely solved. Then at $t_{2}$ the coefficients of (C) can also be solved [^2] We found that this computation breaks when $x\lesssim 10^{-5}$.. The results are extremely tedious, but we can simple them under some approximations, see appendix C. The results should be real because the perturbations become classical on super-horizon. So we can only consider the real parts
$\text{Re}[E_{\zeta}]$ $\simeq\frac{1}{27}\left[-5+32\cosh{\left(3\delta N\right)}\right],$
$\text{Re}[F_{\zeta}]$ $\simeq\frac{32}{27}\sinh\left(3\delta N\right),$
$\text{Re}[G_{\zeta}]$ $\simeq\frac{e^{-3\delta N/2}}{3\sqrt{3}}\left(2e^{3\delta N}+1\right)\sin{\left(2\sqrt{6}h\delta N\right)},$
$\text{Re}[H_{\zeta}]$ $\simeq\frac{e^{-3\delta N/2}}{3\sqrt{3}}\left(2e^{3\delta N}-1\right)\sin{\left(2\sqrt{6}h\delta N\right)}.$ (4.18)
We found that, for fixed $\delta N$, $|G_{\zeta}-H_{\zeta}|$ oscillates around zero when changing $h$. The oscillation is characterized by $\sin(2\sqrt{6}h\delta N)$ and is not significant on amplitude. This is because, on large scales, the dynamics of curvature perturbation is almost a constant and weakly depends on the frequency. For large $h$, the dynamics is determined by an EFT of massless curvature perturbation with sound speed $c_{s}\simeq-1/3$. On the other hand, the amplitudes of $|E_{\zeta}-F_{\zeta}|$ and $|G_{\zeta}-H_{\zeta}|$ exponentially depend on $\delta N$. We can smooth out the oscillation and only estimate the envelope of the power spectrum as
$\Delta_{\text{env}}^{\text{(LS)}}=\frac{1}{729}\left(25+1024e^{-6\delta N}-212e^{-3\delta N}\right),$ (4.19)
as Figure 5 shows. We can see $\Delta_{\zeta}<1$, in other words, the amplitude of power spectrum on large scales is always smaller than the one of single-field inflation and is not sensitive to $h$ for sharp case. We should mention that this result is also valid for board case with top-hat profile of $h$. For $\delta N\gg 1$, we have $\Delta_{\text{env}}^{\text{(LS)}}\simeq 0.034$, which is not sensitive to $\delta N$ either.
4.3 PBH formation
After treating the power spectrum on large and small scales for board and sharp case, we can also compare them with the amplitude of the curvature power spectrum required for PBH formation. The fraction of PBHs at formation time can be estimated in a model-independent way [56]
$\beta_{\text{PBH}}=3.7\times 10^{-9}\left(\frac{\gamma}{0.2}\right)^{-1/2}\left(\frac{g_{*,f}}{10.75}\right)^{1/4}\left(\frac{M_{\text{PBH}}}{M_{\odot}}\right)^{1/2}f_{\text{PBH}},$ (4.20)
where $\gamma$ can be calculated analytically as $\gamma=0.2$ [50], $g_{*,f}$ is the effective number of relativistic degrees of freedom at formation time, $M_{\text{PBH}}$ is the mass of PBHs at formation time and $f_{\text{PBH}}\equiv\Omega_{\text{PBH}}/\Omega_{\text{DM}}$ is the present-day fraction of PBH density.
On the other hand, the $\beta$ can be related to initial conditions of matter power spectrum through collapse theory of Newtonian gravity. The well-used model is the Press-Schechter one [57]. After assuming the Gaussian distribution of overdensity and simply taking the critical density contrast as the threshold $\delta_{c}=0.4$ [58], the $\beta$ can be obtained to
$\beta_{\text{PBH}}\simeq\frac{\sigma}{\sqrt{2\pi}\delta_{c}}\text{Exp}{\left(-\frac{\delta_{c}^{2}}{2\sigma^{2}}\right)},$ (4.21)
where $\sigma^{2}\ll\delta_{c}^{2}$ is the variance of the distribution. Smaller $\sigma$ implies less contrast exceeding the critical value, resulting in less $\beta_{\text{PBH}}$. The variance can be calculated from primordial power spectrum as [59]
$\sigma^{2}(R)\simeq\frac{16}{81}\int d\ln q\ W^{2}(q,R)(qR)^{4}\mathcal{P}_{\zeta}(q),$ (4.22)
where the window function $W(q,R)$ is selected to Gaussian or top-hat usually and the comoving horizon size of PBH formed is related to the mass as $R^{-1}=k_{p}$, where [60]
$M_{\text{PBH}}=30M_{\odot}\left(\frac{\gamma}{0.2}\right)\left(\frac{g_{*,f}}{10.75}\right)^{-1/6}\left(\frac{k_{p}}{2.9\times 10^{5}\text{Mpc}^{-1}}\right)^{-2}.$ (4.23)
For inflation with a triad of vector fields, where both large-scale and PBH-scale are not trivial, the enhancement at PBH scale comparing with CMB scale is
$\Delta\mathcal{P}_{\zeta}\equiv\frac{\mathcal{P}_{\zeta}(k_{p})}{\mathcal{P}_{\zeta}(k_{\text{CMB}})}=\frac{\Delta_{\zeta}(k_{p})}{\Delta_{\zeta}^{\text{(LS)}}}.$ (4.24)
Tne power spectrum at CMB scale is $\mathcal{P}_{\zeta}(k_{\text{CMB}})=2\times 10^{-9}$. For board case this is given by (4.4) and (4.7), while for sharp case, the rapid oscillation will be smoothed out by window function and we can use (4.14) and (4.19).
Figure 6: The fraction of PBHs at formation $\beta_{\text{PBH}}$ v.s. height of peaks $h$. Solid curves denote the cases of $\delta N=2$ for $h_{\text{G}}$ and $\delta N=0.4$ for $h_{\text{TH}}$, while dashed curves denote the cases of smaller $\delta N$. Orange region denotes regime that $f_{\text{PBH}}=1$ between $10^{-17}M_{\odot}<M_{\text{PBH}}<10^{-12}M_{\odot}$.
If we consider the window where the PBHs can be the candidate of dark matter: $10^{-17}M_{\odot}<M_{\text{PBH}}<10^{-12}M_{\odot}$ and $f_{\text{PBH}}=1$, from (4.20) we obtain $1.2\times 10^{-17}<\beta_{\text{PBH}}<3.7\times 10^{-15}$. Then from (4.21) the variance for PBH formation in this mass regimes is $0.0510.047<\sigma<$ 0.051. When the power spectrum is peaked at $k_{p}$, the variance is reduced to $\sigma^{2}\simeq\mathcal{P}_{\zeta}(k_{p})/5$. It’s seem that we need $\mathcal{P}_{\zeta}(k_{p})\sim 10^{-2}$ between $10^{12}\text{Mpc}^{-1}<k_{p}<10^{15}\text{Mpc}^{-1}$ to generate sufficient PBHs. However, because the profile of the peak is very sensitive to $h$ and $\delta N$ and the large-scale power spectrum also depends on these parameters. And the fraction $\beta_{\text{PBH}}$ is exponentially dependent on the power spectrum. Slightly changing profile of the peak will lead to significant change in $\beta_{\text{PBH}}$. We show how $\beta_{\text{PBH}}$ depend on parameters in Figure 6 for board Gaussian peaks and sharp top-hat peaks. We have used $R=k_{p}^{-1}$, which are the scales of peaks and can be separately obtained through (4.6) and (4.15) for Gaussian and top-hat examples. The $\beta_{\text{PBH}}$ decays rapidly for smaller parameter, i.e., the smaller height and fatness of the peaks. Also, the difference of using different window functions is also distinguishable.
The mass spectrum of PBHs are shown in Figure 7 for board Gaussian peaks and in Figure 8 for sharp top-hat peaks. We choose the parameters of these examples as following table.
Figure 7: Abundance of PBHs as dark matter for board($\delta N\gtrsim 1$) Gaussian profile of peak $h_{\text{G}}$.
Figure 8: Abundance of PBHs as dark matter for shape($\delta N\lesssim 1$) top-hat profile of peak $h_{\text{TH}}$.
($h_{\text{G}}$, $W_{\text{G}}$) ($h_{\text{G}}$, $W_{\text{TH}}$) ($h_{\text{TH}}$, $W_{\text{G}}$) ($h_{\text{TH}}$, $W_{\text{TH}}$)
Solid $h$ 7.258 5.997 24.546 19.285
$\delta N$ 2 2 0.4 0.4
Dashed $h$ 7.200 5.997 24.246 19.285
$\delta N$ 2 1.5 0.4 0.395
For solid curves, we choose the parameters so that the PBHs are responsible for total dark matter $f_{\text{PBH}}=1$ in our universe. The time of non-vanishing vector fields are chosen as $N_{f}=31.5$ for all solid examples to avoid constraints from observations of evaporation of PBHs due to Hawking radiation (EVAP) [56] and microlensing of stars in M31 by Subaru Hyper SuprimeCam (HSC) [61]. For dashed curves, we choose different parameters so that $f_{\text{PBH}}<1$ and also different $N_{f}=25.3$, which corresponds to larger PBH mass formation. We have seen in Figure 6 the abundance is sensitive to these parameter. As the first and the third column show, a mere 1% decrease in $h$ leads to a drastic drop in abundance from 1 to $10^{-5}$. For top-hat profile of peak $h_{\text{TH}}$, the abundance is also sensitive to $\delta N$ (4.14). So we can see in the last column of table, 1% decrease in $\delta N$ also reduces $f_{\text{PBH}}$ to $10^{-5}$.
We also compare the results of using Gaussian and top-hat window function. The choice of way of smoothing out the fluctuations increases the uncertainty in the results of PBH formation [62, 63]. We demonstrate a significant difference in the parameters required when using different window functions. For example, the first row in table, To generate the same abundance of PBHs, the required $h$ differs by up to 20% when using Gaussian and top-hat window functions. We usually need a larger primordial density in Gaussian smoothing. Furthermore, we also find that using different window functions results in variations in the peak position of $f_{\text{PBH}}$. The mass of formation is usually greater when using top-hat window function. In the examples, this difference reaches up to $M_{\text{peak,TH}}/M_{\text{peak,G}}\sim 10$.
5 Discussion
[…]
In addition to driving inflation, the enhancement of the power spectrum can also serve as a mechanism for the generation of primordial black holes (PBHs). This requires a strong excitation of the vector field during the late stages of inflation. We discussed two toy examples of $h$: Gaussian and top-hat. The Gaussian is employed to discuss the broad case, while the top-hat is used to address the sharp case, as both yield more precise analytical results. Due to the non-decaying entropy perturbations, the large-scale power spectrum is shifted compared to the single-field one. The amplitude of the shift depends on the details of vector fields dynamics, i.e., $\delta N$ and $h$, as shown in (4.7) and (4.19). For the shape top-hat case, this amplitude is only a function of $\delta N$, allowing for an analytical calculation of the shape of the power spectrum. Given the exponential dependence of the power spectrum shape on these parameters and the subsequent exponential dependence of PBH abundance on the power spectrum shape, the PBH abundance is highly sensitive to these parameters. Whether it’s $h$ or $\delta N$, even slight variations can significantly reduce the abundance of PBHs. Certainly, this sensitivity is also related to the specific details of the PBH collapse. For instance, we demonstrated that under the same $\delta_{c}$, smoothing the same fluctuations in different ways can lead to vastly different results, not only in PBH abundance but also in the masses of PBH formation. For example, compared to a Gaussian window function, using a top-hat window function typically results in PBHs with smaller masses. Further investigation is required to address this question comprehensively.
Now, we delve into another regime, characterized by a relatively large energy of the vector field. In this scenario, the contributions to inflationary predictions differ significantly from the case where $h$ is very small. Many questions need to be reconsidered in this regime, such as the anisotropy for one vector field and non-Gaussianity. We posit that the discussions in this paper are also applicable to a single vector field. In this case, the $IN^{2}$ term in power spectrum, which served as a contribution to anisotropy for small $h$ case, is replaced by other constant modes. It is imperative for us to calculate the anisotropy in this regime. Additionally, when considering a charged field as the inflaton, whether the perturbative Schwinger effect from the electric field is crucial to inflationary predictions is also a question that warrants consideration. Regarding PBHs, we are discussing the isotropic configuration of vector fields as an example. Although it bears similarities to hyperbolic inflation, in general, vector fields themselves exhibit phenomenology distinct from scalar fields, such as anisotropy for a single vector field and chirality for symmetry breaking. These differences manifest in the generated PBHs and induced gravitational waves. We intend to present these studies in the near future.
Acknowledgments
I would like to thank Jiro Soda for helpful discussions. C-B. C. was supported by Japanese Government (MEXT) Scholarship and China Scholarship Council (CSC).
Appendix A Derivation of the quadratic action
In this appendix we show the full quadratic action of the inflation with a triad of vector fields. We will fix the gauge freedoms of gravity as comoving gauge and then eliminate the non-dynamical degrees of freedom.
We can write down the Lagrangian of gravity, scalar field and $U(1)$ fields as following
$\mathcal{L}_{g}=N\sqrt{q}\left[\frac{M_{\text{pl}}^{2}}{2}\left({}^{(3)}R+K_{ij}K^{ij}-K^{2}\right)\right],$
$\mathcal{L}_{\phi}=\sqrt{q}\left[\frac{1}{2N}\pi^{2}-\frac{N}{2}\partial_{i}\phi\partial^{i}\phi-NV(\phi)\right],$
$\mathcal{L}_{A}=\sqrt{q}\left[\frac{f_{ab}}{2N}q^{ik}(E^{a}_{\ i}+F^{a}_{\ ij}N^{j})(E^{b}_{\ k}+F^{b}_{\ kl}N^{l})-\frac{Nf_{ab}}{4}q^{ik}q^{jl}F^{a}_{ij}F^{b}_{kl}\right],$
where we have defined $\pi=\dot{\phi}-N^{j}\phi_{|j}$ and $q_{ij}$ is the induced metric of space-like hypersurface and $E^{a}_{i}\equiv F^{a}_{0i}$. The extrinsic curvature and the spatial curvature are
$K_{ij}=\frac{1}{2N}\left(\dot{q}_{ij}-2N_{(i|j)}\right),$
${{}^{(3)}R}=\left(q_{ij,kl}+q_{mn}{{}^{(3)}\Gamma^{m}_{ij}}{{}^{(3)}\Gamma^{n}_{kl}}\right)\left(q^{ik}q^{jl}-q^{ij}q^{kl}\right)$
The quadratic action is given by
$\mathcal{L}^{(2)}=$ $\frac{a^{3}}{2}\Bigg{\{}\Big{[}-6M_{\text{pl}}^{2}H^{2}\alpha+\dot{\phi}^{2}\alpha-2V_{\phi}\delta\phi-2\dot{\phi}\delta\dot{\phi}+3f^{2}\frac{\dot{\mathbb{A}}^{2}}{a^{2}}\alpha-6ff_{\phi}\frac{\dot{\mathbb{A}}^{2}}{a^{2}}\delta\phi$
$-\sqrt{2}f\frac{\dot{\mathbb{A}}}{a}\Big{(}3\delta\dot{Q}+3(H-\frac{\dot{f}}{f})\delta Q\Big{)}+2f^{2}\frac{\dot{\mathbb{A}}}{a^{2}}\partial^{2}\mathbb{Y}\Big{]}\alpha+\frac{2}{a^{2}}\Big{(}\dot{\phi}\delta\phi-2M_{\text{pl}}^{2}HA$
$+\sqrt{2}f\frac{\dot{\mathbb{A}}}{a}\delta Q\Big{)}\partial^{2}\beta+(\delta\dot{\phi})^{2}-\frac{1}{a^{2}}(\partial_{i}\delta\phi)^{2}-V_{\phi\phi}\left(\delta\phi\right)^{2}+\frac{3}{2}\Big{[}(\delta\dot{Q})^{2}$
$+2(H-\frac{\dot{f}}{f})\delta Q\delta\dot{Q}+(H-\frac{\dot{f}}{f})^{2}(\delta Q)^{2}\Big{]}-\frac{1}{a^{2}}(\partial_{i}\delta Q)^{2}+\frac{f^{2}}{a^{2}}\partial^{2}\mathbb{Y}\partial^{2}\mathbb{Y}$
$-\sqrt{2}\frac{f}{a}\Big{(}\delta\dot{Q}+(H-\frac{\dot{f}}{f})\delta Q\Big{)}\partial^{2}\mathbb{Y}+2\sqrt{2}{a^{2}}f_{\phi}\frac{\dot{\mathbb{A}}}{a}\Big{[}3\delta\dot{Q}+3(H-\frac{\dot{f}}{f})\delta Q\Big{]}$
$-4ff_{\phi}\frac{\dot{\mathbb{A}}}{a^{2}}\partial^{2}\mathbb{Y}\delta\phi+\frac{3}{a^{2}}\big{(}f_{\phi}^{2}+ff_{\phi\phi}\big{)}\mathbb{A}^{2}(\delta\phi)^{2}$ (A.1)
$+\Big{[}\Big{(}12H\dot{\zeta}-\frac{4}{a^{2}}\partial^{2}\zeta\Big{)}M_{\text{pl}}^{2}+6f^{2}\frac{\dot{\mathbb{A}}^{2}}{a^{2}}\zeta\Big{]}A+4M_{\text{pl}}^{2}\frac{1}{a^{2}}\dot{\zeta}\partial^{2}\beta$
$-\Big{(}6\dot{\zeta}^{2}+36\dot{\zeta}\zeta+54h^{2}\zeta^{2}+\frac{2}{a^{2}}\zeta\partial^{2}\zeta\Big{)}M_{\text{pl}}^{2}+6\zeta\Big{[}\dot{\phi}-V_{\phi}\delta\phi+ff_{\phi}\frac{\dot{\mathbb{A}}^{2}}{a^{2}}\delta\phi$
$+\sqrt{2}f\frac{\dot{\mathbb{A}}}{2a}\Big{(}\delta\dot{Q}+(H-\frac{\dot{f}}{f})\delta Q\Big{)}-f^{2}\frac{\dot{\mathbb{A}}}{3a^{2}}\partial^{2}\mathbb{Y}+3\zeta^{2}\Big{(}3\dot{\phi}^{2}+5f^{2}\frac{\dot{\mathbb{A}}^{2}}{a^{2}}\Big{)}\Big{]}\Bigg{\}},$ (A.2)
where (A.1) is the parts which do not contain curvature perturbation $\zeta$ while (A.2) do. After imposing the comoving gauge (2.13), we can eliminate the inflaton perturbation $\delta\phi$. Then the equations of motion of non-dynamical degrees of freedom $\alpha$, $\beta$ and $\mathbb{Y}$ are
$\alpha=\frac{\dot{\zeta}}{H},$ $\partial^{2}\mathbb{Y}=\dot{\mathbb{A}}\left(\zeta-\frac{\dot{\zeta}}{H}\right)+\frac{a}{\sqrt{2}f}\left[\delta\dot{Q}+\left(H-\frac{\dot{f}}{f}+4f_{\phi}\frac{\dot{\mathbb{A}}}{a}h\right)\delta Q\right],$
$\frac{4}{a^{2}}M_{\text{pl}}^{2}H\partial^{2}\beta=$ $-4M_{\text{pl}}^{2}H^{2}\left(3-\epsilon_{\phi}-\frac{3}{2}\epsilon_{A}\right)\alpha-M_{\text{pl}}H^{2}\bigg{[}-12\sqrt{2}\frac{f_{\phi}}{f}M_{\text{pl}}\epsilon_{A}h$
$+2\sqrt{2}\sqrt{2\epsilon_{\phi}}(3+\eta_{\phi})h+3\sqrt{2\epsilon_{A}}\left(1-\frac{\dot{f}}{fH}\right)\bigg{]}\delta Q$
$+M_{\text{pl}}H\sqrt{2\epsilon_{A}}\left(-\delta\dot{Q}+2H\eta_{h}\delta Q+\sqrt{2}\frac{f}{a}\partial^{2}\mathbb{Y}\right)$
$+M_{\text{pl}}^{2}\left(12H\dot{\zeta}-\frac{4}{a^{2}}\partial^{2}\zeta\right)+6M_{\text{pl}}^{2}H^{2}\epsilon_{A}\zeta.$ (A.3)
After inserting these solutions into the quadratic action and imposing a series of tedious integral by parts, the quadratic action can be written as (2.2), where the mass $m_{\sigma}$, $m_{s}$ and the coupling $\mathcal{A}$, $\mathcal{B}$ are
$\ \ \ \ \ \ \ \ \mathcal{A}=4+\mathcal{O}(\epsilon),\ \ \ \ \ \ \ \mathcal{B}=\frac{16h^{2}}{1+2h^{2}}+\mathcal{O}(\epsilon),\ \ \ \ \ \ \ m_{\sigma}^{2}=\frac{16h^{2}}{1+2h^{2}}H^{2},$
$m_{s}^{2}=\frac{8h^{4}-40h^{2}-4}{1+2h^{2}}H^{2}-\frac{12h^{4}-8h^{2}-2}{1+2h^{2}}\frac{f_{\phi\phi}}{f}M_{\text{pl}}^{2}H^{2}\epsilon_{\phi}+\frac{2h^{2}}{1+2h^{2}}V_{\phi\phi}+\mathcal{O}(\epsilon).$ (A.4)
All coefficients except $m_{\sigma}$ have slow-roll corrections $\mathcal{O}(\epsilon)$. Here $\epsilon$ denotes all kinds of slow-roll parameter, i.e., $\epsilon$, $\epsilon_{\phi}$, $\epsilon_{A}$ and their higher derivative of time. We consider decoupling limit in this paper hence these corrections can be discarded. Also, we assume the slow-roll condition that $V_{\phi\phi}\ll H^{2}$ hence the third term in $m_{s}^{2}$ also can be ignored.
Appendix B Growing solution of the adiabatic mode
The negative mass square of entropy fluctuation may lead to a imaginary low frequency $\omega_{-}$. Then the curvature fluctuation start to exponentially grow at some time $\tilde{N}_{k}$, which is determined by $\omega_{-}(\tilde{N}_{k})=0$: $p^{2}(\tilde{N}_{k})=-m_{s}^{2}$. If we neglect the slow rolling, the power spectrum is scale invariant on large scales and we can rescale it as $p/H=e^{-N}$. The growing solution of the curvature fluctuation can be written as [54]
[…]
where $\omega_{-}$ is given by (3.10)
$\frac{|\omega_{-}|}{H}=\sqrt{-\mu^{2}-\frac{1}{2}u+\frac{1}{2}\sqrt{u^{2}+128h^{2}\mu^{2}}}$ (B.2)
and $\mu\equiv p/H$ and $u\equiv m_{s}^{2}/H^{2}+32h^{2}$. It’s useful to define the variable $y$ as $u^{2}y^{2}=u^{2}+128h^{2}\mu^{2}$ Then the integration in (B.1) becomes
$\mathcal{I}(N)\equiv\frac{u}{8\sqrt{2}h}\int_{y_{1}}^{y_{2}}\sqrt{b\left(y-1\right)-\left(y^{2}-1\right)}\frac{ydy}{y^{2}-1}$ (B.3)
where the parameter $b\equiv 64h^{2}/u$, $c\equiv\lambda/u$ and the limits are
$y_{1}=\sqrt{1+\frac{128h^{2}\mu^{2}}{u^{2}}},\ \ \ \ \ \ \ y_{2}=\sqrt{1+\frac{128h^{2}|m_{s}|^{2}}{u^{2}H^{2}}}.$ (B.4)
The primitive function above can be calculated and given by
$F(y)=$ $\frac{u}{8\sqrt{2}h}\Bigg{\{}\sqrt{1-b+by-y^{2}}-\frac{b}{2}\arctan\left[\frac{b-2y}{2\sqrt{1-b+by-y^{2}}}\right]$
$-\frac{\sqrt{2b}}{2}\arctan\left[\frac{2(1-y)+b(y-1-2)}{2\sqrt{2b}\sqrt{1-b+by-y^{2}}}\right],$ (B.5)
For two $\arctan x$ in $F(y)$ at $y_{2}$, the first one is negative $-\pi/2$ while the second one is positive $\pi/2$ due to $b-2y_{2}<0$ and $02(1-y_{2})+b(y_{2})-1-2c>$ 0. On the other hand, the modes will be frozen-in when crossing the horizon hence we take the upper limit at $k|c_{s}|=aH$, i.e.,
$y_{1}=\sqrt{1+\frac{128h^{2}}{u^{2}|c_{s}|^{2}}}$ (B.6)
Summarizing the above computation, we finally have
[…]
$-\frac{\sqrt{2b}}{2}\left[\frac{\pi}{2}-\arctan\left[\frac{2(1-y_{1})+b(y_{1}-3)}{2\sqrt{2b}\sqrt{1-b+by_{1}-y_{1}^{2}}}\right]\right]\Bigg{\}}.$ (B.7)
We can simple take the super-horizon limit $\mu\to 0$ and in the slow-roll limit,
$u^{2}\simeq\frac{8h^{2}(1+6h^{2})}{1+2h^{2}},\ \ \ \ \ \ \ b\simeq\frac{8(1+2h^{2})}{1+6h^{2}},\ \ \ \ \ \ \ y_{1}\simeq 1.$ (B.8)
Then the function $\mathcal{I}$ can be reduced to a simply form
$\mathcal{I}_{\text{approx}}=\sqrt{2}\pi\left(2-\sqrt{\frac{1+6h^{2}}{1+2h^{2}}}\right)h.$ (B.9)
[…]
Appendix C Bogoliubov coefficients
For the shape case, we use the top-hat profile of the $h(N)$ for analytical computations,
$h(N)=h_{0}\left[\theta(t-t_{1})-\theta(t-t_{2})\right].$ (C.1)
Before $t_{1}$, where we denote as region-I, the curvature and entropy modes are decoupled. Hence the solutions are simply the mode function with Bunch-Davies initial conditions, as the same as the single-field case,
$\hat{\zeta}_{c}^{\text{I}}(t,\boldsymbol{k})=$ $u_{k}(t)\hat{a}_{\zeta}(\boldsymbol{k})+\text{h.c.}(-\boldsymbol{k}),$
^$\hat{\mathcal{F}}^{\text{I}}(t,\boldsymbol{k})=$ $u_{k}(t)\hat{a}_{\mathcal{F}}(\boldsymbol{k})+\text{h.c.}(-\boldsymbol{k}),$ (C.2)
where $u_{k}$ is given by (4.10).
After $t>t_{1}$, which is denoted as region-II, these two scalar fields are coupled to each other hence their modes are also mixed with each other. The governing equations of motion are given by (3.1). We discuss two different non-trivial momentum regimes: near PBHs peak $k\sim k_{f}$ and large scales $k\ll k_{f}$, respectively. In order to preserve the continuity of the commutation relations, we should impose the boundary conditions at $t=t_{1}$ as
$\hat{\zeta}_{c}^{\text{I}}(t_{1}^{-})=$ $\hat{\zeta}_{c}^{\text{II}}(t_{1}^{+}),\ \ \ \ \ \ \ \dot{\hat{\zeta}}_{c}^{\text{I}}(t_{1}^{-})=D_{t}\hat{\zeta}_{c}^{\text{II}}(t_{1}^{+}),$
^$\hat{\mathcal{F}}^{\text{I}}(t_{1}^{-})=$$\hat{\mathcal{F}}^{\text{II}}(t_{1}^{+}),\ \ \ \ \ \ \ \dot{\hat{\mathcal{F}}}^{\text{I}}(t_{1}^{-})=\dot{\hat{\mathcal{F}}}^{\text{II}}(t_{1}^{+}),$ (C.3)
where $D_{t}\hat{\zeta}_{c}\equiv\dot{\hat{\zeta}}_{c}-\sqrt{2}h\mathcal{A}H\hat{\mathcal{F}}$.
During region-III $t>t_{2}$, where the curvature and entropy modes are decoupled again, the mode functions are again the ones with Bunch-David vacuum but mixing these two scalar modes,
$\hat{\zeta}_{c}^{\text{III}}(t,\boldsymbol{k})=$ $\Big{[}E_{\zeta}u_{k}(t)+F_{\zeta}u_{k}^{*}(t)\Big{]}\hat{a}_{\zeta}(\boldsymbol{k})+\Big{[}G_{\zeta}u_{k}(t)+H_{\zeta}u_{k}^{*}(t)\Big{]}\hat{a}_{\mathcal{F}}(\boldsymbol{k})$
[…]
^$\hat{\mathcal{F}}^{\text{III}}(t,\boldsymbol{k})=$ $\Big{[}E_{\mathcal{F}}u_{k}(t)+F_{\mathcal{F}}u_{k}^{*}(t)\Big{]}\hat{a}_{\zeta}(\boldsymbol{k})+\Big{[}G_{\mathcal{F}}u_{k}(t)+H_{\mathcal{F}}u_{k}^{*}(t)\Big{]}\hat{a}_{\mathcal{F}}(\boldsymbol{k})$
[…]
We then also impose the boundary conditions at $t=t_{2}$ as
$\hat{\zeta}_{c}^{\text{II}}(t_{2}^{-})=$ $\hat{\zeta}_{c}^{\text{III}}(t_{2}^{+}),\ \ \ \ \ \ \ D_{t}\hat{\zeta}_{c}^{\text{II}}(t_{2}^{-})=\dot{\hat{\zeta}}_{c}^{\text{III}}(t_{2}^{+}),$
^$\hat{\mathcal{F}}^{\text{II}}(t_{2}^{-})=$$\hat{\mathcal{F}}^{\text{III}}(t_{2}^{+}),\ \ \ \ \ \ \ \dot{\hat{\mathcal{F}}}^{\text{II}}(t_{2}^{-})=\dot{\hat{\mathcal{F}}}^{\text{III}}(t_{2}^{+}).$ (C.5)
We have eight coefficients $E_{i}$, $F_{i}$, $G_{i}$ and $H_{i}$, where $i=\zeta,\mathcal{F}$ and can be determined by the eight boundary conditions at $t_{2}$.
For modes with $k\gtrsim k_{f}$, the coefficients of curvature perturbations are solved to
$E_{\zeta}=$ $\frac{e^{i2\kappa\sinh\left(\frac{\delta N}{2}\right)}}{4\kappa\left(S_{+}^{2}-S_{-}^{2}\right)}\sum_{\pm}\pm\frac{S_{\mp}^{2}-\kappa^{2}}{S_{\pm}}\Bigg{\{}e^{iS_{\pm}\delta N}\left[1+\left(S_{\pm}-\kappa\right)^{2}\right]-e^{-iS_{\pm}\delta N}\left[1+\left(S_{\pm}+\kappa\right)^{2}\right]\Bigg{\}},$
$F_{\zeta}=$ $\frac{e^{i2\kappa\cosh\left(\frac{\delta N}{2}\right)}}{2\kappa\left(S_{+}^{2}-S_{-}^{2}\right)}\sum_{\pm}\pm\frac{i\sin{\left(S_{\pm}\delta N\right)}}{S_{\pm}}\left(S_{\mp}^{2}-\kappa^{2}\right)\left[S_{\pm}^{2}+\left(1-\kappa^{2}-i2\kappa\right)\right],$
$G_{\zeta}=$ $\frac{i\sqrt{2}h\mathcal{A}e^{i2\kappa\sinh\left(\frac{\delta N}{2}\right)}}{4\kappa\left(S_{+}^{2}-S_{-}^{2}\right)}\sum_{\pm}\pm\frac{1}{S_{\pm}}\Bigg{\{}e^{iS_{\pm}\delta N}\left[\kappa^{2}-S_{\pm}(i+\kappa)\right]\left(i-\kappa+S_{\pm}\right)$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ -e^{-iS_{\pm}\delta N}\left[\kappa^{2}+S_{\pm}\left(i+\kappa\right)\right]\left(i-\kappa-S_{\pm}\right)\Bigg{\}},$
$H_{\zeta}=$ $\frac{i\sqrt{2}h\mathcal{A}e^{i2\kappa\cosh\left(\frac{\delta N}{2}\right)}}{4\kappa\left(S_{+}^{2}-S_{-}^{2}\right)}\sum_{\pm}\pm\frac{1}{S_{\pm}}\Bigg{\{}e^{iS_{\pm}\delta N}\left[\kappa^{2}-S_{\pm}(i+\kappa)\right]\left(i+\kappa+S_{\pm}\right)$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ -e^{-iS_{\pm}\delta N}\left[\kappa^{2}+S_{\pm}\left(i+\kappa\right)\right]\left(i+\kappa-S_{\pm}\right)\Bigg{\}},$ (C.6)
where $\kappa\equiv k/k_{f}$, $\delta N\equiv H(t_{2}-t_{1})$ and $S_{\pm}\equiv\omega_{\pm}/H$.
[…]
$E_{\zeta}=$ $\frac{1}{27}\left[32\cosh{\left(3\delta N\right)}-5\right]+i\frac{4\left(1-e^{-3\delta N}\right)}{9x_{2}^{3}}+i\frac{e^{-\delta N/2}}{3x_{2}}\bigg{[}\sinh\left(\frac{\delta N}{2}\right)$
$+4\sinh\left(\frac{5\delta N}{2}\right)\bigg{]}+\cdots,$
$F_{\zeta}=$ $\frac{32}{27}\sinh\left(3\delta N\right)+i\frac{4\left(1-e^{-3\delta N}\right)}{9x_{2}^{3}}+i\frac{e^{-\delta N/2}}{3x_{2}}\left[\sinh\left(\frac{\delta N}{2}\right)+4\sinh\left(\frac{5\delta N}{2}\right)\right]+\cdots,$
$G_{\zeta}=$ $\frac{e^{-3\delta N/2}}{3\sqrt{3}}\left(2e^{3\delta N}+1-i\frac{6}{x_{2}^{3}}-ie^{2\delta N}\frac{3}{x_{2}}\right)\sin{\left(2\sqrt{6}h\delta N\right)}+\cdots,$
$H_{\zeta}=$ $\frac{e^{-3\delta N/2}}{3\sqrt{3}}\left(2e^{3\delta N}-1-i\frac{6}{x_{2}^{3}}-ie^{2\delta N}\frac{3}{x_{2}}\right)\sin{\left(2\sqrt{6}h\delta N\right)}+\cdots.$ (C.7)
We find that although the results are complex and the imaginary parts are the functions of $x_{2}$, in $(E_{\zeta}-F_{\zeta})$ and $(G_{\zeta}-H_{\zeta})$, the imaginary parts will be killed to each other. Because perturbations $\zeta$ becomes classical and is a constant after horizon-crossing.

### v2

@@
 We revisit inflation coupled with vector fields employing kinetic coupling in the comoving gauge.
-It is known that there is a cumulative effect $IN^{2}$ on curvature power spectrum.
-For a large number of e-foldings $N$, this contribution is so significant that it could violate observational constraints when the ratio of kinetic energy between vector fields and inflaton $I$ is not extremely small.
-In this paper, we explore the regime where $I\gg 1$, a realm that has not been extensively explored due to the limitations of perturbative methods.
+It is known that there is a cumulative effect $\tilde{I}N^{2}$ on curvature power spectrum.
+For a large number of e-foldings $N$, this contribution is so significant that it could violate observational constraints when the ratio of kinetic energy between vector fields and inflaton $\tilde{I}$ is not extremely small.
+In this paper, we explore the regime where $\tilde{I}\gg 1$, a realm that has not been extensively explored due to the limitations of perturbative methods.
 We found that the entropy perturbation becomes heavy in this regime and the cumulative effect decays away on super-horizon scales.
@@
 The inflationary phase undergoes a transition to new attractors if the potential is steep enough [20], which is very similar to the hyperbolic inflation [8].
-This kind of models was firstly constructed to explain the statistical anisotropy in the CMB [21, 22, 23, 24].
+This kind of models was firstly constructed to explain the statistical anisotropy in the CMB [21, 22, 23, 24] (or read the review [25]).
 The vector field is coupled to gravity through $f^{2}(\phi)F_{\mu\nu}F^{\mu\nu}$.
@@
 Under this new attractor, the power spectrum of curvature perturbation is sourced by the vector field, yielding an additional contribution
-$\mathcal{P}_{\zeta}=\mathcal{P}_{\zeta}^{(0)}\left[1+\Theta(\theta)IN^{2}\right],$ (1.1)
-where $N\equiv\ln{(-k\tau_{e})}$ represents the e-folding number of observed mode elapsed since horizon-crossing and $I=2\epsilon(\rho_{A}/V)$ is the fraction kinetic energy of the vector field.
+$\mathcal{P}_{\zeta}=\mathcal{P}_{\zeta}^{(0)}\left[1+\Theta(\theta)IN_{k}^{2}\right],\ \ \ \ \ \ (I\ll 1)$ (1.1)
+where $N_{k}\equiv\ln{(-k\tau_{e})}$ represents the e-folding number of observed mode elapsed since horizon-crossing, $\tau_{e}$ is time of end of the inflation.
+$I$ is the parameter related to specific models.
 For one U(1) vector field, $\Theta(\theta)=24\sin^{2}\theta$, where $\theta$ is the angle between mode and direction of the vector field [22, 26].
-The term $IN^{2}$ accountable for the amplitude of statistical anisotropy and should be rigorously constrained by CMB observations to $I<10^{-7}$ for a sufficiently lengthy inflationary period $N\sim 60$ [22, 25].The exceedingly small parameter implies the presence of a very small but non-zero vector field, potentially giving rise to fine-tuning challenges [26, 27].
+The term $IN_{k}^{2}$ accountable for the amplitude of statistical anisotropy and should be rigorously constrained by CMB observations to $I<10^{-7}$ for a sufficiently lengthy inflationary period $N_{k}\sim 60$ [22, 26].The exceedingly small parameter implies the presence of a very small but non-zero vector field, potentially giving rise to fine-tuning challenges [27, 28].
 Studies have revealed that employing multiple vector fields can mitigate the anisotropy, eventually converging to an isotropic attractor [24].
@@
 This is a cumulative effect, i.e., the curvature is sourced by non-zero vector fields and evolves on super-horizon scales.
-However, this does not imply that results for large $I$ are undesirable.
+However, this does not imply that results for $I\sim\mathcal{O}(1)$ are undesirable.
 If we require vector fields to handle a sufficiently large field displacement, indicating that $I$ is not a small parameter, it appears that the second term in (1.1) becomes significant and violates the adiabatic and scale-invariant initial conditions.
@@
 Numerical findings indicate that the power spectrum does not actually evolve on super-horizon scales [31, 20].
-In this paper, we revisit this dilaton-gauge coupling model in comoving gauge and thoroughly explore the $I\gg 1$ case with the assistance of effective field theory (EFT).
+In this paper, we revisit this dilaton-gauge coupling model in comoving gauge and thoroughly explore the significant $I$ case with the assistance of effective field theory (EFT).
 We will show, the entropy perturbation is super heavy in this regime and the theory can be described by a low-energy EFT of a massless adiabatic perturbation.
-$IN^{2}$ does not occur in this regime; instead, it is replaced by an exponentially enhanced power spectrum, dependent solely on the parameter $I$.
+$IN_{k}^{2}$ does not occur in this regime; instead, it is replaced by an exponentially enhanced power spectrum, dependent solely on one parameter.
 This enhancement also provides a mechanism for generating primordial black holes.
 2 Inflation with vector fields
-We are considering the spatially flat Friedmann-Lemaitre-Robertson-Walker (FLRW) universe, which is characterized by the scale factor $a(t)$ and the Hubble variable is $H(t)=\dot{a}/a$.
+We are considering the spatially flat Friedmann-Lemaitre-Robertson-Walker (FLRW) universe, which is characterized by the scale factor $a(t)$ and the Hubble variable is $H(t)=\dot{a}/a$, where ˙$\dot{}\equiv d/dt$ is cosmological time derivative.
 The Arnowitt-Deser-Misner(ADM) formalism of the metric is [32]
 $ds^{2}=-N^{2}dt^{2}+q_{ij}\left(N^{i}dt+dx^{i}\right)\left(N^{j}dt+dx^{j}\right),$ (2.1)
-where $N$ is the lapse function, $N^{i}$ is the shift vector and $g_{ij}$ is the metric of hypersurface.
+where $N$ is the lapse function and we choose the cosmological time $N=1$ for background, $N^{i}$ is the shift vector and $g_{ij}$ is the metric of hypersurface.
 We consider the model of inflaton and $U(1)$ gauge fields, which is called dilaton-gauge inflation in this paper
-$S=\int d^{4}x\sqrt{-{g}}\left[-\frac{1}{2}\partial^{\mu}\phi\partial_{\mu}\phi-\frac{1}{4}f_{ab}(\phi)F^{a}_{\ \mu\nu}F^{b\mu\nu}\right],$ (2.2)
-where $f_{ab}$ is the kinetic coupling and only depends on the inflaton.
+$S=\int d^{4}x\sqrt{-{g}}\left[-\frac{1}{2}\partial^{\mu}\phi\partial_{\mu}\phi-V(\phi)-\frac{1}{4}f_{ab}(\phi)F^{a}_{\ \mu\nu}F^{b\mu\nu}\right],$ (2.2)
+where $g$ is the determinant of metric, $V(\phi)$ is the inflationary potential and $f_{ab}$ is the kinetic coupling and only depends on the inflaton.
+The strength of vector is $F^{a}_{\ \mu\nu}=\partial_{\mu}A^{a}_{\ \nu}-\partial_{\nu}A^{a}_{\ \mu}$.
 2.1 Background
@@
 The equations of motion of the matter fields are
-$\ddot{\phi}+3H\dot{\phi}+V_{\phi}-3f_{\phi}f\frac{\dot{\mathbb{A}}^{2}}{a^{2}}=0,\ \ \ \ \ \ \ \ddot{\mathbb{A}}+\left(2\frac{\dot{f}}{f}+H\right)\dot{\mathbb{A}}=0.$ (2.6)
+$\ddot{\phi}+3H\dot{\phi}+V_{\phi}-3f_{\phi}f\frac{\dot{\mathbb{A}}^{2}}{a^{2}}=0,\ \ \ \ \ \ \ \ddot{\mathbb{A}}+\left(2\frac{\dot{f}}{f}+H\right)\dot{\mathbb{A}}=0,$ (2.6)
+where $V_{\phi}\equiv\partial V/\partial\phi$ and $f_{\phi}\equiv\partial f/\partial\phi$.
 In this paper we assume that $\dot{\phi}<0$ and $\dot{\mathbb{A}}>0$.
-We also define a useful variable, the ratio of the kinetic energy of $U(1)$ fields and dilaton [^1] The parameter $I$ in [28] is related to $h$ in our paper as $I\simeq 2h^{2}$.
+We also define a useful variable, the ratio of the kinetic energy of $U(1)$ fields and dilaton [^1] The parameter $I$ in [29] or in (1.1) is related to $h$ in our paper as $I\simeq 2h^{2}$ for $h\ll 1$.
 $h\equiv\sqrt{\frac{\epsilon_{A}}{2\epsilon_{\phi}}}.$ (2.7)
@@
 Taking time derivative on the Friedmann equation one can obtain
-$V_{\phi}=M_{\text{pl}}H^{2}\left(6\epsilon+2\epsilon_{\phi}\eta_{\phi}+\frac{3}{2}\epsilon_{A}\eta_{A}-3\epsilon\epsilon_{A}\right).$ (2.8)
+$\sqrt{2\epsilon_{\phi}}V_{\phi}=M_{\text{pl}}H^{2}\left(6\epsilon+2\epsilon_{\phi}\eta_{\phi}+\frac{3}{2}\epsilon_{A}\eta_{A}-3\epsilon\epsilon_{A}\right).$ (2.8)
 Inserting $V_{\phi}$ into the equation of motion of $\phi$ we obtain
@@
 The scalar fluctuations of gravity and scalar fields can be written as
-$N=1+\alpha,\ \ \ \ N_{i}=\partial_{i}\beta,\ \ \ \ g_{ij}=a^{2}(t)e^{2\zeta}\delta_{ij},\ \ \ \ \phi=\phi(t)+\delta\phi.$ (2.11)
+$N=1+\alpha,\ \ \ \ N_{i}=\partial_{i}\beta,\ \ \ \ g_{ij}=a^{2}(t)\left(e^{2\zeta}\delta_{ij}+2\partial_{i}\partial_{j}E\right),\ \ \ \ \phi=\phi(t)+\delta\phi.$ (2.11)
 There are four gauge freedoms of the time and spatial coordinate choice.
@@
 We choose the comoving gauge
-$Q_{\sigma}=0.$ (2.13)
+$Q_{\sigma}=0,\ \ \ \ \ \ E=0.$ (2.13)
 This is the unitary gauge in the gauge theory, whose Nambu-Goldstone boson is eaten by the metric, as we will see in the next section.
@@
 $m_{s}^{2}=\frac{8h^{4}-40h^{2}-4}{1+2h^{2}}H^{2}-\frac{12h^{4}-8h^{2}-2}{1+2h^{2}}\frac{f_{\phi\phi}}{f}M_{\text{pl}}^{2}H^{2}\epsilon_{\phi},$ (2.19)
-where we have discard slow-roll suppressed terms in the slow-low limit and also assumed $|V_{\phi\phi}|/H^{2}\ll 1$.
+where we have discarded slow-roll suppressed terms in the slow-low limit and also assumed $|V_{\phi\phi}|/H^{2}\ll 1$.
 This is the decoupling limit $\epsilon\to 0$ in the EFT.
@@
 $\mathcal{F}_{0}=-\frac{2\dot{\sigma}h\mathcal{B}H}{m_{s}^{2}}\zeta_{0}.$ (2.23)
-These constant solutions exist for any value of $h$. And it’s obvious that these mode solutions are not the correction $IN^{2}$ in (1.1). We next explore the Lagrangian for small $h$.
+These constant solutions exist for any value of $h$. And it’s obvious that these mode solutions are not the correction $IN_{k}^{2}$ in (1.1). We next explore the Lagrangian for small $h$.
 […]
@@
 […]
-where $N_{c}$ is the time of horizon-crossing and $N_{k}\equiv N_{e}-N_{c}$ is the e-folding number elapsed since horizon-crossing.
+where $N_{c}$ is the time of horizon-crossing, $N_{e}$ is the time of end of inflation and $N_{k}\equiv N_{e}-N_{c}$ is the e-folding number elapsed since horizon-crossing.
 For small $h$ this is a small correction to the adiabatic mode.
@@
 $\delta\mathcal{P}_{\zeta}\simeq 32h^{2}N^{2}\cdot\frac{H^{2}}{8\pi^{2}\epsilon M_{\text{pl}}^{2}},\ \ \ \ \ (h\ll 1)$ (2.28)
-which is consistent with the result calculated through in-in formalism [28].
-These correction is proportional to square of e-folding number $N\sim 60$, which is a large number for inflation.
+which is consistent with the result calculated through in-in formalism [29] and reproduces the result (1.1).
+These correction is proportional to square of e-folding number $N_{k}\sim 60$, which is a large number for inflation.
 Hence for anisotropic configuration of vector fields, there is very strong constrain from CMB on the parameter $h$.
@@
 3 EFT under massive entropy mode
-In the EFT formalism, the physical scalar d.o.f at low-energy scale is the Nambu-Goldstone (NG) boson particle coming from breaking the time translation symmetry $t\to t+\xi^{0}$ of the de Sitter space by a clock field $\phi(t)$, i.e., the homogeneous inflaton.
-One can choose the slice of $t$ that the inflaton has no excitation all the time, i.e.
-the unitary gauge.
-In this gauge, the only dynamical scalar d.o.f at low-energy is the scalar part of graviton $g_{ij}=a^{2}(t)e^{2\zeta}\delta_{ij}$, where the scalar $\zeta(t,\boldsymbol{x})$ is the spatial curvature of the time slice.
-The EFT of a single scalar field in the unitary gauge can be written as [38]
-$\mathcal{L}_{\text{EFT}}=\sqrt{-g}\left[\frac{M_{\text{pl}}^{2}}{2}R+M_{\text{pl}}^{2}\dot{H}g^{00}-M_{\text{pl}}^{2}\left(3H^{2}+\dot{H}\right)+\cdots\right]+\mathcal{L}_{\text{n-u}},$ (3.1)
-where the dots stand for the higher-order terms of $(1+g^{00})$ and we have ignored the extrinsic curvature fluctuation $\delta K_{\mu\nu}$.
-$\mathcal{L}_{\text{n-u}}$ is the non-universal parts and for $\omega^{2}\ll M^{2}+p^{2}$ [39]
-$\mathcal{L}_{\text{n-u}}=\sqrt{-g}\sum_{n=2}\frac{M_{n}^{4}}{n!}\left[(1+g^{00})\frac{M^{2}}{M^{2}-\tilde{\partial}^{2}}\right]^{n-1}(1+g^{00})+\cdots,$ (3.2)
-where $M$ is the mass of heavy fields and ~$\tilde{\partial}\equiv\partial/a$.
-These operators are non-local generally due to the hierarchy between the frequency and momentum when the speed of sound is small $c_{s}\ll 1$.
-There exist regime in energy scales that the dispersion is non-linear [39, 40].
-If $\omega\sim p$ the momentum is also $p\ll M$ and we can ignore the non-local operators.
-In order to transform to the spatial flat gauge $g_{ij}=a^{2}(t)\delta_{ij}$, one can perform a time diffeomorphisms $t\to\tilde{t}=t-\pi(t,\boldsymbol{x})$.
-Then at linear-order the NG boson is related to the curvature fluctuation as $\zeta=-H\pi$.
-The time component of the metric transforms as
-$g^{00}=\frac{\partial x^{0}}{\partial\tilde{x}^{\mu}}\frac{\partial x^{0}}{\partial\tilde{x}^{\nu}}\tilde{g}^{\mu\nu}=\left(1+\dot{\pi}\right)^{2}\tilde{g}^{00}+2\left(1+\dot{\pi}\right)\partial_{i}\pi\tilde{g}^{0i}+\tilde{g}^{ij}\partial_{i}\pi\partial_{j}\pi$ (3.3)
-In the low-energy scale $\omega\sim p\ll M$, we can write down the action in the re-parameterization of time. Dropping the tildes for simplicity the Lagrangian is given by
-$\mathcal{L}_{\text{EFT}}=$ $\sqrt{-g}\Bigg{\{}\frac{M_{\text{pl}}^{2}}{2}R-M_{\text{pl}}^{2}\left[3H^{2}(t+\pi)+\dot{H}(t+\pi)\right]$
-$+M_{\text{pl}}^{2}\dot{H}(t+\pi)\left[\left(1+\dot{\pi}\right)^{2}g^{00}+2\left(1+\dot{\pi}\right)\partial_{i}\pi g^{0i}+\partial^{i}\pi\partial_{i}\pi\right]+$
-$+\frac{M_{2}^{4}}{2}\left[1+(1+\dot{\pi})^{2}g^{00}+2(1+\dot{\pi})\partial_{i}\pi g^{0i}+\partial^{i}\pi\partial_{i}\pi)\right]^{2}+\cdots\Bigg{\}}$ (3.4)
-Expanding the action up to second order of fluctuations, the constrain of the lapse function is solved as $\alpha=\epsilon H\pi$ [41].
-In the dilaton-gauge theory, from the quadratic action in the spatial flat gauge (A), constrain $\alpha$ is given by $\alpha\propto Q_{\sigma}$, which means that the unitary gauge is the vanishing adiabatic fluctuation at linear order.
-In the decoupling limit, the NG boson is decoupled with gravity and becomes a dynamical scalar mode.
-For the inflation, this decoupling limit is the slow-roll limit $\epsilon\to 0$.
-Then the quadratic Lagrangian of NG boson reduces to
-$\mathcal{L}_{\text{EFT}}=a^{3}M_{\text{pl}}^{2}H^{2}\frac{\epsilon}{c_{s}^{2}}\left[\dot{\pi}^{2}-\frac{c_{s}^{2}}{a^{2}}(\partial\mathcal{\pi})^{2}\right],$ (3.5)
-where the speed of sound is defined by
-$\frac{1}{c_{s}^{2}}\equiv 1-\frac{2M_{2}^{4}}{M_{\text{pl}}^{2}\dot{H}}.$ (3.6)
-The modification of the speed of sound is provided by the higher operator $M_{2}$ and we will show soon that this can be obtained by integrating out the heavy fluctuation at low energy in the dilaton-gauge model.
 3.1 Modified dispersion relation
@@
 The heavy entropy fluctuation implies the hierarchy of the two frequencies $\omega_{-}\ll\omega_{+}$, which requires $4(1-c_{s}^{2})p^{2}\ll m_{s}^{2}c_{s}^{-2}$.
-Hence we have a cut-off scale of EFT (3.5) on momentum $p_{\text{UV}}$ and the EFT is only valid below this scale.
+Hence we have a cut-off scale of EFT (B.8) on momentum $p_{\text{UV}}$ and the EFT is only valid below this scale.
 The UV scale is given by
@@
 Figure 1: Evolution of entropy perturbation $\mathcal{F}$ for $h=10$. An instability before horizon-crossing and constant evolution on super-horizon are presented. The scales are arbitrary.
-We have noted that the speed of sound is $|c_{s}|\sim\mathcal{O}(1)$ for large $h$ hence the non-linear dispersion regime is narrower, and we can directly investigate the linear dispersion regime.
+We have noted that the speed of sound is $|c_{s}|\sim\mathcal{O}(1)$ for large $h$ hence the non-linear dispersion regime is narrower, and we can directly investigate the linear dispersion regime (see Appendix B for reviews).
 In the heavy limit $|m_{s}^{2}|\gg\omega^{2}$ we can expand the operator
@@
 And these PBHs can be one of candidates of the dark matter [50, 51, 52] and supermassive black holes in our universe [53].
+The enhancement of the power spectrum due to the heavy fields is studied recently in multiple scalar-field models, and is used to discuss the PBHs formation [56, 54, 55].
+In previous section, we have realized that it also can be achieved by using vector fields.
 We approach to the PBHs without any specific models ($V(\phi)$ and $f(\phi)$).
@@
 4.1 Board case
-For this case, we have enough e-folding of $h(N)$ for exponential growth of the modes. The EFT is valid at time $N_{k}$ satisfying
+For this case, we have enough e-folding of $h(N)$ for exponential growth of the modes. The EFT is valid at time $N_{v}$ satisfying
 $\frac{k}{a(N_{v})}=\frac{3}{4}|m_{s}(N_{v})|.$ (4.2)
@@
 Taking the super-horizon limit, the leading contribution of the power spectrum is
-$\Delta_{\zeta}(k)\equiv\frac{\mathcal{P}_{\zeta}(k)}{\mathcal{P}_{\zeta}^{(0)}(k)}\simeq e^{2\mathcal{I}(h)}\big{|}_{N_{k}},$ (4.4)
-where the exponential factor is given by (see Appendix B for derivation)
+$\Delta_{\zeta}(k)\equiv\frac{\mathcal{P}_{\zeta}(k)}{\mathcal{P}_{\zeta}^{(0)}(k)}\simeq e^{2\mathcal{I}(h)}\big{|}_{N_{v}},$ (4.4)
+where the exponential factor is given by (see Appendix C for derivation [^2] Although this analytical result fixes well with the numerical one.
+This derivation ignores the Hubble friction in the equations of motion.
+After completing this paper, [58] proposed a rigorous derivation when take into account the Hubble fraction terms.)
 $\mathcal{I}(h)\simeq\sqrt{2}\pi\left(2-\sqrt{\frac{1+6h^{2}}{1+2h^{2}}}\right)h.$ (4.5)
@@
 We also discuss large-scale modes $k\ll k_{f}$ and small-scale modes $k\sim k_{f}$ separately.
-Figure 4: The numerical and analytical results (C) of power spectrum for the top-hat case, where $\delta N=0.4$ and $h_{0}=24$. The envelope (4.14) is shown as orange dashed curve.
+Figure 4: The numerical and analytical results (D) of power spectrum for the top-hat case, where $\delta N=0.4$ and $h_{0}=24$. The envelope (4.14) is shown as orange dashed curve.
 Modes with $\boldsymbol{k\gtrsim k_{f}}$
@@
 We have eight unknown coefficients $A_{\pm}^{\zeta}$, $B_{\pm}^{\zeta}$, $C_{\pm}^{\zeta}$ and $D_{\pm}^{\zeta}$ in the region-II and we have eight boundary conditions at $t_{1}$ hence all these coefficients can be completely solved.
-Then at $t_{2}$ we also have eight boundary condictions and the coefficients of (C) can be solved to (see Appendiex C)
+Then at $t_{2}$ we also have eight boundary condictions and the coefficients of (4.2) can be solved to (see Appendiex D)
 $E_{\zeta}$ $\simeq\frac{\left(S_{-}+\kappa\right)^{2}}{4\kappa S_{-}}e^{i2\kappa\sinh(\delta N/2)-iS_{-}\delta N},$
@@
 where $\kappa\equiv k/k_{f}$, $\delta N\equiv H(t_{2}-t_{1})$ and $S_{\pm}\equiv\omega_{\pm}/H$.
-We show the agreement between numerical and completed analytical results (see appendix C) in Figure 4.
+We show the agreement between numerical and completed analytical results (see appendix D) in Figure 4.
 The power spectrum has exponential enhancement near $k_{f}$ and also rapidly oscillate during these scales.
@@
 We have eight unknown coefficients $c_{0}^{i}$, $c_{1}^{i}$, $c_{+}^{i}$ and $c_{-}^{i}$ ($i=\zeta,\mathcal{F}$) in the region-II and eight boundary conditions at $t_{1}$ hence all these coefficients can be completely solved.
-Then at $t_{2}$ the coefficients of (C) can also be solved [^2] We found that this computation breaks when $x\lesssim 10^{-5}$..
-The results are extremely tedious, but we can simple them under some approximations, see appendix C.
+Then at $t_{2}$ the coefficients of (4.2) can also be solved [^3] We found that this computation breaks when $x\lesssim 10^{-5}$..
+The results are extremely tedious, but we can simple them under some approximations, see appendix D.
 The results should be real because the perturbations become classical on super-horizon.
@@
 These differences manifest in the generated PBHs and induced gravitational waves.
+Recently observations from 15-year NANOGrav data set reveals information about low-frequency stochastic gravitational wave background [67].
+PBHs can be one of the sources of this background [68].
+The formation of PBHs in single-field up to one-loop correction is still controversial [69, 70, 71, 72, 73, 74, 75, 76].
+The single-field EFT with a imaginary $c_{s}$ should also be discussed.
 We intend to present these studies in the near future.
@@
 Also, we assume the slow-roll condition that $V_{\phi\phi}\ll H^{2}$ hence the third term in $m_{s}^{2}$ also can be ignored.
-Appendix B Growing solution of the adiabatic mode
+Appendix B Effective field theory of inflation
+In the EFT formalism, the physical scalar d.o.f at low-energy scale is the Nambu-Goldstone (NG) boson particle coming from breaking the time translation symmetry $t\to t+\xi^{0}$ of the de Sitter space by a clock field $\phi(t)$, i.e., the homogeneous inflaton.
+One can choose the slice of $t$ that the inflaton has no excitation all the time, i.e.
+the unitary gauge.
+In this gauge, the only dynamical scalar d.o.f at low-energy is the scalar part of graviton $g_{ij}=a^{2}(t)e^{2\zeta}\delta_{ij}$, where the scalar $\zeta(t,\boldsymbol{x})$ is the spatial curvature of the time slice.
+The EFT of a single scalar field in the unitary gauge can be written as [39]
+$\mathcal{L}_{\text{EFT}}=\sqrt{-g}\left[\frac{M_{\text{pl}}^{2}}{2}R+M_{\text{pl}}^{2}\dot{H}g^{00}-M_{\text{pl}}^{2}\left(3H^{2}+\dot{H}\right)+\cdots\right]+\mathcal{L}_{\text{n-u}},$ (B.1)
+where the dots stand for the higher-order terms of $(1+g^{00})$ and we have ignored the extrinsic curvature fluctuation $\delta K_{\mu\nu}$.
+$\mathcal{L}_{\text{n-u}}$ is the non-universal parts and for $\omega^{2}\ll M^{2}+p^{2}$ [40]
+$\mathcal{L}_{\text{n-u}}=\sqrt{-g}\sum_{n=2}\frac{M_{n}^{4}}{n!}\left[(1+g^{00})\frac{M^{2}}{M^{2}-\tilde{\partial}^{2}}\right]^{n-1}(1+g^{00})+\cdots,$ (B.2)
+where $M$ is the mass of heavy fields and ~$\tilde{\partial}\equiv\partial/a$.
+These operators are non-local generally due to the hierarchy between the frequency and momentum when the speed of sound is small $c_{s}\ll 1$.
+There exist regime in energy scales that the dispersion is non-linear [40, 41].
+If $\omega\sim p$ the momentum is also $p\ll M$ and we can ignore the non-local operators.
+In order to transform to the spatial flat gauge $g_{ij}=a^{2}(t)\delta_{ij}$, one can perform a time diffeomorphisms $t\to\tilde{t}=t-\pi(t,\boldsymbol{x})$.
+Then at linear-order the NG boson is related to the curvature fluctuation as $\zeta=-H\pi$.
+The time component of the metric transforms as
+$g^{00}=\frac{\partial x^{0}}{\partial\tilde{x}^{\mu}}\frac{\partial x^{0}}{\partial\tilde{x}^{\nu}}\tilde{g}^{\mu\nu}=\left(1+\dot{\pi}\right)^{2}\tilde{g}^{00}+2\left(1+\dot{\pi}\right)\partial_{i}\pi\tilde{g}^{0i}+\tilde{g}^{ij}\partial_{i}\pi\partial_{j}\pi$ (B.3)
+In the low-energy scale $\omega^{2}\ll p^{2}+M^{2}$, we can write down the action in the re-parameterization of time. Dropping the tildes for simplicity, the Lagrangian of NG boson is given by [40]
+$\mathcal{L}_{\pi}=$ $\sqrt{-g}\Bigg{\{}-M_{\text{pl}}^{2}\dot{H}\left[\dot{\pi}\left(1+\frac{2M_{2}^{4}}{M_{\text{pl}}^{2}|\dot{H}|}\frac{M^{2}}{M^{2}-\tilde{\partial}^{2}}\right)\dot{\pi}-(\tilde{\partial}\pi)^{2}\right]$
+$+2M_{2}^{4}\left[\dot{\pi}^{2}-(\tilde{\partial}\pi)^{2}\right]\frac{M^{2}}{M^{2}-\tilde{\partial}^{2}}\dot{\pi}-\frac{4}{3}M_{3}^{4}\left(\dot{\pi}\frac{M^{2}}{M^{2}-\tilde{\partial}^{2}}\right)^{2}\dot{\pi}+\cdots\Bigg{\}},$ (B.4)
+where we have expanding the action up to third order of fluctuations.
+In the decoupling limit, the NG boson is decoupled with gravity and becomes a dynamical scalar mode.
+For the inflation, this decoupling limit is the slow-roll limit $\epsilon\to 0$.
+Then the quadratic Lagrangian of NG boson reduces to
+$\mathcal{L}_{\pi}=a^{3}M_{\text{pl}}^{2}\epsilon\left[\dot{\pi}\left(\frac{M^{2}c_{s}^{-2}-\tilde{\partial}^{2}}{M^{2}-\tilde{\partial}^{2}}\right)\dot{\pi}-(\tilde{\partial}\mathcal{\pi})^{2}\right],$ (B.5)
+where the speed of sound is defined by
+$\frac{1}{c_{s}^{2}}\equiv 1-\frac{2M_{2}^{4}}{M_{\text{pl}}^{2}\dot{H}}.$ (B.6)
+The modification of the speed of sound is provided by the higher operator $M_{2}$ and we will show soon that this can be obtained by integrating out the heavy fluctuation at low energy in the dilaton-gauge model.
+The dispersion relation of the propagation of the light field $\pi$ should be written as
+$\omega^{2}(p)=\frac{M^{2}+p^{2}}{M^{2}c_{s}^{-2}+p^{2}}p^{2},$ (B.7)
+which is generally non-linear. However, if the speed of sound $c_{s}$ is not very small, The region with nonlinear dispersion relations is very narrow so the Lagrangian of NG boson shortly reduces to
+$\mathcal{L}_{\pi}=a^{3}M_{\text{pl}}^{2}H^{2}\frac{\epsilon}{c_{s}^{2}}\left[\dot{\pi}^{2}-\frac{c_{s}^{2}}{a^{2}}(\partial\mathcal{\pi})^{2}\right],$ (B.8)
+where the dispersion relation is almost linear $\omega(p)\simeq c_{s}p$ throughout the entire process.
+Appendix C Growing solution of the adiabatic mode
 The negative mass square of entropy fluctuation may lead to a imaginary low frequency $\omega_{-}$.
@@
 where $\omega_{-}$ is given by (3.4)
-$\frac{|\omega_{-}|}{H}=\sqrt{-\mu^{2}-\frac{1}{2}u+\frac{1}{2}\sqrt{u^{2}+128h^{2}\mu^{2}}}$ (B.2)
-and $\mu\equiv p/H$ and $u\equiv m_{s}^{2}/H^{2}+32h^{2}$. It’s useful to define the variable $y$ as $u^{2}y^{2}=u^{2}+128h^{2}\mu^{2}$ Then the integration in (B.1) becomes
-$\mathcal{I}(N)\equiv\frac{u}{8\sqrt{2}h}\int_{y_{1}}^{y_{2}}\sqrt{b\left(y-1\right)-\left(y^{2}-1\right)}\frac{ydy}{y^{2}-1}$ (B.3)
+$\frac{|\omega_{-}|}{H}=\sqrt{-\mu^{2}-\frac{1}{2}u+\frac{1}{2}\sqrt{u^{2}+128h^{2}\mu^{2}}}$ (C.2)
+and $\mu\equiv p/H$ and $u\equiv m_{s}^{2}/H^{2}+32h^{2}$. It’s useful to define the variable $y$ as $u^{2}y^{2}=u^{2}+128h^{2}\mu^{2}$ Then the integration in (C.1) becomes
+$\mathcal{I}(N)\equiv\frac{u}{8\sqrt{2}h}\int_{y_{1}}^{y_{2}}\sqrt{b\left(y-1\right)-\left(y^{2}-1\right)}\frac{ydy}{y^{2}-1}$ (C.3)
 where the parameter $b\equiv 64h^{2}/u$, $c\equiv\lambda/u$ and the limits are
-$y_{1}=\sqrt{1+\frac{128h^{2}\mu^{2}}{u^{2}}},\ \ \ \ \ \ \ y_{2}=\sqrt{1+\frac{128h^{2}|m_{s}|^{2}}{u^{2}H^{2}}}.$ (B.4)
+$y_{1}=\sqrt{1+\frac{128h^{2}\mu^{2}}{u^{2}}},\ \ \ \ \ \ \ y_{2}=\sqrt{1+\frac{128h^{2}|m_{s}|^{2}}{u^{2}H^{2}}}.$ (C.4)
 The primitive function above can be calculated and given by
 $F(y)=$ $\frac{u}{8\sqrt{2}h}\Bigg{\{}\sqrt{1-b+by-y^{2}}-\frac{b}{2}\arctan\left[\frac{b-2y}{2\sqrt{1-b+by-y^{2}}}\right]$
-$-\frac{\sqrt{2b}}{2}\arctan\left[\frac{2(1-y)+b(y-1-2)}{2\sqrt{2b}\sqrt{1-b+by-y^{2}}}\right],$ (B.5)
+$-\frac{\sqrt{2b}}{2}\arctan\left[\frac{2(1-y)+b(y-1-2)}{2\sqrt{2b}\sqrt{1-b+by-y^{2}}}\right],$ (C.5)
 For two $\arctan x$ in $F(y)$ at $y_{2}$, the first one is negative $-\pi/2$ while the second one is positive $\pi/2$ due to $b-2y_{2}<0$ and $02(1-y_{2})+b(y_{2})-1-2c>$ 0.
 On the other hand, the modes will be frozen-in when crossing the horizon hence we take the upper limit at $k|c_{s}|=aH$, i.e.,
-$y_{1}=\sqrt{1+\frac{128h^{2}}{u^{2}|c_{s}|^{2}}}$ (B.6)
+$y_{1}=\sqrt{1+\frac{128h^{2}}{u^{2}|c_{s}|^{2}}}$ (C.6)
 Summarizing the above computation, we finally have
 […]
-$-\frac{\sqrt{2b}}{2}\left[\frac{\pi}{2}-\arctan\left[\frac{2(1-y_{1})+b(y_{1}-3)}{2\sqrt{2b}\sqrt{1-b+by_{1}-y_{1}^{2}}}\right]\right]\Bigg{\}}.$ (B.7)
+$-\frac{\sqrt{2b}}{2}\left[\frac{\pi}{2}-\arctan\left[\frac{2(1-y_{1})+b(y_{1}-3)}{2\sqrt{2b}\sqrt{1-b+by_{1}-y_{1}^{2}}}\right]\right]\Bigg{\}}.$ (C.7)
 We can simple take the super-horizon limit $\mu\to 0$ and in the slow-roll limit,
-$u^{2}\simeq\frac{8h^{2}(1+6h^{2})}{1+2h^{2}},\ \ \ \ \ \ \ b\simeq\frac{8(1+2h^{2})}{1+6h^{2}},\ \ \ \ \ \ \ y_{1}\simeq 1.$ (B.8)
+$u^{2}\simeq\frac{8h^{2}(1+6h^{2})}{1+2h^{2}},\ \ \ \ \ \ \ b\simeq\frac{8(1+2h^{2})}{1+6h^{2}},\ \ \ \ \ \ \ y_{1}\simeq 1.$ (C.8)
 Then the function $\mathcal{I}$ can be reduced to a simply form
-$\mathcal{I}_{\text{approx}}=\sqrt{2}\pi\left(2-\sqrt{\frac{1+6h^{2}}{1+2h^{2}}}\right)h.$ (B.9)
+$\mathcal{I}_{\text{approx}}=\sqrt{2}\pi\left(2-\sqrt{\frac{1+6h^{2}}{1+2h^{2}}}\right)h.$ (C.9)
 […]
-Appendix C Bogoliubov coefficients
+Appendix D Bogoliubov coefficients
 For the shape case, we use the top-hat profile of the $h(N)$ for analytical computations,
-$h(N)=h_{0}\left[\theta(t-t_{1})-\theta(t-t_{2})\right].$ (C.1)
+$h(N)=h_{0}\left[\theta(t-t_{1})-\theta(t-t_{2})\right].$ (D.1)
 Before $t_{1}$, where we denote as region-I, the curvature and entropy modes are decoupled.
@@
 $\hat{\zeta}_{c}^{\text{I}}(t,\boldsymbol{k})=$ $u_{k}(t)\hat{a}_{\zeta}(\boldsymbol{k})+\text{h.c.}(-\boldsymbol{k}),$
-^$\hat{\mathcal{F}}^{\text{I}}(t,\boldsymbol{k})=$ $u_{k}(t)\hat{a}_{\mathcal{F}}(\boldsymbol{k})+\text{h.c.}(-\boldsymbol{k}),$ (C.2)
+^$\hat{\mathcal{F}}^{\text{I}}(t,\boldsymbol{k})=$ $u_{k}(t)\hat{a}_{\mathcal{F}}(\boldsymbol{k})+\text{h.c.}(-\boldsymbol{k}),$ (D.2)
 where $u_{k}$ is given by (4.10).
@@
 $\hat{\zeta}_{c}^{\text{I}}(t_{1}^{-})=$ $\hat{\zeta}_{c}^{\text{II}}(t_{1}^{+}),\ \ \ \ \ \ \ \dot{\hat{\zeta}}_{c}^{\text{I}}(t_{1}^{-})=D_{t}\hat{\zeta}_{c}^{\text{II}}(t_{1}^{+}),$
-^$\hat{\mathcal{F}}^{\text{I}}(t_{1}^{-})=$$\hat{\mathcal{F}}^{\text{II}}(t_{1}^{+}),\ \ \ \ \ \ \ \dot{\hat{\mathcal{F}}}^{\text{I}}(t_{1}^{-})=\dot{\hat{\mathcal{F}}}^{\text{II}}(t_{1}^{+}),$ (C.3)
+^$\hat{\mathcal{F}}^{\text{I}}(t_{1}^{-})=$$\hat{\mathcal{F}}^{\text{II}}(t_{1}^{+}),\ \ \ \ \ \ \ \dot{\hat{\mathcal{F}}}^{\text{I}}(t_{1}^{-})=\dot{\hat{\mathcal{F}}}^{\text{II}}(t_{1}^{+}),$ (D.3)
 where $D_{t}\hat{\zeta}_{c}\equiv\dot{\hat{\zeta}}_{c}-\sqrt{2}h\mathcal{A}H\hat{\mathcal{F}}$.
@@
 $\hat{\zeta}_{c}^{\text{II}}(t_{2}^{-})=$ $\hat{\zeta}_{c}^{\text{III}}(t_{2}^{+}),\ \ \ \ \ \ \ D_{t}\hat{\zeta}_{c}^{\text{II}}(t_{2}^{-})=\dot{\hat{\zeta}}_{c}^{\text{III}}(t_{2}^{+}),$
-^$\hat{\mathcal{F}}^{\text{II}}(t_{2}^{-})=$$\hat{\mathcal{F}}^{\text{III}}(t_{2}^{+}),\ \ \ \ \ \ \ \dot{\hat{\mathcal{F}}}^{\text{II}}(t_{2}^{-})=\dot{\hat{\mathcal{F}}}^{\text{III}}(t_{2}^{+}).$ (C.5)
+^$\hat{\mathcal{F}}^{\text{II}}(t_{2}^{-})=$$\hat{\mathcal{F}}^{\text{III}}(t_{2}^{+}),\ \ \ \ \ \ \ \dot{\hat{\mathcal{F}}}^{\text{II}}(t_{2}^{-})=\dot{\hat{\mathcal{F}}}^{\text{III}}(t_{2}^{+}).$ (D.5)
 We have eight coefficients $E_{i}$, $F_{i}$, $G_{i}$ and $H_{i}$, where $i=\zeta,\mathcal{F}$ and can be determined by the eight boundary conditions at $t_{2}$.
@@
 $H_{\zeta}=$ $\frac{i\sqrt{2}h\mathcal{A}e^{i2\kappa\cosh\left(\frac{\delta N}{2}\right)}}{4\kappa\left(S_{+}^{2}-S_{-}^{2}\right)}\sum_{\pm}\pm\frac{1}{S_{\pm}}\Bigg{\{}e^{iS_{\pm}\delta N}\left[\kappa^{2}-S_{\pm}(i+\kappa)\right]\left(i+\kappa+S_{\pm}\right)$
-$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ -e^{-iS_{\pm}\delta N}\left[\kappa^{2}+S_{\pm}\left(i+\kappa\right)\right]\left(i+\kappa-S_{\pm}\right)\Bigg{\}},$ (C.6)
+$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ -e^{-iS_{\pm}\delta N}\left[\kappa^{2}+S_{\pm}\left(i+\kappa\right)\right]\left(i+\kappa-S_{\pm}\right)\Bigg{\}},$ (D.6)
 where $\kappa\equiv k/k_{f}$, $\delta N\equiv H(t_{2}-t_{1})$ and $S_{\pm}\equiv\omega_{\pm}/H$.
@@
 $G_{\zeta}=$ $\frac{e^{-3\delta N/2}}{3\sqrt{3}}\left(2e^{3\delta N}+1-i\frac{6}{x_{2}^{3}}-ie^{2\delta N}\frac{3}{x_{2}}\right)\sin{\left(2\sqrt{6}h\delta N\right)}+\cdots,$
-$H_{\zeta}=$ $\frac{e^{-3\delta N/2}}{3\sqrt{3}}\left(2e^{3\delta N}-1-i\frac{6}{x_{2}^{3}}-ie^{2\delta N}\frac{3}{x_{2}}\right)\sin{\left(2\sqrt{6}h\delta N\right)}+\cdots.$ (C.7)
+$H_{\zeta}=$ $\frac{e^{-3\delta N/2}}{3\sqrt{3}}\left(2e^{3\delta N}-1-i\frac{6}{x_{2}^{3}}-ie^{2\delta N}\frac{3}{x_{2}}\right)\sin{\left(2\sqrt{6}h\delta N\right)}+\cdots.$ (D.7)
 We find that although the results are complex and the imaginary parts are the functions of $x_{2}$, in $(E_{\zeta}-F_{\zeta})$ and $(G_{\zeta}-H_{\zeta})$, the imaginary parts will be killed to each other.

### v3

@@
 We revisit inflation coupled with vector fields employing kinetic coupling in the comoving gauge.
-It is known that there is a cumulative effect $\tilde{I}N^{2}$ on curvature power spectrum.
-For a large number of e-foldings $N$, this contribution is so significant that it could violate observational constraints when the ratio of kinetic energy between vector fields and inflaton $\tilde{I}$ is not extremely small.
-In this paper, we explore the regime where $\tilde{I}\gg 1$, a realm that has not been extensively explored due to the limitations of perturbative methods.
+It is known that there is a cumulative effect $IN^{2}$ on the curvature power spectrum.
+For a large number of e-foldings $N$, this contribution is so significant that it could violate observational constraints when the ratio of kinetic energy between the vector fields and the inflaton $I$ is not extremely small.
+In this paper, we explore a regime where $I\gg 1$.
+This regime has not been extensively explored due to the limitations of perturbative methods.
 We found that the entropy perturbation becomes heavy in this regime and the cumulative effect decays away on super-horizon scales.
@@
 1 Introduction
-Inflationary physics has proven successful in phenomenologically explaining the large-scale structure of our universe.
+Inflationary physics has proven to be successful in phenomenologically explaining the large-scale structure of our universe.
 The simplest inflationary model, single-field inflation, predicts nearly scale-invariant, Gaussian, and adiabatic primordial perturbations.
@@
 These inflationary models are expected to be effective field theories derived from UV completed theories, such as string theory.
-Despite the anticipated suppression of higher-dimension irrelevant operators by the UV cutoff, they play a crucial role at the low-energy scale during inflation.
+Despite an anticipated suppression of higher-dimension irrelevant operators by UV cutoff, they play a crucial role at the low-energy scale during inflation.
 The UV sensitivity of inflation gives rise to issues like the $\eta$-problem [4, 5] and the super-Planckian problem of a single large field [6, 7, 5], disrupting the flatness of the scalar field potential.
-The study of introducing extra fields during inflation has attracted attention to evade the issues of single-field inflation.
+Study of introducing extra fields during inflation has attracted attention to evade these issues of single-field inflation.
 In particular, the study of multi-scalar fields exhibiting strongly non-geodesic motion has been extensively examined in recent years [8, 9, 10, 11, 12, 13, 14].
 In contrast, our focus is on U(1) gauge fields, commonly known as vector fields, as additional fields during inflation.
-This choice stems not only from the ubiquity of gauge fields as essential components of the standard model (SM) and beyond but also from their rich phenomenological implications in cosmology [15, 16, 17, 18, 19].
-The breaking of the conformal coupling between vector fields and gravity is necessary to generate sufficiently strong vector fields [19].
-Analogous to the non-geodesic motion of scalar fields, vector fields should exhibit background dynamics and participate alongside scalar fields in driving inflation.
+This choice stems not only from ubiquity of gauge fields as essential components of the standard model (SM) and beyond but also from their rich phenomenological implications in cosmology [15, 16, 17, 18, 19].
+Breaking of conformal couplings between vector fields and gravity is necessary to generate sufficiently strong vector fields [19].
+Analogous to the non-geodesic motion of scalar fields, vector fields should exhibit background dynamics and participate alongside the scalar fields in driving inflation.
 The inflationary phase undergoes a transition to new attractors if the potential is steep enough [20], which is very similar to the hyperbolic inflation [8].
@@
 The vector field is coupled to gravity through $f^{2}(\phi)F_{\mu\nu}F^{\mu\nu}$.
-After suitable choices of $f(\phi)$, the dynamics of background vector fields provides an effective potential for the slow-roll scalar field.
-Under this new attractor, the power spectrum of curvature perturbation is sourced by the vector field, yielding an additional contribution
-$\mathcal{P}_{\zeta}=\mathcal{P}_{\zeta}^{(0)}\left[1+\Theta(\theta)IN_{k}^{2}\right],\ \ \ \ \ \ (I\ll 1)$ (1.1)
-where $N_{k}\equiv\ln{(-k\tau_{e})}$ represents the e-folding number of observed mode elapsed since horizon-crossing, $\tau_{e}$ is time of end of the inflation.
-$I$ is the parameter related to specific models.
+After suitable choices of $f(\phi)$, dynamics of background vector fields provides an effective potential for the slow-roll scalar field.
+Under this new attractor, the power spectrum of the curvature perturbation is sourced by the vector fields, yielding an additional contribution
+$\mathcal{P}_{\zeta}=\mathcal{P}_{\zeta}^{(0)}\left[1+\Theta(\theta)\mathcal{I}N_{k}^{2}\right],\ \ \ \ \ \ (\mathcal{I}\ll 1)$ (1.1)
+where $N_{k}\equiv\ln{(-k\tau_{e})}$ represents the e-folding number of observed modes elapsed since horizon-crossing, $\tau_{e}$ is the end time of the inflation.
+$\mathcal{I}$ is a parameter that related to specific models.
 For one U(1) vector field, $\Theta(\theta)=24\sin^{2}\theta$, where $\theta$ is the angle between mode and direction of the vector field [22, 26].
-The term $IN_{k}^{2}$ accountable for the amplitude of statistical anisotropy and should be rigorously constrained by CMB observations to $I<10^{-7}$ for a sufficiently lengthy inflationary period $N_{k}\sim 60$ [22, 26].The exceedingly small parameter implies the presence of a very small but non-zero vector field, potentially giving rise to fine-tuning challenges [27, 28].
+The term $\mathcal{I}N_{k}^{2}$ accountable for the amplitude of statistical anisotropy and should be rigorously constrained by CMB observations to $\mathcal{I}<10^{-7}$ for a sufficiently lengthy inflationary period $N_{k}\sim 60$ [22, 26].
+The exceedingly small parameter implies the presence of a tiny but non-zero vector field, potentially giving rise to fine-tuning challenges [27, 28].
 Studies have revealed that employing multiple vector fields can mitigate the anisotropy, eventually converging to an isotropic attractor [24].
-For the isotropic configuration of a triad of vector fields, $\Theta(\theta)=16$, relaxing the constraint on $I$ [29].
-However, it still needs to be constrained to $I<10^{-5}$ due to the tilt of the power spectrum caused by the e-folding number $N$.
-This is a cumulative effect, i.e., the curvature is sourced by non-zero vector fields and evolves on super-horizon scales.
-However, this does not imply that results for $I\sim\mathcal{O}(1)$ are undesirable.
-If we require vector fields to handle a sufficiently large field displacement, indicating that $I$ is not a small parameter, it appears that the second term in (1.1) becomes significant and violates the adiabatic and scale-invariant initial conditions.
+For the isotropic configuration of a triad of vector fields, $\Theta(\theta)=16$, relaxing the constraint on $\mathcal{I}$ [29].
+However, it still needs to be constrained to $\mathcal{I}<10^{-5}$ due to the tilt of the power spectrum caused by the e-folding number $N_{k}$.
+This is a cumulative effect, i.e., the curvature is sourced by the non-zero vector fields and evolves on super-horizon scales.
+However, this does not imply that results for $\mathcal{I}\sim\mathcal{O}(1)$ are undesirable.
+If we require vector fields to handle a sufficiently large field displacement, indicating that $\mathcal{I}$ is not a small parameter, it appears that the second term in (1.1) becomes significant and violates adiabatic and scale-invariant initial conditions.
 However, this result lacks reliability because the perturbative method breaks down in this regime.
-It has been demonstrated that for $I\gtrsim 0.2$, the calculation by using in-in formalism starts to deviate from the numerical results and it is difficult to track the dynamics near and outside the horizon analytically [30, 31].
+It has been demonstrated that for $\mathcal{I}\gtrsim 0.2$, the calculation by using the in-in formalism starts to deviate from the numerical results and it is difficult to track the dynamics near and outside the horizon analytically [30, 31].
 Numerical findings indicate that the power spectrum does not actually evolve on super-horizon scales [31, 20].
-In this paper, we revisit this dilaton-gauge coupling model in comoving gauge and thoroughly explore the significant $I$ case with the assistance of effective field theory (EFT).
-We will show, the entropy perturbation is super heavy in this regime and the theory can be described by a low-energy EFT of a massless adiabatic perturbation.
-$IN_{k}^{2}$ does not occur in this regime; instead, it is replaced by an exponentially enhanced power spectrum, dependent solely on one parameter.
+In this paper, we revisit this dilaton-gauge coupling model in the comoving gauge and thoroughly explore the significant $\mathcal{I}$ case with the assistance of the effective field theory (EFT).
+We will show, that the entropy perturbation is super heavy in this regime and the theory can be described by a low-energy EFT of a massless adiabatic perturbation.
+$\mathcal{I}N_{k}^{2}$ does not occur in this regime; instead, it is replaced by an exponentially enhanced power spectrum, dependent solely on one parameter.
 This enhancement also provides a mechanism for generating primordial black holes.
 2 Inflation with vector fields
-We are considering the spatially flat Friedmann-Lemaitre-Robertson-Walker (FLRW) universe, which is characterized by the scale factor $a(t)$ and the Hubble variable is $H(t)=\dot{a}/a$, where ˙$\dot{}\equiv d/dt$ is cosmological time derivative.
-The Arnowitt-Deser-Misner(ADM) formalism of the metric is [32]
+We are considering a spatially flat Friedmann-Lemaitre-Robertson-Walker (FLRW) universe, which is characterized by the scale factor $a(t)$ and the Hubble variable is $H(t)=\dot{a}/a$, where ˙$\dot{}\equiv d/dt$ isthe cosmological time derivative.
+The Arnowitt-Deser-Misner(ADM) formalism for the metric is [32]
 $ds^{2}=-N^{2}dt^{2}+q_{ij}\left(N^{i}dt+dx^{i}\right)\left(N^{j}dt+dx^{j}\right),$ (2.1)
-where $N$ is the lapse function and we choose the cosmological time $N=1$ for background, $N^{i}$ is the shift vector and $g_{ij}$ is the metric of hypersurface.
-We consider the model of inflaton and $U(1)$ gauge fields, which is called dilaton-gauge inflation in this paper
+where $N$ is the lapse function and we choose the cosmological time $N=1$ for background, $N^{i}$ is the shift vector and $g_{ij}$ is the metric of the hypersurface.
+We consider a model of an inflaton and $U(1)$ gauge fields, which is called dilaton-gauge inflation in this paper
 $S=\int d^{4}x\sqrt{-{g}}\left[-\frac{1}{2}\partial^{\mu}\phi\partial_{\mu}\phi-V(\phi)-\frac{1}{4}f_{ab}(\phi)F^{a}_{\ \mu\nu}F^{b\mu\nu}\right],$ (2.2)
-where $g$ is the determinant of metric, $V(\phi)$ is the inflationary potential and $f_{ab}$ is the kinetic coupling and only depends on the inflaton.
-The strength of vector is $F^{a}_{\ \mu\nu}=\partial_{\mu}A^{a}_{\ \nu}-\partial_{\nu}A^{a}_{\ \mu}$.
+where $g$ is the determinant of the metric, $V(\phi)$ is the inflationary potential and $f_{ab}$ is the kinetic coupling, which only depends on the inflaton.
+The strength of the vector fields is $F^{a}_{\ \mu\nu}=\partial_{\mu}A^{a}_{\ \nu}-\partial_{\nu}A^{a}_{\ \mu}$.
 2.1 Background
-We consider one dilaton field $\phi$ as inflaton and a triplet of $U(1)$ fields and take an isotropic configuration of background evolution, which can be realized by the choice of $U(1)$ fields as following [33, 34]
+The configuration of the gauge fields is chosen for preserving isotropy [33, 34].
+One can consider a large number of randomly vector fields.
+These vector fields have about the same magnitude of order $A$ initially.
+Then, for example, [34]
+$\sum_{a}^{N\gg 1}A^{a}_{\ i}A^{a}_{\ j}\simeq\frac{N}{3}A^{2}\delta_{ij}+\mathcal{O}(1)\sqrt{N}A^{2}$ (2.3)
+is dominated by the $\delta_{ij}$ parts.
+The corrections proportional to $\sqrt{N}$ are due to stochastic random distribution of directions of the fields and they do not vanish for $i\neq j$.
+Moreover, if the coupling $f_{ab}$ are uniform, the isotropic configuration is an attractor in the phase space for the number of vector fields greater than two.
+This has been confirmed numerically in [24].
+If the field is a SU(2) non-Abelian gauge field, it’s natural to have this uniform coupling.
+The gauge coupling can be ignored during inflation thus SU(2) gauge field looks exactly like a triplet of U(1) gauge fields [35].
+Therefore in this paper we assume one dilaton field $\phi$ as the inflaton and a triplet of $U(1)$ fields and take an isotropic configuration of the background evolution, which can be realized by the following choice of $U(1)$ fields [36, 37]
 […]
-The identification of the spatial indices and internal indices of the field space allows us to achieve the internal gauge transformation of three $U(1)$ gauge fields by the $O(3)$ rotation of the three spatial dimensions in real space.
-We mention that the discussion doesn’t rely on the number of the $U(1)$ fields.
-If we have only one $U(1)$ field and kinetic coupling with dilaton, there may be anisotropy left in our universe when the energy density of $U(1)$ gauge field is not so small [22, 26].
-Although we will see, the theoretical parameterization of power spectrum is different when this energy density is large enough, we still consider isotropic case so that we can only treat scalar modes at linear order.
+The identification of the spatial indices and the internal indices of the field space allows us to achieve an internal gauge transformation of the three $U(1)$ gauge fields by a $O(3)$ rotation of the three spatial dimensions in real space.
+We note that the discussion doesn’t rely on the number of the $U(1)$ fields.
+If we have only one $U(1)$ field and kinetic coupling with dilaton, there may be anisotropy left in our universe when the energy density of the $U(1)$ gauge field is not so small [22, 26].
 Under this configurations the Friedmann equations are given by
@@
 It’s also useful to define $\dot{\sigma}^{2}\equiv\dot{\phi}^{2}/2+f^{2}\dot{\mathbb{A}}^{2}/a^{2}$ as the “kinetic” energy of the matter fields.
-Then another gravitational equation can be written as
+Another gravitational equation can then be written as
 $H^{2}M_{\text{pl}}^{2}\epsilon=\dot{\sigma}^{2}.$ (2.6)
@@
 In this paper we assume that $\dot{\phi}<0$ and $\dot{\mathbb{A}}>0$.
-We also define a useful variable, the ratio of the kinetic energy of $U(1)$ fields and dilaton [^1] The parameter $I$ in [29] or in (1.1) is related to $h$ in our paper as $I\simeq 2h^{2}$ for $h\ll 1$.
+We also define a useful variable, the ratio of the kinetic energy of $U(1)$ fields and the dilaton [^1] The parameter $\mathcal{I}$ in [29] or in (1.1) is related to $h$ in our paper as $\mathcal{I}\simeq 2h^{2}$ for $h\ll 1$.
 $h\equiv\sqrt{\frac{\epsilon_{A}}{2\epsilon_{\phi}}}.$ (2.8)
-But here the ratio $h$ needn’t to be a small parameter.
+But here, the ratio $h$ needn’t to be a small parameter.
 In fact, we will consider the larger $h\gg 1$ on small scales for PBH’s production.
 We will also define the slow-roll parameter $\epsilon_{h}\equiv\dot{h}/(Hh)$, $\eta_{h}\equiv\dot{\epsilon}_{h}/(H\epsilon_{h})$, $\eta_{\phi}\equiv\ddot{\phi}/(H\dot{\phi})$ and $\eta_{A}\equiv\dot{\epsilon}_{A}/(H\epsilon_{A})$.
-Taking time derivative on the Friedmann equation one can obtain
+Taking the time derivative of the Friedmann equation one obtains
 $\sqrt{2\epsilon_{\phi}}V_{\phi}=M_{\text{pl}}H^{2}\left(6\epsilon+2\epsilon_{\phi}\eta_{\phi}+\frac{3}{2}\epsilon_{A}\eta_{A}-3\epsilon\epsilon_{A}\right).$ (2.9)
@@
 2.2 Quadratic action in comoving gauge
-Thanks to the isotropic background and we will work in the quadratic action and the linear fluctuations of the theory hence the vector and tensor fluctuations are decoupled from scalar ones at linear order.
+Thanks to the isotropic background and we will work in the quadratic action and the linear fluctuations of the theory.
+Hence, the vector and tensor fluctuations are decoupled from the scalar ones at linear order.
 After using the helicity decomposition the scalar parts of the $U(1)$ fields can be written as
 […]
-We do not distinguish indices $a$ and $i$ here because we identified the spatial rotation symmetry with the internal global $O(3)$ symmetry of the space of the gauge fields.
-The gauge fields enjoys local $U(1)$ symmetry $A^{a}_{\ \mu}\rightarrow A^{a}_{\ \mu}+\partial_{\mu}\rho^{a}$, where $\rho^{a}$ is a arbitrary function.
-One can fix the scalar gauge freedom of $U(1)$ fields by $\mathbb{M}=0$ [20].
-The $\mathbb{Y}$ is the constrain d.o.f and we will eliminate it latter after fixing the gauge freedoms of gravity.
-The scalar fluctuations of gravity and scalar fields can be written as
+We do not distinguish between the indices $a$ and $i$ here because we identified the spatial rotation symmetry with the internal global $O(3)$ symmetry of the space of the gauge fields.
+The scalar fluctuations of gravity and the scalar field can be written as
 $N=1+\alpha,\ \ \ \ N_{i}=\partial_{i}\beta,\ \ \ \ g_{ij}=a^{2}(t)\left(e^{2\zeta}\delta_{ij}+2\partial_{i}\partial_{j}E\right),\ \ \ \ \phi=\phi(t)+\delta\phi.$ (2.12)
+We have analysed the vector and tensor perturbations in our previous work [20].
+The $U$ and the vector perturbations are massless (see (5.10) of [20]).
+Therefore, there is no instability in these perturbations so we didn’t discuss therm in this work.
+The tensor perturbations are almost the same as the single-filed inflation thus they are not interesting for our discussion (see (5.14) of [20]).
+We have a total of nine scalar degrees of freedom, where four from the metric and five from the matter sector.
 There are four gauge freedoms of the time and spatial coordinate choice.
-If we have scalar field, in addition to vector fields we still have one physical d.o.f in the theory.
+If we have one scalar field, there are two scalar gauge freedoms in the theory.
+On the other hand, if a triple of gauge fields is considered, it enjoys a local $U(1)$ symmetry $A^{a}_{\ \mu}\rightarrow A^{a}_{\ \mu}+\partial_{\mu}\rho^{a}$, where $\rho^{a}$ is an arbitrary function.
+This allows for another one gauge freedom.
+Fixing each gauge freedom removes one constraint.
+In conclusion, we can eliminate $2\times(2+1)=6$ d.o.f, leaving $3$ remaining.
+One can fix the scalar gauge freedom of the $U(1)$ fields by $\mathbb{M}=0$ [20].
 The spatial flat gauge is usually used to study the dilaton-gauge theory.
-In this gauge, this physical d.o.f is the fluctuation of the scalar field $\delta\phi$.
+In this gauge, the physical d.o.f are the fluctuation of the scalar field $\delta\phi$.
 On the other hand, the “comoving” gauge is also used in the two-scalar fields [38, 39].
@@
 The physical d.o.f are the curvature and entropy fluctuations.
-In this paper we adopt the similar comoving gauge for the dilaton-gauge model.
+In this paper we adopt a similar comoving gauge for the dilaton-gauge model.
 The adiabatic fluctuation is given by $Q_{\sigma}\equiv-\sqrt{2}\dot{\sigma}\delta u$, where $\delta u$ is the velocity potential given by $\delta T^{0}_{i}\equiv(\rho+p)\partial_{i}\delta u$.
 For the isotropic one-form gauge fields, we have $\delta T^{0}_{\ i}=-\dot{\phi}\partial_{i}\delta\phi-(2f^{2}\dot{\mathbb{A}}/a^{2})\partial_{i}\delta\mathbb{A}$ [20].
-Here $\rho$ and $p$ are the total energy density and pressure of the system and in our model we have $\rho+p=\dot{\phi}^{2}+2f^{2}\dot{\mathbb{A}}^{2}/a^{2}=2\dot{\sigma}^{2}$.
-It’s useful to define the variable $\delta Q=\sqrt{2}f\delta\mathbb{A}/a$.
+Here $\rho$ and $p$ are the total energy density and pressure of the system, and in our model, we have $\rho+p=\dot{\phi}^{2}+2f^{2}\dot{\mathbb{A}}^{2}/a^{2}=2\dot{\sigma}^{2}$.
+It’s useful to define a variable $\delta Q=\sqrt{2}f\delta\mathbb{A}/a$.
 Then the adiabatic fluctuation is given by
@@
 $Q_{\sigma}=0,\ \ \ \ \ \ E=0.$ (2.14)
-This is the unitary gauge in the gauge theory, whose Nambu-Goldstone boson is eaten by the metric, as we will see in the next section.
+This is the unitary gauge in the gauge theory, whose Nambu-Goldstone boson is eaten by the metric.
 Then we have fixed all gauge freedom of the theory.
-The d.o.f $\mathbb{U}$ in the vector fields is the magnetic fluctuation and decouple with any other modes in our configuration.
-It only contribute to the isocurvature fluctuation hence we ignore it in the following discussion.
+The d.o.f $\mathbb{U}$ in the vector fields is the magnetic fluctuation and decouples from any other modes in our configuration.
+It only contributes to the isocurvature fluctuation hence we ignore it in the following discussion.
 We also define the entropy fluctuations.
@@
 $\mathcal{F}\equiv\cos{\vartheta}\delta Q-\sin{\vartheta}\delta\phi,$ (2.16)
-which is similar to the entropy fluctuation in the two-scalar fields.
-But here the entropy fluctuations is not equivalent to the isocurvature one because we still have another contribution $\mathbb{U}$, which is decoupled with $\zeta$ and $\mathcal{F}$.
-After imposing the comoving gauge, the equation of motion of the non-dynamical d.o.f $\alpha$ and $\mathbb{Y}$ can be derived and the solutions read
-$\alpha=$ $\frac{\dot{\zeta}}{H},$ (2.16)
-$\partial^{2}\mathbb{Y}=$ $\dot{\mathbb{A}}\left(\zeta-\frac{\dot{\zeta}}{H}\right)-\frac{a}{\sqrt{2}f}\left(\frac{\mathcal{F}}{\sqrt{1+2h^{2}}}\right)^{.}$
-$-\frac{a}{\sqrt{2}f}\left(H-\frac{\dot{f}}{f}+4f_{\phi}\frac{\dot{\mathbb{A}}}{a}h\right)\frac{\mathcal{F}}{\sqrt{1+2h^{2}}}.$ (2.17)
-The d.o.f $\beta$ is totally eliminated from the quadratic action when we use the constrain equation of $\alpha$ hence we don’t need to write down its specific expression.
+which is similar to the entropy fluctuation in systems of two-scalar fields.
+But here the entropy fluctuations is not equivalent to the isocurvature one because we still have another contribution $\mathbb{U}$, which is decoupled from the $\zeta$ and $\mathcal{F}$.
+After imposing the comoving gauge, the solutions of the equations of motion of the non-dynamical d.o.f $\alpha$, $\beta$ and $\mathbb{Y}$ can be derived (see Appendix A).
 Then after using the background equations (2.7) and (2.10) to eliminate the $V$, $V_{\phi}$ and $f_{\phi}$, and inserting the constrain equations into the quadratic Lagrangian, we obtain
@@
 where we have discarded slow-roll suppressed terms in the slow-low limit and also assumed $|V_{\phi\phi}|/H^{2}\ll 1$.
-This is the decoupling limit $\epsilon\to 0$ in the EFT.
-Then $\epsilon_{\phi}$ and $\epsilon_{A}$ are also vanishing because they are positive quantities.
+This is the decoupling limit $\epsilon\to 0$ so that the higher-order terms of $\epsilon$ are discarded.
+The higher order of $\epsilon_{\phi}$ and $\epsilon_{A}$ are also vanishing because they are positive quantities as we have $\epsilon=\epsilon_{\phi}+\epsilon_{A}$.
 We find that the curvature and entropy fluctuations are both massive.
 The mass term of $\zeta$ is only dependent on $h$.
-For the mass term of $\mathcal{F}$, the second term can be regards as the contribution from the curved field space, whose curvature is given by $R_{\text{fs}}=-f_{\phi\phi}/f$.
+For the mass term of $\mathcal{F}$, the second term can be regarded as the contribution from the curved field space, whose curvature is given by $R_{\text{fs}}=-f_{\phi\phi}/f$.
 If we assume $h=0$ initially, where we are in the single-field inflation regime, the entropy mass $m_{s}^{2}$ is still non-vanishing and have large negative contribution if the curvature of field space is negatively large enough.
@@
 $m_{s}^{2}\simeq-\frac{8h^{2}\left(3+2h^{2}\right)}{1+2h^{2}}H^{2}.$ (2.19)
-Therefore the entropy fluctuation has imaginary mass.
+Therefore the entropy fluctuation has an imaginary mass.
 This is similar to the so-called geometrical destabilization [11, 12], which is firstly studied in the multi-scalar inflation and then in the inflation with multiple vector fields [20].
-The differences are that the curvature fluctuation $\zeta$ is massive and we have an additional coupling $\zeta\mathcal{F}$.
+The differences are that the curvature fluctuation $\zeta$ is massive, and we have an additional coupling $\zeta\mathcal{F}$.
 Hence we next investigate the mode solutions of this Lagrangian.
@@
 […]
-where $N_{c}$ is the time of horizon-crossing, $N_{e}$ is the time of end of inflation and $N_{k}\equiv N_{e}-N_{c}$ is the e-folding number elapsed since horizon-crossing.
+where $N_{c}$ is the time of horizon-crossing, $N_{e}$ is the end time of inflation and $N_{k}\equiv N_{e}-N_{c}$ is the e-folding number elapsed since horizon-crossing.
 For small $h$ this is a small correction to the adiabatic mode.
@@
 3 EFT under massive entropy mode
-3.1 Modified dispersion relation
-From the quadratic Lagrangian (2.2) we can write down the linear equations of motion
+In this system we have two coupled d.o.f. From the quadratic Lagrangian (2.2) we can write down the linear equations of motion of the system:
 $\ddot{\zeta}+\left(3+\eta\right)H\dot{\zeta}+\left(\frac{k^{2}}{a^{2}}+m_{\sigma}^{2}\right)\zeta=$ $\frac{1}{a^{3}\epsilon}\left(a^{3}\frac{\dot{\sigma}h\mathcal{A}}{M_{\text{pl}}^{2}}\mathcal{F}\right)^{.}-\frac{\dot{\sigma}hH\mathcal{B}}{\epsilon M_{\text{pl}}^{2}}\mathcal{F},$
 ¨$\ddot{\mathcal{F}}+3H\dot{\mathcal{F}}+\left(\frac{k^{2}}{a^{2}}+m_{s}^{2}\right)\mathcal{F}=$ $-2\dot{\sigma}h\left(\mathcal{A}\dot{\zeta}+\mathcal{B}H\zeta\right).$ (3.1)
-We are interested in the decoupling limit where the nontrivial solution $\zeta=\text{constant}$ is permitted hence we drop out the small slow-roll parameters.
-Then we have $\mathcal{A}=4$ and $\mathcal{B}=m_{\sigma}/H^{2}$.
-We regard $\dot{\sigma}$ and $h$ as constants.
-To find the dispersion relation we study the short-wavelength limit where the Hubble friction terms can be disregarded.
-We have noted that there is also a mass term for curvature fluctuation.
-The upper bound of the mass is $m_{\sigma}^{2}\lesssim 8H^{2}$, i.e., the mass of $\zeta$ is at most of the same order as the energy scale of inflation.
-In the short-wavelength limit the contribution of $m_{\sigma}$ term can also be ignored.
-Then the equations of motion read
-$\ddot{\zeta_{c}}+p^{2}\zeta_{c}=$ $\sqrt{2}hH\mathcal{A}\dot{\mathcal{F}},$
-¨$\ddot{\mathcal{F}}+p^{2}\mathcal{F}+m_{s}^{2}\mathcal{F}=$ $-\sqrt{2}hH\mathcal{A}\dot{\zeta}_{c},$ (3.2)
-where $p=k/a$ is the physical momentum and $\zeta_{c}\equiv M_{\text{pl}}\sqrt{2\epsilon}\zeta$ is the canonical variable. The schematic solutions of these equations can be written as [43]
-$\zeta_{c}=\zeta_{+}e^{i\omega_{+}t}+\zeta_{-}e^{i\omega_{-}t},$
-$\mathcal{F}=\mathcal{F}_{+}e^{i\omega_{+}t}+\mathcal{F}_{-}e^{i\omega_{-}t},$ (3.3)
-where the two frequencies $\omega_{\pm}$ are given by
-$\omega_{\pm}^{2}=\frac{m_{s}^{2}}{2c_{s}^{2}}+p^{2}\pm\frac{m_{s}^{2}}{2c_{s}^{2}}\sqrt{1+\frac{4\left(1-c_{s}^{2}\right)p^{2}}{m_{s}^{2}c_{s}^{-2}}}$ (3.4)
-and the speed of sound is given by
-$\frac{1}{c_{s}^{2}}=1+\frac{32H^{2}h^{2}}{m_{s}^{2}}.$ (3.5)
-The modification of the dispersion relation of fluctuations is provided by the interaction with $\dot{\zeta}{\mathcal{F}}$.
-The heavy entropy fluctuation implies the hierarchy of the two frequencies $\omega_{-}\ll\omega_{+}$, which requires $4(1-c_{s}^{2})p^{2}\ll m_{s}^{2}c_{s}^{-2}$.
-Hence we have a cut-off scale of EFT (B.8) on momentum $p_{\text{UV}}$ and the EFT is only valid below this scale.
-The UV scale is given by
-$p_{\text{UV}}^{2}=\frac{|m_{s}^{2}c_{s}^{-2}|}{4\left(1-c_{s}^{2}\right)}.$ (3.6)
-In the low energy scales, the two scalar modes, which are corresponded to two frequencies, oscillate coherently and we integrate out the heavy mode to eliminate the high-frequency mode from the dynamics of the curvature fluctuation.
-3.2 Integrating out the heavy $\mathcal{F}$
-The fluctuation is heavy for $h\gg 1$ and we also have $|m_{s}^{2}|\gg m_{\sigma}^{2}$ as we have mentioned before.
-For positive $m_{s}^{2}$, the operator from coupling $\dot{\zeta}\mathcal{F}$ would reduce the speed of sound of the fluctuation [43].
-However from (2.20) the mass square is negative due to the contribution from kinetic coupling $f(\phi)$.
-We can ignored the slow-roll suppressed terms and the entropy mass is given by (2.20).
-Then the speed of sound in this decoupling limit is
-$\frac{1}{c_{s}^{2}}=1-\frac{4\left(1+2h^{2}\right)}{3+2h^{2}}.$ (3.7)
-The speed of sound $c_{s}^{2}\simeq-1/3$ is imaginary for heavy entropy mass $h\gg 1$.
-The imaginary speed of sound has been studied in the two-scalar inflation recently [12, 14, 44].
-And the entropy mass-square in the two-scalar models is also imaginary hence is tachyonic before the sound horizon crossing $p^{2}c_{s}^{2}\sim H^{2}$.
-The curvature fluctuation sourced by the entropy fluctuation also experiences transient instability before horizon crossing.
-After the horizon-crossing, the dynamics of fluctuations are influenced by the background.
-In other words, the mass term $m_{\sigma}$ and the Hubble terms become important outside the horizon.
-Figure 1: Evolution of entropy perturbation $\mathcal{F}$ for $h=10$. An instability before horizon-crossing and constant evolution on super-horizon are presented. The scales are arbitrary.
-We have noted that the speed of sound is $|c_{s}|\sim\mathcal{O}(1)$ for large $h$ hence the non-linear dispersion regime is narrower, and we can directly investigate the linear dispersion regime (see Appendix B for reviews).
-In the heavy limit $|m_{s}^{2}|\gg\omega^{2}$ we can expand the operator
+Figure 1: Evolution of the entropy perturbation $\mathcal{F}$ for $h=10$. An instability before horizon-crossing and a constant evolution on super-horizon are presented. The scales are arbitrary.
+In the heavy limit $|m_{s}^{2}|\gg\omega^{2}$ we can expand the operator as
 $\frac{1}{m_{s}^{2}-\square}=\frac{1}{m_{s}^{2}}+\frac{\square}{m_{s}^{4}}+\cdots,\ \ \ \ \ \ \text{where}\ \ \square\equiv-\frac{\partial^{2}}{\partial t^{2}}-3H\frac{\partial}{\partial t}+\frac{\partial^{2}}{a^{2}}.$ (3.2)
-Then the leading-order of the linear equation of motion of $\mathcal{F}$ is given by
+The solution of $\mathcal{F}$ is dominated by the $1/m_{s}^{2}$ term, which is given by
 $\mathcal{F}_{\text{LO}}=-\frac{2\dot{\sigma}h(\mathcal{A}\dot{\zeta}+\mathcal{B}H\zeta)}{m_{s}^{2}}.$ (3.3)
 We show the consistent of the leading-order contribution of $\mathcal{F}$ in Figure 1 for large enough $h$.
-The evolution of curvature perturbation on large scales is sourced by non-adiabatic perturbation $\dot{\zeta}=-\delta P_{\text{nad}}H/(\rho+p)$ [45, 46, 47].
+In the slow-roll limit, the entropy mass $m_{s}$ can be calculated by the (2.19). From (B.6) the EFT is only valid under the scale
+$p_{\text{UV}}=\frac{3}{4}|m_{s}|.$ (3.4)
+Then after inserting the equation (3.3) into the quadratic action (2.2) we obtain
+$\mathcal{L}^{(2)}_{\text{LO}}=a^{3}M_{\text{pl}}^{2}\frac{\epsilon}{c_{s}^{2}}\left[\dot{\zeta}^{2}-\frac{c_{s}^{2}}{a^{2}}\left(\partial\zeta\right)^{2}\right],$ (3.5)
+where $c_{s}$ is the speed of sound of $\zeta$
+$\frac{1}{c_{s}^{2}}=1-\frac{4\left(1+2h^{2}\right)}{3+2h^{2}}.$ (3.6)
+The operator from the coupling $\dot{\zeta}\mathcal{F}$ would reduce the speed of sound of the fluctuation [45].
+We have also used integration by parts to eliminate the $\dot{\zeta}\zeta$ terms in the action.
+Then we found that the mass term of the curvature fluctuation is cancelled and $\zeta$ is conserved on super-horizon scales.
+We can find that in the EFT of single field, we have only this constant curvature perturbation on large scales.
+Figure 2: The adiabaticity of the super-horizon ($|k\tau_{e}|=10^{-15}$) curvature perturbation.
+As $h$ increases from very small values, the perturbations become non-adiabatic.
+However, when $h$ exceeds $\mathcal{O}(0.1)$, the perturbation quickly becomes adiabatic.
+[…]
+Let’s make some comments on this result:
+The evolution of curvature perturbation on large scales is sourced by the non-adiabatic perturbation $\dot{\zeta}=-\delta P_{\text{nad}}H/(\rho+p)$ [47, 48, 49].
 Then from the leading-order solution of entropy perturbation, the non-adiabatic mode can be represented as
@@
 should be relatively small enough on large scales.
-We can see that the entropy fluctuation doesn’t decay outside the horizon, but is proportional to the curvature fluctuation if $\zeta=\text{constant}$ on super-horizon.
-We can find that in the EFT of single field, we have only this constant curvature perturbation on large scales.
-In the slow-roll limit the entropy mass $m_{s}$ can be calculated by the (2.20).
-From (3.6) the EFT is only valid under scale
-$p_{\text{UV}}=\frac{3}{4}|m_{s}|.$ (3.11)
-Then after inserting the equation (3.9) into the quadratic action (2.2) we obtain
-$\mathcal{L}^{(2)}_{\text{LO}}=a^{3}M_{\text{pl}}^{2}\frac{\epsilon}{c_{s}^{2}}\left[\dot{\zeta}^{2}-\frac{c_{s}^{2}}{a^{2}}\left(\partial\zeta\right)^{2}\right],$ (3.12)
-where $c_{s}$ is given by (3.7).
-We have used the integration by parts to eliminate the $\dot{\zeta}\zeta$ terms in the action.
-Then we found that the mass term of curvature fluctuation is cancelled and $\zeta$ is conserved on super-horizon scales.
-Figure 2: The adiabaticity of super-horizon ($|k\tau_{e}|=10^{-15}$) curvature perturbation.
-As $h$ increases from very small values, the perturbations become non-adiabatic.
-However, when $h$ exceeds $\mathcal{O}(0.1)$, the perturbation quickly becomes adiabatic.
-Different from multi-scalar inflation, where the non-decaying non-adibatic modes will source to curvature modes so that $\zeta$ evolves on super-horizon.
+We can see that the entropy fluctuation doesn’t decay outside the horizon but is proportional to the curvature fluctuation if $\zeta=\text{constant}$ on the super-horizon.
+Unlike multi-scalar inflation, where non-decaying non-adiabatic modes will source to curvature modes so that $\zeta$ evolves on the super-horizon.
 In dilation-gauge inflation, we have constant $\zeta$ for non-decaying $\mathcal{F}$.
-In other words, $\mathcal{F}$ doesn’t play the role of non-adiabatic mode in dilaton-gauge inflation.
-For mid-value $h\lesssim\mathcal{O}(0.1)$, where $|m_{s}|\sim H$, the sub-leading terms in $\delta P_{\text{nad}}$ become important and the curvature perturbation evolves on super-horizon.
-These non-adiabatic contributions are from the cumulative modes (2.27).
+In other words, For the large $h$ case, $\mathcal{F}$ doesn’t play the role of non-adiabatic mode in dilaton-gauge inflation.
 We show this non-adiabatic occurrence for mild $h\lesssim\mathcal{O}(0.1)$ in Figure 2.
 One of the most important findings from this result is that the cumulative modes (2.26) decay away for large $h$.
-Because for large $h$, the modes (3.9) on super-horizon are exactly the constant solutions of the adiabatic mode (2.23).
-The decaying modes can be also obtained through solving the equation of motion (3.1), which has been studied in spatially flat gauge in [20].
+Because for large $h$, the modes (3.3) on the super-horizon are exactly the constant solutions of the adiabatic mode (2.22).
+The decaying modes can be also obtained through solving the equation of motion (3), which has been studied in the spatially flat gauge in [20].
 The decaying mode is $\zeta\propto(-k\tau)^{(3-\sqrt{9-96h^{2}})/2}$.
@@
 4 Primordial black holes
-The remained adiabatic mode experiences a exponential growth before horizon-crossing due to the imaginary speed of sound.
+If gauge fields are not important to the inflation, which means that inflation is driven by a single scalar field, it may be ruinous because the higher-order loop corrections to the scalar-field’s mass state indicate that the potential has structure on scales of order of the cut-off of inflation [5].
+If the potential has sufficiently steep structure at some points in inflation, i.e., $V_{\phi}>3H\dot{\phi}$ [20], the gauge fields become unstable and can be switched on later.
+The parameter space for switching on gauge fields is wide making it not difficult for this to occur [50].
+We show a simple example in Figure 3.
+If the dynamics of inflation fall into the parameter space of single-field inflation, the gauge fields will be turned off.
+Figure 3: Example of switching gauge fields on during inflation.
+The model with the potential $V=m^{2}\phi^{2}/2$ and coupling $f=\text{Exp}(c\phi^{2}/M_{\text{pl}}^{2})$, where $m=10^{-6}M_{\text{pl}}$ and $c=3$.
+After switching on the gauge fields, the curvature mode experiences a exponential growth before horizon-crossing due to the imaginary speed of sound.
 Therefore the power spectrum is enhanced exponentially.
-If vector fields switch on only at late time, The power spectrum is enhanced on small scales so the formations of PBHs in radiation era is possible [48, 49].
-And these PBHs can be one of candidates of the dark matter [50, 51, 52] and supermassive black holes in our universe [53].
-The enhancement of the power spectrum due to the heavy fields is studied recently in multiple scalar-field models, and is used to discuss the PBHs formation [56, 54, 55].
-In previous section, we have realized that it also can be achieved by using vector fields.
+If the vector fields switch on only at late time, the power spectrum is enhanced on small scales, making the formations of PBHs in the radiation era is possible [51, 52].
+These PBHs can be candidates for dark matter [53, 54, 55] and supermassive black holes in our universe [56].
+The enhancement of the power spectrum due to heavy fields is studied recently in multiple scalar-field models, and is used to discuss PBHs formation [59, 57, 58].
+In the previous section, we have realized that it also can be achieved by using vector fields.
 We approach to the PBHs without any specific models ($V(\phi)$ and $f(\phi)$).
-We here only consider the switch-on of the vector fields at late time $N_{f}$ during inflation and see the features of power spectrum on small scales for generating PBHs.
+We here only consider the switch-on/off of vector fields at late time $N_{f}$ $only\ once$ and see the features of the power spectrum on small scales for generating PBHs.
 We choose the horizon-crossing time of the CMB scale $k_{\text{CMB}}=0.05\text{Mpc}^{-1}$ as $N_{\text{CMB}}=0$.
 Assuming that $h$ has peak at $N_{f}$.
-The curvature modes with different scales that cross the horizon at different time have different contributions to the power spectrum.
+The curvature modes with different scales that cross the horizon at different times have different contributions to the power spectrum.
 We can define the corresponding scale
 $k_{f}|c_{s}|=a(N_{f})H(N_{f}).$ (4.1)
-The modes $k\gg k_{f}$ that cross the horizon far after the switch-off time experienced nothing hence the power spectrum is the same as the single-field inflation $\mathcal{P}_{\zeta}^{(0)}=H^{2}/(8\pi\epsilon M_{\text{pl}}^{2})$ on small scales.
-On the other hand, the modes $k\ll k_{f}$ that cross the horizon far before the switch-on time will be impacted by the entropy fluctuation on super-horizon.
-The entropy fluctuation doesn’t decay outside the horizon hence always transfer to the curvature fluctuation.
-Hence on these large scales, the power spectrum is shifted so different from the single-field one.
-The interesting scales are the modes $k\sim k_{f}$ that cross the horizon near the switch-on time $N_{f}$ have large enhancements due to the large instability of the entropy fluctuation.
-Hence we expect that the power spectrum of the curvature fluctuation has exponential enhancements near the scale $k_{f}$.
-We can see from Figure 1 that the mode takes a few e-folding to grow before frozen-in outside the horizon.
-Hence there are two kinds of power spectrum, the broad ($\delta N\gtrsim\ln{(h_{0})}$) and the sharp ($\delta N\lesssim\ln{(h_{0})}$) switch-on of $h(N)$ [56].
-For the shape case, the curvature modes stop growing before crossing the horizon.
-4.1 Board case
-For this case, we have enough e-folding of $h(N)$ for exponential growth of the modes. The EFT is valid at time $N_{v}$ satisfying
+The modes $k\gg k_{f}$ that cross the horizon far after the switch-off time experienced nothing.
+Hence the power spectrum is the same as that for single-field inflation $\mathcal{P}_{\zeta}^{(0)}=H^{2}/(8\pi\epsilon M_{\text{pl}}^{2})$ on small scales.
+On the other hand, the modes $k\ll k_{f}$ that cross the horizon far before the switch-on time will be impacted by the entropy fluctuation on super-horizon scales.
+The entropy fluctuation doesn’t decay outside the horizon and thus always transfer to the curvature fluctuation.
+Hence on these large scales, the power spectrum is shifted and different from that of single-field one.
+The interesting scales are those where $k\sim k_{f}$; These modes cross the horizon near the switch-on time $N_{f}$ have large enhancements due to the large instability of the entropy fluctuation.
+We expect that the power spectrum of the curvature fluctuation has exponential enhancements near the scale $k_{f}$.
+As shown in Figure 1, the mode takes a few e-folding to grow before freezing outside the horizon.
+Hence there are two kinds of power spectrum, the broad and the sharp of $h(N)$ [59].
+For the sharp case, the curvature modes stop growing before crossing the horizon.
+4.1 Broad case
+The EFT is valid at time $N_{v}$ satisfying
 $\frac{k}{a(N_{v})}=\frac{3}{4}|m_{s}(N_{v})|.$ (4.2)
-Then the most contributions to the power spectrum are the modes with $N_{v}\sim N_{f}$. We here consider the Gaussian profile of the function
+Then the majority of contributions to the power spectrum come from modes with $N_{v}\sim N_{f}$. We here consider the Gaussian profile of the function
 $h_{G}(N)=h_{0}e^{-(N-N_{f})^{2}/(2\delta N^{2})},$ (4.3)
 where the profile is controlled by two parameters $(h_{0},\delta N)$.
-On super-horizon, the entropy modes do not decay and still couple to the curvature ones.
+The "broad" means $\delta N\gtrsim\ln{(h_{0})}$, indicating that we have enough e-folding of non-zero $h(N)$ for exponential growth of the modes.
+On super-horizon scales, the entropy modes do not decay and still couple to the curvature ones.
 Hence we should discuss large-scale modes $k\ll k_{f}$ and small-scale modes $k\sim k_{f}$ separately.
 Modes with $\boldsymbol{k\gtrsim k_{f}}$
-To characterize the peak of power spectrum with these modes, it’s also useful to introduce a dimensionless parameter $\mathcal{I}$ [44, 57].
-The quantization of the curvature fluctuation with imaginary speed of sound is quite different from the real one [44].
-The difference is that unlike the Bunch-Davies vacuum, the negative frequency modes are non-zero for initial state of EFT due to the different commutation conditions of the mode functions.
-Taking the super-horizon limit, the leading contribution of the power spectrum is
+To characterize the peak of the power spectrum for these modes, it’s also useful to introduce a dimensionless parameter $\mathcal{I}$ [46, 60].
+The quantization of the curvature fluctuation with imaginary speed of sound differs significantly from that of with real one [46].
+The difference lies in the fact that, unlike the Bunch-Davies vacuum, the negative frequency modes are non-zero for initial state of the EFT due to the different commutation conditions of the mode functions.
+Taking the super-horizon limit, the leading contribution of the the power spectrum is
 $\Delta_{\zeta}(k)\equiv\frac{\mathcal{P}_{\zeta}(k)}{\mathcal{P}_{\zeta}^{(0)}(k)}\simeq e^{2\mathcal{I}(h)}\big{|}_{N_{v}},$ (4.4)
@@
 $\mathcal{I}(h)\simeq\sqrt{2}\pi\left(2-\sqrt{\frac{1+6h^{2}}{1+2h^{2}}}\right)h.$ (4.5)
-For modes who start to experiences transient instability earlier, the amplitude will be larger after crossing the horizon.
+For modesthat begin to experience transient instability earlier, the amplitude will be larger after crossing the horizon.
 And we need to use a more UV EFT to describe these modes on deeper sub-horizon scales.
-Figure 3: The numerical and analytical results (4.4) of power spectrum for the Gaussian case, where $h_{0}=8$ and $\delta N=2$.
-The scale of peak $k_{p}$ is the one that has maximal $h$ at time when the modes is growing at $N_{v}$, i.e., $N_{p_{k}}=N_{v}$. Then we have
+Figure 4: The numerical and analytical results (4.4) of the power spectrum for the Gaussian case, where $h_{0}=8$ and $\delta N=2$.
+The scale of the peak $k_{p}$ is the one that has maximal $h$ at the time when the modes are growing at $N_{v}$, i.e., $N_{p_{k}}=N_{v}$. Thus, we have
 $k_{p}=k_{f}\cdot\frac{3}{4}\sqrt{\frac{8h_{0}^{2}\left(3+2h_{0}^{2}\right)}{1+2h_{0}^{2}}}.$ (4.6)
-We show the numerical and analytical results of our calculation in the Figure 3.
-One thing we should mention is that (4.4) is not valid for all range of $h(N)$ because the computation is under the vaild of EFT so the mass term of curvature is small compared to the heavy entropy one (or other, large $h(N)$).
-Hence this estimation work more well for $|m_{s}^{2}(N)|\gtrsim 8H^{2}$ in our Gaussian-profile model.
+We present the numerical and analytical results of our calculation in the Figure 4.
+One point to mention is that (4.4) is not valid for all ranges of $h(N)$ because the computation is under the valid of the EFT, where the mass term of curvature is small compared to that of the heavy entropy (or in other words, for large $h(N)$).
+Hence, this estimation works better for $|m_{s}^{2}(N)|\gtrsim 8H^{2}$ in our Gaussian-profile model.
 Modes with $\boldsymbol{k\ll k_{f}}$
-As Figure 3 shown, for the l.h.s of the peak, where the switch occurs after crossing the horizon, the power spectrum is governed by the long range contributions.
-Hence we can see a shift on the large scale due to the coupling with entropy fluctuation on large scales.
-We numerically found the amplitude of power spectrum on large scale as the function of $h$ and $\delta N$
+As Figure 4 shows, for the l.h.s of the peak, where the switch-on occurs after crossing the horizon, the power spectrum is governed by the long-range contributions.
+Thus we can observe a shift on large scales due to the coupling with entropy fluctuation on large scales.
+We have numerically found the amplitude of the power spectrum on large scale as a function of $h$ and $\delta N$
 $\Delta_{\zeta}^{\text{(LS)}}=\left(-0.00585h\cdot\delta N+0.0537h+0.333\delta N+1.95\right)^{2}.$ (4.7)
-The amplitude of large-scale curvature perturbation depends on $h$ and $\delta N$ linearly and the amplitude of power spectrum $\Delta_{\zeta}^{\text{(LS)}}\gtrsim 10$, i.e., large-scale amplitude is always greater than the one of single-field inflation.
+The amplitude of the large-scale curvature perturbation depends linearly on $h$ and $\delta N$, and the amplitude of the power spectrum $\Delta_{\zeta}^{\text{(LS)}}\gtrsim 10$, i.e., the large-scale amplitude is always greater than that of single-field inflation.
 4.2 Sharp case
-For very small $\delta N$ the Gaussian profile can be approximated to a delta function. For analytically computable, we use the top-hat profile
+For very small $\delta N$ the Gaussian profile can be approximated by a delta function. For analytically computation, we use the top-hat profile
 $h_{\text{TH}}=h_{0}\left[\theta(t-t_{1})-\theta(t-t_{2})\right].$ (4.8)
-The wave number $k_{f}\simeq He^{H(t_{1}+t_{2})/2}$ is the scale that the modes cross the horizon during the $h\neq 0$.
-After the switch-off, the curvature and entropy modes are decoupled.
-Then the solutions of curvature modes are [55]
+The wave number $k_{f}\simeq He^{H(t_{1}+t_{2})/2}$ is the scale at which the modes cross the horizon during the $h\neq 0$.
+After the switch-off, the curvature and entropy modes decouple.
+Then the solutions for the curvature modes are [58]
 $\hat{\zeta}_{c}(t,\boldsymbol{k})=$ $\Big{[}E_{\zeta}u_{k}(t)+F_{\zeta}u_{k}^{*}(t)\Big{]}\hat{a}_{\zeta}(\boldsymbol{k})$
 $+\Big{[}G_{\zeta}u_{k}(t)+H_{\zeta}u_{k}^{*}(t)\Big{]}\hat{a}_{\mathcal{F}}(\boldsymbol{k})+\text{h.c.}(-\boldsymbol{k}),$ (4.9)
-where the annihilation and creation operators of two scalar modes satisfying the commutation relations $[\hat{a}_{\alpha}(\boldsymbol{k}),\hat{a}^{\dagger}_{\beta}(\boldsymbol{k}^{\prime})]=(2\pi)^{3}\delta_{\alpha\beta}\delta(\boldsymbol{k}-\boldsymbol{k}^{\prime})$ , $[\hat{a}^{\dagger}_{\alpha}(\boldsymbol{k}),\hat{a}^{\dagger}_{\beta}(\boldsymbol{k}^{\prime})]=[\hat{a}_{\alpha}(\boldsymbol{k}),\hat{a}_{\beta}(\boldsymbol{k}^{\prime})]=0$ and $u_{k}$ are the mode functions with Bunch-Davies initial conditons
+where the annihilation and creation operators of the two scalar modes satisfy the commutation relations $[\hat{a}_{\alpha}(\boldsymbol{k}),\hat{a}^{\dagger}_{\beta}(\boldsymbol{k}^{\prime})]=(2\pi)^{3}\delta_{\alpha\beta}\delta(\boldsymbol{k}-\boldsymbol{k}^{\prime})$ , $[\hat{a}^{\dagger}_{\alpha}(\boldsymbol{k}),\hat{a}^{\dagger}_{\beta}(\boldsymbol{k}^{\prime})]=[\hat{a}_{\alpha}(\boldsymbol{k}),\hat{a}_{\beta}(\boldsymbol{k}^{\prime})]=0$ and $u_{k}$ are the mode functions with Bunch-Davies initial conditions
 $u_{k}(t)=\frac{iH}{\sqrt{2k^{3}}}\Big{[}1+ik\tau(t)\Big{]}e^{-ik\tau(t)}.$ (4.10)
-Then the power spectrum of curvature perurtbation is given
+The power spectrum of the curvature perturbation is then given
 $\Delta_{\zeta}=|E_{\zeta}-F_{\zeta}|^{2}+|G_{\zeta}-H_{\zeta}|^{2}.$ (4.11)
-The non-trivial period is the $h_{\text{TH}}\neq 0$ (which is denoted as region-II), where two kinds of modes strongly couple to each other.
+The non-trivial period is when $h_{\text{TH}}\neq 0$ (which is denoted as region-II), during which the two types of modes strongly couple.
 We also discuss large-scale modes $k\ll k_{f}$ and small-scale modes $k\sim k_{f}$ separately.
-Figure 4: The numerical and analytical results (D) of power spectrum for the top-hat case, where $\delta N=0.4$ and $h_{0}=24$. The envelope (4.14) is shown as orange dashed curve.
+Figure 5: The numerical and analytical results (D) of the power spectrum for the top-hat case, where $\delta N=0.4$ and $h_{0}=24$. The envelope (4.14) is shown as orange dashed curve.
 Modes with $\boldsymbol{k\gtrsim k_{f}}$
-We simply use the conformal time as the function of $t$: $\tau(t)=-e^{-Ht}/H$.
-During $t_{1}$ and $t_{2}$ with $\delta t\equiv t_{2}-t_{1}\ll H^{-1}$, one can discard the Hubble friction of the equation of motions.
+We use the conformal time as a function of $t$: $\tau(t)=-e^{-Ht}/H$.
+During $t_{1}$ and $t_{2}$ with $\delta t\equiv t_{2}-t_{1}\ll H^{-1}$, one can ignore the Hubble friction in the equation of motions.
 After $t>t_{1}$, the coupling is switch-on.
-The duration of $h\neq 0$ is too short $\delta t\ll H^{-1}$ that we can ignore the evolution of the scale factor.
+The duration of $h\neq 0$ is too short, $\delta t\ll H^{-1}$, that we can ignore the evolution of the scale factor.
 Hence we can ignore the Hubble frictions and curvature mass.
-Then we can use the equations of motion (3.1) and the WKB-type solutions
+We then use the equations of motion (B) and the WKB-type solutions:
 $\hat{\zeta}_{c}^{\text{II}}(t,\boldsymbol{k})=\sum_{\pm}$ $\Big{(}A_{\pm}^{\zeta}e^{i\omega_{\pm}t}+B_{\pm}^{\zeta}e^{-i\omega_{\pm}t}\Big{)}\hat{a}_{\zeta}(\boldsymbol{k})+\Big{(}C_{\pm}^{\zeta}e^{i\omega_{\pm}t}+D_{\pm}^{\zeta}e^{-i\omega_{\pm}t}\Big{)}\hat{a}_{\mathcal{F}}(\boldsymbol{k})$
@@
 […]
-where the frequencies are given by (3.4). From the equations of motion, we have relations of the coefficients
+where the frequencies are given by (B.3). From the equations of motion, we have relations for the coefficients
 $A_{\pm}^{\mathcal{F}}(C_{\pm}^{\mathcal{F}})=i\frac{\omega_{\pm}^{2}-p^{2}}{\sqrt{2}h\mathcal{A}H\omega_{\pm}}A_{\pm}^{\zeta}(C_{\pm}^{\zeta}),$
 $B_{\pm}^{\mathcal{F}}(D_{\pm}^{\mathcal{F}})=-i\frac{\omega_{\pm}^{2}-p^{2}}{\sqrt{2}h\mathcal{A}H\omega_{\pm}}B_{\pm}^{\zeta}(D_{\pm}^{\zeta}).$
-During the short region-II, where the scale factor is nearly invariant, we have $a=a(N_{f})=k_{f}/H$ hence $p=(k/k_{f})H$ in the above equations.
-We have eight unknown coefficients $A_{\pm}^{\zeta}$, $B_{\pm}^{\zeta}$, $C_{\pm}^{\zeta}$ and $D_{\pm}^{\zeta}$ in the region-II and we have eight boundary conditions at $t_{1}$ hence all these coefficients can be completely solved.
-Then at $t_{2}$ we also have eight boundary condictions and the coefficients of (4.2) can be solved to (see Appendiex D)
+During the short region-II, where the scale factor is nearly invariant, we have $a=a(N_{f})=k_{f}/H$, hence $p=(k/k_{f})H$ in the above equations.
+There are eight unknown coefficients $A_{\pm}^{\zeta}$, $B_{\pm}^{\zeta}$, $C_{\pm}^{\zeta}$ and $D_{\pm}^{\zeta}$ in the region-II, and eight boundary conditions at $t_{1}$, so these coefficients can be completely solved.
+Then at $t_{2}$, we also have eight boundary conditions and the coefficients of (4.2) can be solved to (see Appendix D)
 $E_{\zeta}$ $\simeq\frac{\left(S_{-}+\kappa\right)^{2}}{4\kappa S_{-}}e^{i2\kappa\sinh(\delta N/2)-iS_{-}\delta N},$
@@
 We show the agreement between numerical and completed analytical results (see appendix D) in Figure 5.
-The power spectrum has exponential enhancement near $k_{f}$ and also rapidly oscillate during these scales.
-We found $E_{\zeta}$ and $F_{\zeta}$ are dominated so the power spectrum can be simply written as $\Delta_{\zeta}\simeq\Delta_{\text{env}}\times\left(\text{rapid oscillation}\right)$, where the envelope is
+The power spectrum has exponential enhancement near $k_{f}$ and also rapidly oscillate at these scales.
+We found that $E_{\zeta}$ and $F_{\zeta}$ are dominated, so the power spectrum can be simply written as $\Delta_{\zeta}\simeq\Delta_{\text{env}}\times\left(\text{rapid oscillation}\right)$, where the envelope is
 $\Delta_{\text{env}}=\left|\frac{e^{-i2S_{-}\delta N}\left(\kappa^{2}-|S_{-}|^{2}\right)}{4|S_{-}|^{2}}\right|$ (4.14)
 and the rapid oscillation is characterized by $\text{Exp}\left(i2\kappa e^{\pm\delta N/2}\right)$, which is only related to the period of non-zero $h_{\text{tp}}$.
-The enhancement of power spectrum is provided by low-frequency $\omega_{-}$ and period $\delta N$ together.
-The scale of peak can be nearly determined by the exponential factor in $\Delta_{\text{env}}$, which is estimated by
+The enhancement of the power spectrum is provided by the low-frequency $\omega_{-}$(see Appendix B) and period $\delta N$ together.
+The scale of the peak can be nearly determined by the exponential factor in $\Delta_{\text{env}}$, which is estimated by
 $k_{p}=k_{f}\cdot\frac{h_{0}\sqrt{30+104h_{0}^{2}+56h_{0}^{4}}}{2(1+2h_{0}^{2})}.$ (4.15)
-These calculations are only valid near scale $k\gtrsim k_{f}$ and for $k\ll k_{f}$ $\Delta_{\zeta}=1$, which miss the matching with the numerical results.
+These calculations are only valid near the scale $k\gtrsim k_{f}$ and for $k\ll k_{f}$ $\Delta_{\zeta}=1$, which miss the matching with the numerical results.
 Modes with $\boldsymbol{k\ll k_{f}}$
-Figure 5: In the top-hat case, the analytical result of large-scale power spectrum as function of $\delta N$ (4), where $h_{0}=24$. The envelope (4.19) is shown as orange dashed curve.
-On large scales, the momentum terms in equations of motion can be ignored and the mass terms and all coupling terms become important. The equations of motion can be solved as
+Figure 6: In the top-hat case, the analytical result of the large-scale power spectrum as a function of $\delta N$ (4), where $h_{0}=24$. The envelope (4.19) is shown as the orange dashed curve.
+On large scales, the momentum terms in the equations of motion can be ignored, and the mass terms and all coupling terms become important. The equations of motion can be solved as
 $\hat{\zeta}_{c}^{\text{II}}(\tau,\boldsymbol{k})=$ $\sum_{i=\zeta,\mathcal{F}}\left[c_{0}^{i}+c_{1}^{i}x^{3}+c_{+}^{i}x^{1+p_{+}}+c_{-}^{i}x^{1+p_{-}}\right]\hat{a}_{i}(\boldsymbol{k})+\text{h.c.}(-\boldsymbol{k}),$
@@
 $d_{+}^{i}=\frac{p_{-}(1+2h^{2})-2}{2\sqrt{2}h(3+4h^{2})}c_{+}^{i},$ $\ \ \ \ \ d_{-}^{i}=\frac{p_{+}(1+2h^{2})-2}{2\sqrt{2}h(3+4h^{2})}c_{-}^{i}.$ (4.17)
-We have eight unknown coefficients $c_{0}^{i}$, $c_{1}^{i}$, $c_{+}^{i}$ and $c_{-}^{i}$ ($i=\zeta,\mathcal{F}$) in the region-II and eight boundary conditions at $t_{1}$ hence all these coefficients can be completely solved.
+We have eight unknown coefficients $c_{0}^{i}$, $c_{1}^{i}$, $c_{+}^{i}$ and $c_{-}^{i}$ ($i=\zeta,\mathcal{F}$) in region-II and eight boundary conditions at $t_{1}$, so all of these coefficients can be completely solved.
 Then at $t_{2}$ the coefficients of (4.2) can also be solved [^4] We found that this computation breaks when $x\lesssim 10^{-5}$..
-The results are extremely tedious, but we can simple them under some approximations, see appendix D.
-The results should be real because the perturbations become classical on super-horizon.
+The results are extremely tedious, but we can simplify them under some approximations, see appendix D.
+The results should be real because the perturbations become classical on the super-horizon.
 So we can only consider the real parts
@@
 We found that, for fixed $\delta N$, $|G_{\zeta}-H_{\zeta}|$ oscillates around zero when changing $h$.
-The oscillation is characterized by $\sin(2\sqrt{6}h\delta N)$ and is not significant on amplitude.
-This is because, on large scales, the dynamics of curvature perturbation is almost a constant and weakly depends on the frequency.
+The oscillation is characterized by $\sin(2\sqrt{6}h\delta N)$ and is not significant in amplitude.
+This is because, on large scales, the dynamics of the curvature perturbation is almost constant and weakly depends on the frequency.
 For large $h$, the dynamics is determined by an EFT of massless curvature perturbation with sound speed $c_{s}\simeq-1/3$.
@@
 as Figure 6 shows.
-We can see $\Delta_{\zeta}<1$, in other words, the amplitude of power spectrum on large scales is always smaller than the one of single-field inflation and is not sensitive to $h$ for sharp case.
-We should mention that this result is also valid for board case with top-hat profile of $h$.
+We can see $\Delta_{\zeta}<1$, in other words, the amplitude of the power spectrum on large scales is always smaller than that of single-field inflation and is not sensitive to $h$ for sharp case.
+We should mention that this result is also valid for broad case with a top-hat profile of $h$.
 For $\delta N\gg 1$, we have $\Delta_{\text{env}}^{\text{(LS)}}\simeq 0.034$, which is not sensitive to $\delta N$ either.
 4.3 PBH formation
-After treating the power spectrum on large and small scales for board and sharp case, we can also compare them with the amplitude of the curvature power spectrum required for PBH formation.
-The fraction of PBHs at formation time can be estimated in a model-independent way [59]
+After treating the power spectrum on large and small scales for the broad and sharp case, we can also compare them with the amplitude of the curvature power spectrum required for PBH formation.
+The fraction of PBHs at the formation time can be estimated in a model-independent way [62]
 $\beta_{\text{PBH}}=3.7\times 10^{-9}\left(\frac{\gamma}{0.2}\right)^{-1/2}\left(\frac{g_{*,f}}{10.75}\right)^{1/4}\left(\frac{M_{\text{PBH}}}{M_{\odot}}\right)^{1/2}f_{\text{PBH}},$ (4.20)
-where $\gamma$ can be calculated analytically as $\gamma=0.2$ [51], $g_{*,f}$ is the effective number of relativistic degrees of freedom at formation time, $M_{\text{PBH}}$ is the mass of PBHs at formation time and $f_{\text{PBH}}\equiv\Omega_{\text{PBH}}/\Omega_{\text{DM}}$ is the present-day fraction of PBH density.
-On the other hand, the $\beta$ can be related to initial conditions of matter power spectrum through collapse theory of Newtonian gravity.
-The well-used model is the Press-Schechter one [60].
-After assuming the Gaussian distribution of overdensity and simply taking the critical density contrast as the threshold $\delta_{c}=0.4$ [61], the $\beta$ can be obtained to
+where $\gamma$ can be calculated analytically as $\gamma=0.2$ [54], $g_{*,f}$ is the effective number of relativistic degrees of freedom at formation time, $M_{\text{PBH}}$ is the mass of PBHs at formation time, and $f_{\text{PBH}}\equiv\Omega_{\text{PBH}}/\Omega_{\text{DM}}$ is the present-day fraction of PBH density.
+On the other hand, the $\beta$ can be related to the initial conditions of the matter power spectrum through the collapse theory of Newtonian gravity.
+Acommonly used model is the Press-Schechter one [63].
+After assuming a Gaussian distribution of overdensity and taking the critical density contrast as the threshold $\delta_{c}=0.4$ [64], $\beta$ can be obtained as
 $\beta_{\text{PBH}}\simeq\frac{\sigma}{\sqrt{2\pi}\delta_{c}}\text{Exp}{\left(-\frac{\delta_{c}^{2}}{2\sigma^{2}}\right)},$ (4.21)
 where $\sigma^{2}\ll\delta_{c}^{2}$ is the variance of the distribution.
-Smaller $\sigma$ implies less contrast exceeding the critical value, resulting in less $\beta_{\text{PBH}}$.
-The variance can be calculated from primordial power spectrum as [62]
+A smaller $\sigma$ implies fewer contrasts exceeding the critical value, resulting in a smaller $\beta_{\text{PBH}}$.
+The variance can be calculated from the primordial power spectrum as [65]
 $\sigma^{2}(R)\simeq\frac{16}{81}\int d\ln q\ W^{2}(q,R)(qR)^{4}\mathcal{P}_{\zeta}(q),$ (4.22)
-where the window function $W(q,R)$ is selected to Gaussian or top-hat usually and the comoving horizon size of PBH formed is related to the mass as $R^{-1}=k_{p}$, where [63]
+where the window function $W(q,R)$ is usually selected as Gaussian or top-hat, and the comoving horizon size of the PBH formed is related to the mass as $R^{-1}=k_{p}$, where [66]
 $M_{\text{PBH}}=30M_{\odot}\left(\frac{\gamma}{0.2}\right)\left(\frac{g_{*,f}}{10.75}\right)^{-1/6}\left(\frac{k_{p}}{2.9\times 10^{5}\text{Mpc}^{-1}}\right)^{-2}.$ (4.23)
-For inflation with a triad of vector fields, where both large-scale and PBH-scale are not trivial, the enhancement at PBH scale comparing with CMB scale is
+For inflation with a triad of vector fields, where both large-scale and PBH-scale are not trivial, the enhancement at the PBH scale comparing to the CMB scale is
 $\Delta\mathcal{P}_{\zeta}\equiv\frac{\mathcal{P}_{\zeta}(k_{p})}{\mathcal{P}_{\zeta}(k_{\text{CMB}})}=\frac{\Delta_{\zeta}(k_{p})}{\Delta_{\zeta}^{\text{(LS)}}}.$ (4.24)
-Tne power spectrum at CMB scale is $\mathcal{P}_{\zeta}(k_{\text{CMB}})=2\times 10^{-9}$.
-For board case this is given by (4.4) and (4.7), while for sharp case, the rapid oscillation will be smoothed out by window function and we can use (4.14) and (4.19).
+The power spectrum at the CMB scale is $\mathcal{P}_{\zeta}(k_{\text{CMB}})=2\times 10^{-9}$.
+For the broad case, this is given by (4.4) and (4.7), while for the sharp case, the rapid oscillation will be smoothed out by the window function and we can use (4.14) and (4.19).
 Figure 7: The fraction of PBHs at formation $\beta_{\text{PBH}}$ v.s.
-height of peaks $h$.
-Solid curves denote the cases of $\delta N=2$ for $h_{\text{G}}$ and $\delta N=0.4$ for $h_{\text{TH}}$, while dashed curves denote the cases of smaller $\delta N$.
-Orange region denotes regime that $f_{\text{PBH}}=1$ between $10^{-17}M_{\odot}<M_{\text{PBH}}<10^{-12}M_{\odot}$.
-If we consider the window where the PBHs can be the candidate of dark matter: $10^{-17}M_{\odot}<M_{\text{PBH}}<10^{-12}M_{\odot}$ and $f_{\text{PBH}}=1$, from (4.20) we obtain $1.2\times 10^{-17}<\beta_{\text{PBH}}<3.7\times 10^{-15}$.
-Then from (4.21) the variance for PBH formation in this mass regimes is $0.0510.047<\sigma<$ 0.051.
+the height of the peaks $h$.
+The solid curves denote the cases of $\delta N=2$ for $h_{\text{G}}$ and $\delta N=0.4$ for $h_{\text{TH}}$, while the dashed curves denote the cases of smaller $\delta N$.
+THe orange region denotes the regime where $f_{\text{PBH}}=1$ between $10^{-17}M_{\odot}<M_{\text{PBH}}<10^{-12}M_{\odot}$.
+If we consider the window where the PBHs can be the a candidate of dark matter: $10^{-17}M_{\odot}<M_{\text{PBH}}<10^{-12}M_{\odot}$ and $f_{\text{PBH}}=1$, from (4.20) we obtain $1.2\times 10^{-17}<\beta_{\text{PBH}}<3.7\times 10^{-15}$.
+Then, from (4.21) the variance for PBH formation in this mass regimes is $0.0510.047<\sigma<$ 0.051.
 When the power spectrum is peaked at $k_{p}$, the variance is reduced to $\sigma^{2}\simeq\mathcal{P}_{\zeta}(k_{p})/5$.
-It’s seem that we need $\mathcal{P}_{\zeta}(k_{p})\sim 10^{-2}$ between $10^{12}\text{Mpc}^{-1}<k_{p}<10^{15}\text{Mpc}^{-1}$ to generate sufficient PBHs.
-However, because the profile of the peak is very sensitive to $h$ and $\delta N$ and the large-scale power spectrum also depends on these parameters.
-And the fraction $\beta_{\text{PBH}}$ is exponentially dependent on the power spectrum.
-Slightly changing profile of the peak will lead to significant change in $\beta_{\text{PBH}}$.
-We show how $\beta_{\text{PBH}}$ depend on parameters in Figure 6 for board Gaussian peaks and sharp top-hat peaks.
-We have used $R=k_{p}^{-1}$, which are the scales of peaks and can be separately obtained through (4.6) and (4.15) for Gaussian and top-hat examples.
-The $\beta_{\text{PBH}}$ decays rapidly for smaller parameter, i.e., the smaller height and fatness of the peaks.
-Also, the difference of using different window functions is also distinguishable.
-The mass spectrum of PBHs are shown in Figure 7 for board Gaussian peaks and in Figure 8 for sharp top-hat peaks. We choose the parameters of these examples as following table.
-Figure 7: Abundance of PBHs as dark matter for board($\delta N\gtrsim 1$) Gaussian profile of peak $h_{\text{G}}$.
-Figure 8: Abundance of PBHs as dark matter for shape($\delta N\lesssim 1$) top-hat profile of peak $h_{\text{TH}}$.
+It seems that we need $\mathcal{P}_{\zeta}(k_{p})\sim 10^{-2}$ between $10^{12}\text{Mpc}^{-1}<k_{p}<10^{15}\text{Mpc}^{-1}$ to generate sufficient PBHs.
+However, because the profile of the peak is very sensitive to $h$ and $\delta N$, and the large-scale power spectrum also depends on these parameters.
+The fraction $\beta_{\text{PBH}}$ is exponentially dependent on the power spectrum.
+Slightly changing the profile of the peak will lead to a significant change in $\beta_{\text{PBH}}$.
+We show how $\beta_{\text{PBH}}$ depends on parameters in Figure 7 for broad Gaussian peaks and sharp top-hat peaks.
+We have used $R=k_{p}^{-1}$, which are the scales of the peaks and can be separately obtained through (4.6) and (4.15) for Gaussian and top-hat examples.
+The $\beta_{\text{PBH}}$ decays rapidly for smaller parameters, i.e., smaller height and fatness of the peaks.
+Also, the difference in using different window functions is also distinguishable.
+Figure 8: Abundance of PBHs as dark matter for broad($\delta N\gtrsim 1$) Gaussian profile of peak $h_{\text{G}}$.
+Figure 9: Abundance of PBHs as dark matter for sharp($\delta N\lesssim 1$) top-hat profile of peak $h_{\text{TH}}$.
+The mass spectrum of PBHs are shown in Figure 8 for broad Gaussian peaks and in Figure 9 for sharp top-hat peaks. We choose the parameters of these examples as following table.
 ($h_{\text{G}}$, $W_{\text{G}}$) ($h_{\text{G}}$, $W_{\text{TH}}$) ($h_{\text{TH}}$, $W_{\text{G}}$) ($h_{\text{TH}}$, $W_{\text{TH}}$)
@@
 $\delta N$ 2 1.5 0.4 0.395
-For solid curves, we choose the parameters so that the PBHs are responsible for total dark matter $f_{\text{PBH}}=1$ in our universe.
-The time of non-vanishing vector fields are chosen as $N_{f}=31.5$ for all solid examples to avoid constraints from observations of evaporation of PBHs due to Hawking radiation (EVAP) [59] and microlensing of stars in M31 by Subaru Hyper SuprimeCam (HSC) [64].
-For dashed curves, we choose different parameters so that $f_{\text{PBH}}<1$ and also different $N_{f}=25.3$, which corresponds to larger PBH mass formation.
+For the solid curves, we choose the parameters so that the PBHs are responsible for the total dark matter $f_{\text{PBH}}=1$ in our universe.
+The time of non-vanishing vector fields are chosen as $N_{f}=31.5$ for all solid examples to avoid constraints from observations of the evaporation of PBHs due to Hawking radiation (EVAP) [62] and microlensing of stars in M31 by Subaru Hyper SuprimeCam (HSC) [67].
+For the dashed curves, we choose different parameters so that $f_{\text{PBH}}<1$ and also different $N_{f}=25.3$, which corresponds to the formation of larger PBH mass.
 We have seen in Figure 7 the abundance is sensitive to these parameter.
 As the first and the third column show, a mere 1% decrease in $h$ leads to a drastic drop in abundance from 1 to $10^{-5}$.
-For top-hat profile of peak $h_{\text{TH}}$, the abundance is also sensitive to $\delta N$ (4.14).
+For the top-hat profile of the peak $h_{\text{TH}}$, the abundance is also sensitive to $\delta N$ (4.14).
 So we can see in the last column of table, 1% decrease in $\delta N$ also reduces $f_{\text{PBH}}$ to $10^{-5}$.
-We also compare the results of using Gaussian and top-hat window function.
-The choice of way of smoothing out the fluctuations increases the uncertainty in the results of PBH formation [65, 66].
+We also compare the results of using Gaussian and top-hat window functions.
+The choice of the method of smoothing out the fluctuations increases the uncertainty in the results of PBH formation [68, 69].
 We demonstrate a significant difference in the parameters required when using different window functions.
-For example, the first row in table, To generate the same abundance of PBHs, the required $h$ differs by up to 20% when using Gaussian and top-hat window functions.
+For example, in the first row of the table, to generate the same abundance of PBHs, the required $h$ differs by up to 20% when using Gaussian and top-hat window functions.
 We usually need a larger primordial density in Gaussian smoothing.
 Furthermore, we also find that using different window functions results in variations in the peak position of $f_{\text{PBH}}$.
-The mass of formation is usually greater when using top-hat window function.
-In the examples, this difference reaches up to $M_{\text{peak,TH}}/M_{\text{peak,G}}\sim 10$.
+The mass of formation is usually greater when using the top-hat window function.
+In these examples, this difference reaches up to $M_{\text{peak,TH}}/M_{\text{peak,G}}\sim 10$.
 5 Discussion
@@
 The Gaussian is employed to discuss the broad case, while the top-hat is used to address the sharp case, as both yield more precise analytical results.
-Due to the non-decaying entropy perturbations, the large-scale power spectrum is shifted compared to the single-field one.
-The amplitude of the shift depends on the details of vector fields dynamics, i.e., $\delta N$ and $h$, as shown in (4.7) and (4.19).
-For the shape top-hat case, this amplitude is only a function of $\delta N$, allowing for an analytical calculation of the shape of the power spectrum.
-Given the exponential dependence of the power spectrum shape on these parameters and the subsequent exponential dependence of PBH abundance on the power spectrum shape, the PBH abundance is highly sensitive to these parameters.
+Due to the non-decaying entropy perturbations, the large-scale power spectrum is shifted compared to the single-field case.
+The amplitude of this shift depends on the dynamics of the vector fields, i.e., $\delta N$ and $h$, as shown in (4.7) and (4.19).
+For the sharp top-hat case, this amplitude is only a function of $\delta N$, allowing for an analytical calculation of the shape of the power spectrum.
+Given the exponential dependence of the power spectrum’s shape on these parameters and the subsequent exponential dependence of PBH abundance on the power spectrum’s shape, the PBH abundance is highly sensitive to these parameters.
 Whether it’s $h$ or $\delta N$, even slight variations can significantly reduce the abundance of PBHs.
-Certainly, this sensitivity is also related to the specific details of the PBH collapse.
-For instance, we demonstrated that under the same $\delta_{c}$, smoothing the same fluctuations in different ways can lead to vastly different results, not only in PBH abundance but also in the masses of PBH formation.
+Certainly, this sensitivity is also related to the specific details of PBH collapse.
+For instance, we demonstrated that, under the same $\delta_{c}$, smoothing the same fluctuations in different ways can lead to vastly different results, not only in PBH abundance but also in the masses of the PBH formation.
 For example, compared to a Gaussian window function, using a top-hat window function typically results in PBHs with smaller masses.
@@
 In this scenario, the contributions to inflationary predictions differ significantly from the case where $h$ is very small.
-Many questions need to be reconsidered in this regime, such as the anisotropy for one vector field and non-Gaussianity.
+Many questions need to be reconsidered in this regime, such as the anisotropy for a single vector field, and non-Gaussianity.
 We posit that the discussions in this paper are also applicable to a single vector field.
-In this case, the $IN^{2}$ term in power spectrum, which served as a contribution to anisotropy for small $h$ case, is replaced by other constant modes.
+In this case, the $IN^{2}$ term in the power spectrum, which served as a contribution to anisotropy in the small $h$ case, is replaced by other constant modes.
 It is imperative for us to calculate the anisotropy in this regime.
 Additionally, when considering a charged field as the inflaton, whether the perturbative Schwinger effect from the electric field is crucial to inflationary predictions is also a question that warrants consideration.
-Regarding PBHs, we are discussing the isotropic configuration of vector fields as an example.
+Regarding PBHs, we focus on the isotropic configuration of the vector fields as an example.
 Although it bears similarities to hyperbolic inflation, in general, vector fields themselves exhibit phenomenology distinct from scalar fields, such as anisotropy for a single vector field and chirality for symmetry breaking.
-These differences manifest in the generated PBHs and induced gravitational waves.
+These differences manifest in the generated PBHs and the induced gravitational waves.
 Recently observations from 15-year NANOGrav data set reveals information about low-frequency stochastic gravitational wave background [70].
@@
 Also, we assume the slow-roll condition that $V_{\phi\phi}\ll H^{2}$ hence the third term in $m_{s}^{2}$ also can be ignored.
-Appendix B Effective field theory of inflation
-In the EFT formalism, the physical scalar d.o.f at low-energy scale is the Nambu-Goldstone (NG) boson particle coming from breaking the time translation symmetry $t\to t+\xi^{0}$ of the de Sitter space by a clock field $\phi(t)$, i.e., the homogeneous inflaton.
-One can choose the slice of $t$ that the inflaton has no excitation all the time, i.e.
-the unitary gauge.
-In this gauge, the only dynamical scalar d.o.f at low-energy is the scalar part of graviton $g_{ij}=a^{2}(t)e^{2\zeta}\delta_{ij}$, where the scalar $\zeta(t,\boldsymbol{x})$ is the spatial curvature of the time slice.
-The EFT of a single scalar field in the unitary gauge can be written as [39]
-$\mathcal{L}_{\text{EFT}}=\sqrt{-g}\left[\frac{M_{\text{pl}}^{2}}{2}R+M_{\text{pl}}^{2}\dot{H}g^{00}-M_{\text{pl}}^{2}\left(3H^{2}+\dot{H}\right)+\cdots\right]+\mathcal{L}_{\text{n-u}},$ (B.1)
-where the dots stand for the higher-order terms of $(1+g^{00})$ and we have ignored the extrinsic curvature fluctuation $\delta K_{\mu\nu}$.
-$\mathcal{L}_{\text{n-u}}$ is the non-universal parts and for $\omega^{2}\ll M^{2}+p^{2}$ [40]
-$\mathcal{L}_{\text{n-u}}=\sqrt{-g}\sum_{n=2}\frac{M_{n}^{4}}{n!}\left[(1+g^{00})\frac{M^{2}}{M^{2}-\tilde{\partial}^{2}}\right]^{n-1}(1+g^{00})+\cdots,$ (B.2)
-where $M$ is the mass of heavy fields and ~$\tilde{\partial}\equiv\partial/a$.
-These operators are non-local generally due to the hierarchy between the frequency and momentum when the speed of sound is small $c_{s}\ll 1$.
-There exist regime in energy scales that the dispersion is non-linear [40, 41].
-If $\omega\sim p$ the momentum is also $p\ll M$ and we can ignore the non-local operators.
-In order to transform to the spatial flat gauge $g_{ij}=a^{2}(t)\delta_{ij}$, one can perform a time diffeomorphisms $t\to\tilde{t}=t-\pi(t,\boldsymbol{x})$.
-Then at linear-order the NG boson is related to the curvature fluctuation as $\zeta=-H\pi$.
-The time component of the metric transforms as
-$g^{00}=\frac{\partial x^{0}}{\partial\tilde{x}^{\mu}}\frac{\partial x^{0}}{\partial\tilde{x}^{\nu}}\tilde{g}^{\mu\nu}=\left(1+\dot{\pi}\right)^{2}\tilde{g}^{00}+2\left(1+\dot{\pi}\right)\partial_{i}\pi\tilde{g}^{0i}+\tilde{g}^{ij}\partial_{i}\pi\partial_{j}\pi$ (B.3)
-In the low-energy scale $\omega^{2}\ll p^{2}+M^{2}$, we can write down the action in the re-parameterization of time. Dropping the tildes for simplicity, the Lagrangian of NG boson is given by [40]
-$\mathcal{L}_{\pi}=$ $\sqrt{-g}\Bigg{\{}-M_{\text{pl}}^{2}\dot{H}\left[\dot{\pi}\left(1+\frac{2M_{2}^{4}}{M_{\text{pl}}^{2}|\dot{H}|}\frac{M^{2}}{M^{2}-\tilde{\partial}^{2}}\right)\dot{\pi}-(\tilde{\partial}\pi)^{2}\right]$
-$+2M_{2}^{4}\left[\dot{\pi}^{2}-(\tilde{\partial}\pi)^{2}\right]\frac{M^{2}}{M^{2}-\tilde{\partial}^{2}}\dot{\pi}-\frac{4}{3}M_{3}^{4}\left(\dot{\pi}\frac{M^{2}}{M^{2}-\tilde{\partial}^{2}}\right)^{2}\dot{\pi}+\cdots\Bigg{\}},$ (B.4)
-where we have expanding the action up to third order of fluctuations.
-In the decoupling limit, the NG boson is decoupled with gravity and becomes a dynamical scalar mode.
-For the inflation, this decoupling limit is the slow-roll limit $\epsilon\to 0$.
-Then the quadratic Lagrangian of NG boson reduces to
-$\mathcal{L}_{\pi}=a^{3}M_{\text{pl}}^{2}\epsilon\left[\dot{\pi}\left(\frac{M^{2}c_{s}^{-2}-\tilde{\partial}^{2}}{M^{2}-\tilde{\partial}^{2}}\right)\dot{\pi}-(\tilde{\partial}\mathcal{\pi})^{2}\right],$ (B.5)
-where the speed of sound is defined by
-$\frac{1}{c_{s}^{2}}\equiv 1-\frac{2M_{2}^{4}}{M_{\text{pl}}^{2}\dot{H}}.$ (B.6)
-The modification of the speed of sound is provided by the higher operator $M_{2}$ and we will show soon that this can be obtained by integrating out the heavy fluctuation at low energy in the dilaton-gauge model.
-The dispersion relation of the propagation of the light field $\pi$ should be written as
-$\omega^{2}(p)=\frac{M^{2}+p^{2}}{M^{2}c_{s}^{-2}+p^{2}}p^{2},$ (B.7)
-which is generally non-linear. However, if the speed of sound $c_{s}$ is not very small, The region with nonlinear dispersion relations is very narrow so the Lagrangian of NG boson shortly reduces to
-$\mathcal{L}_{\pi}=a^{3}M_{\text{pl}}^{2}H^{2}\frac{\epsilon}{c_{s}^{2}}\left[\dot{\pi}^{2}-\frac{c_{s}^{2}}{a^{2}}(\partial\mathcal{\pi})^{2}\right],$ (B.8)
-where the dispersion relation is almost linear $\omega(p)\simeq c_{s}p$ throughout the entire process.
+Appendix B Modified dispersion relation
+We are interested in the decoupling limit where the nontrivial solution $\zeta=\text{constant}$ is permitted hence we drop out the small slow-roll parameters.
+Then we have $\mathcal{A}=4$ and $\mathcal{B}=m_{\sigma}/H^{2}$.
+We regard $\dot{\sigma}$ and $h$ as constants.
+To find the dispersion relation we study the short-wavelength limit where the Hubble friction terms can be disregarded.
+We have noted that there is also a mass term for curvature fluctuation.
+The upper bound of the mass is $m_{\sigma}^{2}\lesssim 8H^{2}$, i.e., the mass of $\zeta$ is at most of the same order as the energy scale of inflation.
+In the short-wavelength limit the contribution of $m_{\sigma}$ term can also be ignored.
+Then the equations of motion read
+$\ddot{\zeta_{c}}+p^{2}\zeta_{c}=$ $\sqrt{2}hH\mathcal{A}\dot{\mathcal{F}},$
+¨$\ddot{\mathcal{F}}+p^{2}\mathcal{F}+m_{s}^{2}\mathcal{F}=$ $-\sqrt{2}hH\mathcal{A}\dot{\zeta}_{c},$ (B.1)
+where $p=k/a$ is the physical momentum and $\zeta_{c}\equiv M_{\text{pl}}\sqrt{2\epsilon}\zeta$ is the canonical variable. The schematic solutions of these equations can be written as [45]
+$\zeta_{c}=\zeta_{+}e^{i\omega_{+}t}+\zeta_{-}e^{i\omega_{-}t},$
+$\mathcal{F}=\mathcal{F}_{+}e^{i\omega_{+}t}+\mathcal{F}_{-}e^{i\omega_{-}t},$ (B.2)
+where the two frequencies $\omega_{\pm}$ are given by
+$\omega_{\pm}^{2}=\frac{m_{s}^{2}}{2c_{s}^{2}}+p^{2}\pm\frac{m_{s}^{2}}{2c_{s}^{2}}\sqrt{1+\frac{4\left(1-c_{s}^{2}\right)p^{2}}{m_{s}^{2}c_{s}^{-2}}},$ (B.3)
+where the speed of sound is given by
+$\frac{1}{c_{s}^{2}}=1+\frac{32H^{2}h^{2}}{m_{s}^{2}}.$ (B.4)
+The speed of sound can be given by the dispersion relation of low-frequency modes at low-energy $4(1-c_{s}^{2})p^{2}\ll m_{s}^{2}c_{s}^{-2}$
+$\omega_{-}^{2}\simeq c_{s}^{2}p^{2}.$ (B.5)
+The modification of the dispersion relation of fluctuations is provided by the interaction with $\dot{\zeta}{\mathcal{F}}$.
+The heavy entropy fluctuation implies the hierarchy of the two frequencies $\omega_{-}\ll\omega_{+}$.
+Hence we have a cut-off scale on momentum
+$p_{\text{UV}}^{2}=\frac{|m_{s}^{2}c_{s}^{-2}|}{4\left(1-c_{s}^{2}\right)}.$ (B.6)
+In the low energy scales, the two scalar modes, which are corresponded to two frequencies, oscillate coherently and we integrate out the heavy mode to eliminate the high-frequency mode from the dynamics of the curvature fluctuation.
 Appendix C Growing solution of the adiabatic mode
@@
 […]
-where $\omega_{-}$ is given by (3.4)
+where $\omega_{-}$ is given by (B.3)
 $\frac{|\omega_{-}|}{H}=\sqrt{-\mu^{2}-\frac{1}{2}u+\frac{1}{2}\sqrt{u^{2}+128h^{2}\mu^{2}}}$ (C.2)
@@
 Appendix D Bogoliubov coefficients
-For the shape case, we use the top-hat profile of the $h(N)$ for analytical computations,
+For the sharp case, we use the top-hat profile of the $h(N)$ for analytical computations,
 $h(N)=h_{0}\left[\theta(t-t_{1})-\theta(t-t_{2})\right].$ (D.1)
@@
 After $t>t_{1}$, which is denoted as region-II, these two scalar fields are coupled to each other hence their modes are also mixed with each other.
-The governing equations of motion are given by (3.1).
+The governing equations of motion are given by (3).
 We discuss two different non-trivial momentum regimes: near PBHs peak $k\sim k_{f}$ and large scales $k\ll k_{f}$, respectively.